### v1

# Ctrl-GenAug: Controllable Generative Augmentation for Medical Sequence Classification

Abstract
In the medical field, the limited availability of large-scale datasets and labor-intensive annotation processes hinder the performance of deep models. Diffusion-based generative augmentation approaches present a promising solution to this issue, having been proven effective in advancing downstream medical recognition tasks. Nevertheless, existing works lack sufficient semantic and sequential steerability for challenging video/3D sequence generation, and neglect quality control of noisy synthesized samples, resulting in unreliable synthetic databases and severely limiting the performance of downstream tasks. In this work, we present Ctrl-GenAug, a novel and general generative augmentation framework that enables highly semantic- and sequential-customized sequence synthesis and suppresses incorrectly synthesized samples, to aid medical sequence classification. Specifically, we first design a multimodal conditions-guided sequence generator for controllably synthesizing diagnosis-promotive samples. A sequential augmentation module is integrated to enhance the temporal/stereoscopic coherence of generated samples. Then, we propose a noisy synthetic data filter to suppress unreliable cases at semantic and sequential levels. Extensive experiments on 3 medical datasets, using 11 networks trained on 3 paradigms, comprehensively analyze the effectiveness and generality of Ctrl-GenAug, particularly in underrepresented high-risk populations and out-domain conditions.
Index Terms:
Classification, Controllable Generative Augmentation, Medical Sequence Synthesis, Diffusion Model.
publicationid: pubid: This work has been submitted to the IEEE for possible publication. Copyright may be transferred without notice, after which this version may no longer be accessible.
I Introduction
Dynamic information in medical imaging (e.g., video sequences) plays a vital role in clinical diagnosis. Recent deep learning-based classifiers have shown the ability to improve the diagnostic accuracy of different diseases [1, 2]. Despite rapid advancements, the performance of current advanced solutions is still limited by several issues: 1) The scarcity of dynamic clinical cases, coupled with the high annotation cost, limits data availability; 2) Imbalanced data distribution, driven by the rarity of high-risk positive cases, skews model training; 3) Deep models are prone to brittle performance degradation when tested on out-domain data, e.g., cases from different medical centers [3, 4]. This risk presents a significant obstacle to deploying models in real-world medical settings. To address these issues, there is growing interest in generative augmentation paradigms, which employ advanced generative models to synthesize medical samples, thereby augmenting relevant diagnostic tasks.
Denoising diffusion probabilistic models (DDPMs), which utilize explicit likelihood estimation and a progressive sampling process for data synthesis, have more well-established mathematical explanations and abilities to achieve stable, controllable, and diverse synthesis than previous generative methods like generative adversarial networks (GANs) [5, 6, 7]. Diffusion models have garnered remarkable success in several natural image fields, including static image generation [8, 9], video synthesis [10, 11, 12, 13, 14, 15, 16, 17], video editing [18, 19, 20], and image animation [21, 22]. Recently, researchers have explored high-fidelity medical sequence synthesis by applying the above approaches within the medical field [23, 24].
Most recently, pioneers deeply investigated the impact of the diffusion-based generative augmentation scheme on solving data scarcity [7, 25] and domain generalization [25] issues in diagnostic tasks. However, they failed to fully ensure the reliability of synthetic medical images for classification due to limited semantic steerability and absent quality control over the generated images. Moreover, their techniques focus on image-level synthesis rather than sequence-level synthesis, which is crucial for medical modalities like MRI, ultrasound (US), etc [7]. Thus, our study plans to devise a diagnosis-reliable controllable generative augmentation framework to facilitate accurate and robust medical sequence classification.
Figure 1: Datasets description: (a) Carotid US videos with various stenosis gradings. Red and yellow boxes represent plaques and residual lumens, respectively. (b) Thyroid US videos with different TI-RADS levels, where blue boxes indicate nodules. (c) Cardiac MRI volumes with distinct diseases. Three key anatomical structures associated with diagnosis are highlighted with masks, including the left ventricle (blue), myocardium (green), and right ventricle (red).
Intuitively, building customized, high-fidelity, and sequentially coherent synthetic medical databases, and effectively utilizing them, is essential for enhancing sequence recognition. However, this task presents several challenges. First, lesions (Fig. 1(a-b)) or structures (Fig. 1(c)) of the same disease category exhibit large visual variances (e.g., shapes, intensities, positions, etc). This may confuse diffusion learning seriously, thus causing uncontrollable and unreliable generation. Second, artifacts and noises (e.g., US speckle noise) may prevent models from accurately perceiving vital anatomical targets, hindering the synthesis of high-fidelity sequences. Third, the complex dynamic changes (Fig. 1(a-b)) and the varying sizes of anatomical regions (Fig. 1(c)) challenge the coherence of synthesized sequences. The fourth challenge lies in the domain gap between synthetic and real samples, where achieving real-domain customization is crucial for effectively utilizing synthetic samples in downstream learning [26]. Lastly, even a well-designed sequence generator cannot always guarantee high-quality synthesis due to random sampling, and unsatisfactory synthetic samples may negatively impact subsequent classifier learning. In summary, from model- and data-centric perspectives, the challenges of the task can be concluded as: 1) How can we design the generator architecture to achieve satisfactory sequence synthesis? 2) How can we discriminatively filter out potentially harmful synthetic samples?
In this work, we propose a novel controllable generative augmentation framework, named Ctrl-GenAug, to facilitate medical sequence classification tasks on different organs and modalities. Concretely, Ctrl-GenAug synthesizes customized, high-quality sequences using conditional diffusion models and applies quality control to the generated sequences, to supplement medical datasets and improve classification. This is a general method for enhancing medical sequence classification. We believe this is the first comprehensive study to analyze the impact of the controllable generative augmentation scheme on promoting multi-modal medical sequence classification. Our contributions are three-fold:
  • We propose a multimodal conditions-guided sequence generator to ensure controllable and high-fidelity synthesis. The introduced image prior serves as conditional guidance to empower the real-domain customization capacity. Moreover, we design a sequential augmentation module to promote dynamics modeling, thus improving the temporal or stereoscopic coherence of generated data.
  • We introduce a noisy synthetic data filter to suppress harmful synthetic sequences at class semantics and sequential levels. By effectively reducing noise and enhancing the reliability of the synthetic databases, we can associate the synthetic samples with downstream tasks, thus better improving the classification performance.
  • Comprehensive experiments on 3 medical datasets, with 11 downstream networks trained on 3 paradigms, show that our proposed framework can improve diagnostic performance, especially within underrepresented high-risk sets. Ctrl-GenAug also increases out-domain robustness, further indicating its practicality in clinical scenarios.
It is noted that the sequence generator extends our prior work presented at MICCAI [23], with significant improvements in controllability at both semantic and sequential levels. We highlight that customized semantics and consistent sequential dynamics are crucial for enhancing the learning capability of downstream classifiers. Hence, we incorporate attribute texts and class labels to provide fine-grained semantic guidance, while employing a computation-efficient motion field to enhance dynamic control. Besides, we propose a sequential augmentation module to further ensure the sequential consistency and smoothness of generated samples.
II Related Works
In this section, we briefly review the approaches for controllable sequence synthesis using diffusion models. Given the strong inter-slice association in 3D medical volumes [27], we treat them as video sequence data in our study and summarize related video-based methods. Finally, we deeply involve existing diffusion-based generative augmentation schemes.
II-A Controllable Video Synthesis with Diffusion Models
Diffusion models, a class of generative models, have recently attracted significant attention, especially text-to-video (T2V) synthesis [28, 29, 10, 11]. However, these methods did not achieve precise control over the visual appearance and dynamics of the synthesized video, resulting in limited practicality. To mitigate this issue, researchers have concentrated on incorporating multimodal conditions into T2V frameworks to guide controllable synthesis. Specifically, AnimateDiff [30] and MoonShot [31] introduced text and image conditions to facilitate the control of visual appearance. Most recently, several studies [19, 32, 33] achieved fine-grained spatial compositional control by integrating additional diverse conditions (i.e., sketch, mask, and depth sequences). Meanwhile, two spatial-sequential (S2) modeling strategies (fully attention layers [19] and S2 condition encoder [33]) were introduced to capture the dynamics of sequential conditions, promoting the inter-frame features interaction. Besides, dense optical flow [34] was adopted to control sequential consistency [35, 36]. The above methods have shown promising controllability in natural video synthesis. However, for medical synthesis, acquiring abundant dense annotations (e.g., mask sequences) as control signals is impractical and unaffordable. Additionally, medical videos have unique attributes compared with natural ones, such as blurred anatomical regions and complex structure variations. Hence, they may not suit the medical video synthesis tasks.
In the medical field, several diffusion-based methods have been proposed to generate photorealistic echocardiography videos using the conditional guidance of a single frame [37], ejection fractions [37], single semantic mask [38, 39], and single sketch image [23]. Zhou et al. [23] additionally leveraged mitral valve (MV) skeletons to control the complex motion trajectories of MV in the generated echocardiograms. Besides, Li et al. [24] proposed an approach to generate endoscopy videos that simulate clinical scenes by integrating a spatial-temporal transformer with 2D vision foundation model priors. However, the above methods may have the following drawbacks: unsteerable semantics due to the lack of attribute guidance (e.g., category or target shape), poor control over sequences, etc. Thus, they may be unsuitable for direct adoption in the generative augmentation tasks.
II-B Generative Augmentation with Diffusion Models
Synthetic data can effectively supplement traditional data augmentation (e.g., rotation) and provide diverse samples to help deep model training [7, 40]. Previous works have succeeded in utilizing synthetic data generated with diffusion models to enhance natural image classification [26, 41, 42, 43], segmentation [44, 45, 46], and detection [47]. Recently, Singh et al. [48] performed an in-depth analysis of models trained with synthetic data across various robustness measures, and verified that they achieved good suitability in real-world settings.
To address data scarcity and promote diagnostic performance, the intelligent medical field has raised increasing interest in exploring diffusion-based generative augmentation for aiding different downstream diagnostic tasks [49, 50, 25, 51, 52, 53]. Luo et al. [7] marked the first focus on comprehensively analyzing the impact of their designed uncertainty-guided diffusion models on downstream diagnostic tasks. Nevertheless, current studies have two main limitations:
On the one hand, due to the lack of adequate semantic and sequential guidance during sampling, coupled with the absence of post-sampling quality control for generated samples, existing works could hardly fully ensure the reliability of synthetic databases for downstream disease recognition. For instance, it is desirable to embed descriptive text (e.g., lesion morphology) to provide more informative guidance associated with disease grading. Unlike previous solutions, we aim to develop a generative augmentation framework that enables highly semantic- and sequential-customized sequence synthesis and suppresses incorrectly synthesized samples to better facilitate downstream classifier learning.
On the other hand, few studies comprehensively explore the effect of generative models on downstream medical tasks, and they solely focus on image-level generation. Considering the value of dynamic information for clinical diagnosis, we concentrate on diagnosis-reliable sequence synthesis. This is more challenging than image-related works as synthetic data not only meets high-fidelity requirements but also ensures temporal or stereoscopic consistency and coherence for effectively promoting downstream S2 modeling capacity. A brief review of studies in the medical field utilizing diffusion models to promote downstream tasks can be found in the Supplementary Material.
III Method
Figure 2: Pipeline of using our proposed framework to facilitate medical sequence recognition, which can be worked with a variety of classifiers. Here, we use the carotid plaque US video sequence as an example to demonstrate the overall process.
III-A Model Overview
Fig. 2 shows the pipeline of using our proposed controllable generative augmentation framework to aid medical sequence classification. First, we propose a sequence generator that enables perceiving multiple semantic and sequential conditions to guide controllable and diverse generation. Then, an efficient noisy data filter is introduced to suppress unsatisfactory synthetic sequences. Last, the quality-controlled synthetic sequences and real ones will work together to improve the performance of arbitrary classifiers. To formulate customized, high-fidelity, and coherent synthetic databases to boost classification, we design the whole framework from model-centric, i.e., sequence generator (Secs. III-BIII-C, and III-D) and data-centric, i.e., data filter (Sec. III-E) perspectives.
III-B Basic Architecture of Sequence Generator
Fig. 3 shows the pipeline of our proposed sequence generator. It supports customized and high-quality medical sequence generation via multimodal conditions. The generator is implemented with a two-stage training scheme. In the pretraining stage, it attends to high-fidelity visual features learning for controllable image synthesis. While in the finetuning stage, the domain-specific visual knowledge acquired from the previous stage is reused and focuses on sequential patterns modeling for customized sequence synthesis. During inference, given a single or combination of multimodal conditions as control signal inputs, high-quality and steerable sequence generation from Gaussian noise can work. We then provide a short background of video diffusion models in Sec. III-B1, followed by a detailed description of the basic architecture of our generator in Secs. III-B2 and III-B3.
III-B1 Preliminaries of Video DDPMs
They are trained to learn complex video distribution by iteratively denoising corrupted inputs using pixel-space diffusion methods [54, 55]. Given a training video $\mathit{x}_{0}\sim\mathit{q}$ and let $\mathit{T}$ the timestep, DDPMs first produce a sequence of noisy inputs via a diffusion process $\mathit{q}(\mathit{x}_{t}|\mathit{x}_{0},\mathit{t}),\mathit{t}\in{1,2,...,\mathit{T}}$, which progressively adds Gaussian noise $\epsilon\in\mathcal{N}(\mathbf{0},\mathbf{I})$ to $\mathit{x}_{0}$. To ease the computational burden due to pixel-space training, video latent diffusion models [28] perform the diffusion process in latent space of a variational autoencoder [56] (VAE). The model is then trained to estimate the parameterized Gaussian transition $\mathit{p}(\mathit{x}_{t-1}|\mathit{x}_{t})$ via a denoising network $\theta$. Mathematically, the optimized objective can be a simplified variant of the variational lower bound:
$\begin{split}\min_{\theta}\mathbb{E}_{\mathit{z}_{0},\epsilon\sim\mathcal{N}(\mathbf{0},\mathbf{I}),\mathbf{c},\mathit{t}}||\epsilon-\epsilon_{\theta}(\mathit{z}_{t},\mathbf{c},\mathit{t})||^{2},\end{split}$ (1)
where $\mathit{z}_{0}$ is the latent code of $\mathit{x}_{0}$. $\epsilon_{\theta}$ and $\epsilon$ represent the predicted and target noise, respectively. Here, $\mathbf{c}$ denotes the (optional) control signal that the model can be conditioned on. For our generator, $\mathbf{c}$ involves multimodal conditions including semantic and sequential ones.
III-B2 Factorized Learning of Visual Features and Sequential Patterns
Developing a generator with pure 3D architecture design is an intuitive scheme for sequence synthesis [28]. However, it is very challenging to simultaneously learn visual features and sequential patterns due to the scarcity and complexity of medical data, while causing high training costs. To this end, we propose to factorize both learning by first pretraining a latent diffusion model (LDM) and then finetuning its sequence counterpart. In this way, the generator enables realistic sequence synthesis, while easing the training burden.
III-B3 2D-to-3D Model Inflation
In the pretraining phase, we develop a 2D UNet [57] in LDM to predict noises for image synthesis. Referring to [13], we extend the 2D UNet to a 3D version via an inflation scheme to build the sequence LDM. Specifically, all the spatial convolution layers are inflated to pseudo-3D counterparts by expanding the kernels at the sequential dimension (e.g., $3\times 3\rightarrow 1\times 3\times 3$ kernel). Besides, we perform sequential insertion by adding sequential attention (SA) layers (see Fig. 3). Fig. 4(a) shows the principle of the SA mechanism, where each patch queries to those at the same spatial position and across frames/slices. This design encourages the generator to model the sequential patterns, while not significantly altering visual feature distribution baked in the LDM [31]. Overall, sequence LDM enables inheriting the rich visual concepts preserved in LDM and focusing on sequential pattern aggregation, making the model learning efficient.
Figure 3: Pipeline of our proposed sequence generator. The MedSAM image encoder [58] is adopted for domain-specific image prior feature extraction.
III-C Multimodal Conditions Guidance
Sequence generators guided by a single condition (e.g., text) have the following drawbacks: 1) poor controllability and 2) sequential inconsistency. To solve these issues, we introduce multimodal conditions to guide our sequence generator for customized, realistic, and sequential-consistent synthesis.
We consider four multimodal conditions to ensure comprehensive and accurate control over the sequence synthesis procedure (see Figs. 2-3). Specifically, these conditions are divided into semantic and sequential ones for visual appearance control and serial guidance, respectively. We highlight that our generator supports composable synthesis by allowing users to flexibly choose any single condition or composition during inference. This flexibility makes our generator particularly user-friendly, as it enables high-quality synthesis even when certain conditions are missing. Details are described below.
III-C1 Semantic Conditions for Visual Appearance Control
As shown in Fig. 3, the sequence generator exploits three semantic conditions to perform visual appearance control, thus achieving controllable and high-fidelity synthesis.
- Descriptive Text: It provides an intuitive indication of the coarse-grained semantic concepts of sequences. In our implementation, we pre-align the visual and textual features by finetuning the widely used CLIP model [59] using our medical datasets to mitigate the large gap between them. Then, semantic information of texts can feasibly query relevant medical visual patterns, thus easing subsequent training. Similar to [8], the text guides the generator via cross-attention layers.
- Class Label: Generating medical data with the expected disease class is crucial for enhancing downstream classification, particularly improving the diagnostic accuracy of underrepresented high-risk sets. Thus, we propose encouraging the model to focus on the grading signal. Specifically, rather than merging the class label with descriptive text, we treat it as a separate tag to directly provide disease-specific guidance, as merging with text may dilute the class signal. We further propose a dedicated insertion method for this purpose. First, similar to the timestep embedding process [8] in diffusion models, we adopt a class label encoder composed of a label discretization layer and an embedding layer to obtain its embeddings. The addition between the timestep and class label embeddings is then input to the spatial convolution layer in each UNet block for visual appearance control.
- Image Prior: Merely using the above conditions faces challenges of insufficient semantic control and an inevitable domain gap between real and synthetic sequences, constraining the capability of generative augmentation. Hence, we introduce the first frame/slice of real-domain sequences as image priors to provide rich semantic guidance and yield real-domain style sequences. As image prior offers global semantics like texts, we align image prior features $f_{i}$ extracted from image prior encoder (see Fig. 3) with text embeddings $f_{t}$ generated by our medical data-specific CLIP text encoder for joint guidance. Specifically, motivated by [60], we replace the text-guided cross-attention layers in original UNet blocks with decoupled counterparts that handle texts and image priors in parallel and then merge the results by addition. The decoupled attention can be formulated as:
[…]
where $W^{Q},W^{K},W^{V}$ denote the trainable linear projection matrices. $\mathbf{t},\mathbf{i}$ represent text and image prior conditions. $\mathit{d}$ denotes dimension of latent features. $\sigma(\cdot)$, softmax function.
III-C2 Sequential Condition for Serial Guidance
Apart from perceiving visual concepts, modeling sequential knowledge is also important in sequence synthesis, with sequential cues playing a vital role in this process. Most existing methods [35, 36] adopted dense optical flow [34] to promote sequential dynamics modeling. However, optical flow extraction requires high computational demands [33]. Thus, we instead introduce computation-efficient motion field as a sequential condition in sequence LDM. It explicitly showcases the pixel-wise motions between adjacent frames/slices (see Figs. 2-3). In our approach, we first extract the motion fields of real sequences using a Python package [61]. Then, a motion encoder (see Fig. 3) receives the motion fields and produces motion features. Last, the features are concatenated with latent representations $\mathit{z}$ from VAE along the channel dimension for serial guidance.
Figure 4: Schematics of three attention mechanisms for sequential modeling (a-c) and our proposed sequential augmentation module in this study (d).
III-D Sequential Augmentation Module
Solely equipping with SA layers and sequential cues in sequence LDM may present inadequate consistency and coherence across synthetic frames/slices. This issue may arise from insufficient sequential modeling of input noisy latents and motion fields, constrained by a small parameter space, overly burdening the SA layers. To solve the issue, we propose a sequential augmentation module (SAM) that enables the generator to more effectively model sequential dependencies. As shown in Fig. 4(d), SAM integrates two attention mechanisms in cascade for sequential augmentation.
III-D1 Key-frame/slice Attention
The common spatial self-attention layers can lead to sequential inconsistency due to the lack of interaction across frames/slices. To augment sequential consistency, we introduce a key-frame/slice attention (KA) mechanism, where two selected key-frames/slices act as references to propagate S2 information throughout the sequence. Specifically, for any frame/slice, we select its previous and the first counterpart of the sequence as the references and transform spatial self-attention into KA, aiming to align the latent features $z_{l}$ of the l-th frame/slice with $z_{1}$ and $z_{l-1}$. We obtain query from $z_{l}$, key and value features from $z_{1}$ and $z_{l-1}$, and compute $Attention\left(\mathbf{Q},\mathbf{K},\mathbf{V}\right)$ using:
$\begin{split}\mathbf{Q}=W^{Q}z_{l},\mathbf{K}=W^{K}\left\{z_{1},z_{l-1}\right\},\mathbf{V}=W^{V}\left\{z_{1},z_{l-1}\right\},\end{split}$ (3)
[…]
III-D2 Motion Field Attention
To further boost the sequential coherence of the generated sequences, we propose to reuse the motion field condition (refer to Sec. III-C2) by introducing a motion field attention (MFA) mechanism after KA. As shown in Fig. 4(c), MFA requires the patches to communicate with those in the same motion field-based pathway including itself, which eliminates flickers of the generated sequences to make the contents visually smooth. Inspired by [62], MFA is implemented in two steps: a) motion field-based patch pathway sampling and b) attention calculation. In step 1, we sample the patch pathways based on the m-scaled downsampled motion fields. Take a video sequence as an example, for a patch $p_{l}$ on the l-th frame of a f-frame video, the $path$ can be derived from the motion field. Since the sampling procedure inevitably generates multiple pathways for the same patch, we randomly sample a pathway to ensure its uniqueness to which each patch belongs. In this setting, let $H,W$ the height and width of the input video frame, then the size of the pathway set after sampling equals $\frac{H\times W}{m^{2}}\times f$. In step 2, we calculate $Attention(\mathbf{Q},\mathbf{K},\mathbf{V})$ with:
$\begin{split}\mathbf{Q}=z_{p_{l}},\mathbf{K}=\mathbf{V}=z_{path},\end{split}$ (4)
where $z_{path}$ represents the latent features of patches on the $path$, as follows:
$\begin{split}z_{path}=\left[z_{p_{0}},z_{p_{1}},\dots,z_{p_{l}},\dots,z_{p_{f-1}},z_{p_{f}}\right].\end{split}$ (5)
The output of KA is directly fed into MFA without being handled by the query, key, and value projection functions. Hence, MFA can comfortably decrease extra computation.
Algorithm 1 Noisy synthetic data filtering.
1:synthetic samples set $\mathrm{S}$, real data-trained classifier $p$.
2:# Stage 1: Class semantics misalignment filtering
3:$\mathrm{S_{1}}=\varnothing$triangleright synthetic samples set obtained after stage 1
4:for $i\in\left\{1,...,n\right\}$ do
5:    $c\leftarrow$ class in the conditions bank of the group $\mathrm{S}[i]$
6:    $l_{x}=-\sum c\mathrm{log}\sigma(p(\mathrm{S}[i][x]))$triangleright loss of the sample indexed $x$
7:    $L_{c}=\frac{1}{M}\sum_{x=1}^{M}l_{x}$triangleright noise identification threshold
8$\triangleright$ $M$ clips in the group
9:    for $x\in\left\{1,...,M\right\}$ do
10:         if $l_{x}>L_{c}$ then
11:             $continue$
12:         else
13:             $\mathrm{S_{1}}\leftarrow\mathrm{S_{1}}\cup\{\mathrm{S}[i][x]\}$triangleright save less noisy sample
14:         end if
15:    end for
16:end for
17:return $\mathrm{S_{1}}$triangleright with $n_{1}(<n)$ groups totaling $N_{1}(<N)$ clips
18:# Stage 2: Inner-sequence filtering
19:$\mathrm{S_{2}}=\varnothing$triangleright synthetic samples set obtained after stage 2
20:$\mathrm{A}=\left\{A_{1},...,A_{N}\right\}$triangleright VAE-Seq values set of $\mathrm{S}$
21:$\mathrm{B}=\left\{B_{1},...,B_{N_{1}}\right\}$triangleright VAE-Seq values set of $\mathrm{S_{1}}$
22:$t_{l},t_{h}\leftarrow\mathbf{KMeans}(\mathrm{A},\mathbf{K}=4)$triangleright boundary thresholds
23:for $j\in\left\{1,...,N_{1}\right\}$ do
24:    if $B_{j}\geq t_{l}\And B_{j}\leq t_{h}$ then
25:         $\mathrm{S_{2}}\leftarrow\mathrm{S_{2}}\cup\{y_{j}|y_{j}\in\mathrm{S_{1}}\}$triangleright save gentle dynamic sample
26:    else
27:         $continue$
28:    end if
29:end for
30:return $\mathrm{S_{2}}$triangleright with $n_{2}(<n_{1})$ groups totaling $N_{2}(<N_{1})$ clips
31:# Stage 3: Inter-sequence filtering
32:$\mathrm{S_{3}}=\varnothing$triangleright synthetic samples set obtained after stage 3
33:for $k\in\left\{1,...,n_{2}\right\}$ do
34:    $\mathrm{S_{3}}\leftarrow\mathrm{S_{3}}\cup\mathrm{S_{2}}[k][1]$
35:    for $q\in\left\{2,...,M^{\prime}\right\}$ do $\triangleright$ $M^{\prime}$ clips in the group $\mathrm{S_{2}}[k]$
[…]
37:             $\mathrm{S_{3}}\leftarrow\mathrm{S_{3}}\cup\mathrm{S_{2}}[k][q]$triangleright save diverse sample
38:         else
39:             $continue$
40:         end if
41:    end for
42:end for
43:return $\mathrm{S_{3}}$triangleright final synthetic database for downstream training
III-E Noisy Synthetic Data Filter
We use the proposed sequence generator to constitute the synthetic samples set. Concretely, assuming there are $n$ training clips in the target dataset, with a bank of conditions derived from each clip, we synthesize a group of clips guided by each conditions bank. Eventually, we obtain $n$ groups of clips to form our synthetic sample set, with a total of $N$ clips. Although visually realistic and smooth (good cases, Fig. 5), sequence synthesis may still suffer from class semantics misalignment, cross-frame/slice inconsistency or over-consistency (i.e., almost static clip), and inter-clip similarity. For instance, in Fig. 5(e), the synthesized carotid clip category is wrong, which should be moderate rather than mild. In Fig. 5(f), the generated clip includes abrupt changes in anatomical structures. Hence, blindly using all synthetic clips for classifier learning will significantly cause a performance drop due to noisy samples. Our work proposes a noisy synthetic data filter to adaptively remove harmful generated clips at class semantics and sequential levels, as illustrated in Algorithm 1. This can also link the synthetic data to downstream tasks, thus potentially achieving a higher performance upper bound.
III-E1 Class Semantics Misalignment Filtering
This scheme is designed to filter the generated clips whose visual contents mismatch the corresponding class labels. Motivated by anomaly detection [63], we consider that such noisy synthetic samples will produce large losses if fed into a powerful classifier trained on real ones. Hence, a loss threshold can be set to identify and filter the anomalistic clips. Specifically, we first adopt a well-trained classifier in the real data domain to calculate case-wise cross-entropy loss values $\left\{l_{x}\right\}_{x=1}^{M}$ based on corresponding category $c$ for all synthetic clips in each group. Then, we set the average loss as the noise identification threshold of $c$. Last, we conduct class semantics misalignment filtering at the group level. If the loss of a synthetic clip with $c$ exceeds the threshold $L_{c}$, it will be viewed as a noisy sample and excluded from the synthetic pool.
III-E2 Sequential Filtering
This strategy aims to filter noisy samples at a) inner-sequence and b) inter-sequence levels. For a), we screen out synthetic clips with gentle dynamic, avoiding those that are either inconsistent or over-consistent affecting downstream learning. Specifically, we retain a synthetic clip whose cross-frame/slice consistency falls in a pre-computed range. It is determined by K-means clustering [64] based on the cross-frame/slice consistency of all synthetic clips. To assess the video sequence consistency, [33] tends to use CLIP [59] image embeddings to compute the average cosine similarity across consecutive frames. However, this method leads to inadequate evaluation due to the usage of limited informative CLIP embeddings. To resolve the problem, we propose a metric termed VAE-Seq that utilizes latent embeddings from pretrained VAE, instead of deriving embeddings from CLIP space, to assess cross-frame/slice consistency. Therefore, with the higher dimension of latents (e.g., 4$\times$32$\times$32 for input size 256$\times$256 compared to 768 in CLIP embeddings), VAE-Seq reflects an accurate evaluation using more fine-grained and informative features. For b), we seek to diversify the samples in the synthetic set to prevent overfitting and avoid wasting computational resources in downstream training. To this end, we perform the inter-sequence filtering based on inter-clip similarity in each group, as shown in Algorithm 1. We quantify the inter-clip similarity by calculating the cosine distance of VAE latent embeddings between frame/slice pairs in two clips and averaging the obtained distance values.
III-F Evaluation
As a common practice, existing sequence synthesis works evaluate the synthetic quality by mostly using Fréchet Inception Distance (FID) [65] and Fréchet Video Distance (FVD) [66]. However, researchers have verified that these metrics do not consistently correlate with performance metrics on downstream tasks [46, 7]. To exhaustively evaluate the values of synthetic samples to downstream tasks, eleven popular classifiers are deployed for comparison with respect to diagnostic accuracy and area under the receiver operating characteristic curve (AUROC). Besides, we adopt the proposed VAE-Seq to assess the cross-frame/slice consistency (refer to Sec. III-E2). Moreover, we consider it crucial to evaluate the smoothness of generated clips. Motivated by [67], we quantify it using the proposed metric (termed Dynamic Smoothness), which is the mean absolute error between the reconstructed frames/slices from the interpolation model [68] and the original real ones.
Figure 5: Typical good and bad synthetic samples of three datasets using different banks of conditions. Conditional controls include class label (C), text (T), image prior (I), and motion field (MF). See more synthetic results in Supplementary Material.
IV Experiments
In this section, we first comprehensively validated the effectiveness and generality of our method on three medical datasets, using eleven popular networks trained on three paradigms. We then tested the performance of our method in underrepresented high-risk sets and out-domain conditions. Last, we analyzed the contributions of our proposed multimodal conditions guidance, sequential augmentation module, and noisy synthetic data filter respectively, as well as the impacts of traditional and generative augmentations.
TABLE I: Experimental datasets and settings.
Carotid TUSC [69] ACDC [70]
Patient 231 167 150
Clip 486 633 314
Mild/Moderate/Severe TI-RADS2-3/4/5 NOR/MINF/DCM/HCM/ARV
Class (train) 141/104/64 146/189/113 40/40/40/40/45
Class (test) 51/27/20 64/73/48 21/20/20/20/28
Class (OD) 37/24/18 - -
Sequence generator Training
Downsampling rate of VAE(8)   $\mathit{T}$(1,000)
AdamW(lr=1e-4)   lr scheduler(Cosine)   Warm(500 iterations)
Epochs(200)   Pretraining batch size(64)   Finetuning batch size(8)
Inference
Classifier-free guidance [71] factor(7.5)   Sampling steps(200)
Eleven classifiers Adam(lr=1e-4)   lr scheduler(Cosine)   Epochs(100)   Batch size(8)
Traditional augmentations(Random Color/Move/Gaussian/Rotation/Flip)
IV-A Datasets and Settings
Carotid dataset was collected by three medical centers with approval from local institutional review boards, including a) The Third Affiliated Hospital of Sun Yat-sen University, b) Shenzhen Longgang District People’s Hospital, and c) The Eighth Affiliated Hospital of Sun Yat-sen University. It consists of 273 US video sequences from 231 patients with carotid plaque-induced stenosis. The severity of carotid stenosis is graded into three stages, i.e., mild, moderate, and severe. We uniformly sampled several non-overlapping 8-frame clips from each sequence. The final dataset involved an in-domain (ID) subset (407 clips from 193 patients collected by Center (a)) and an out-domain (OD) subset (79 clips from 38 patients acquired from Center (b)-(c)). The ID/OD subset was randomly split into 309/39 and 98/40 clips for training and testing at the patient level. Each clip has up to two plaques and was labeled with stenosis grading and descriptive text by experienced sonographers using Pair annotation software package [72]. The text annotations indicated the plaque characteristics of echogenicity and location (see Fig. 2). We cropped the region of interest based on the pretrained carotid vessel detector [73] for easing the model learning.
TUSC [69] contains 192 thyroid US video sequences from 167 patients collected by the Stanford University Medical Center, with each including one nodule. Each sequence provided a TI-RADS level and nodule descriptors [74] (e.g., shape), the latter of which was used to form descriptive text. The TI-RADS level is ranging from 1 to 5. Considering the imbalanced distribution of TI-RADS levels, as suggested by the experienced sonographers, we rebuilt the dataset by removing the TI-RADS1 video sequence and combining the levels 2-3 sequences into a unified category named TI-RADS2-3. We uniformly sampled the sequences, obtaining 633 8-frame clips, with 448 for training and 185 for testing. Each sampled clip inherited the same TI-RADS level and text as the original sequence. The bounding box of the nodule region for each frame was extracted by radiologists, similar to [75].
ACDC [70] includes 4D cardiac cine-MRI from 150 patients acquired at the University Hospital of Dijon. The patients were evenly divided into five classes: normal patients (NOR), myocardial infarction (MINF), dilated cardiomyopathy (DCM), hypertrophic cardiomyopathy (HCM), and abnormal right ventricle (ARV). We first extracted 3D volume sequences at end-systole and diastole phases from each case and assigned corresponding classes. Then, the 8-slice clips were uniformly sampled from each sequence along the slice axis. Finally, 205 and 109 clips were constructed for training and testing. Text prompts were set to empty strings during generator training.
Table I presents a detailed description of experimental datasets and settings. Using all available conditions, we synthesized 300 clips per category for the ACDC dataset and 500 clips per class for others. All frames/slices were resized to 256$\times$256. By default, we set m=8 for the MFA mechanism. For the sequence generator, the pretraining LDM was built upon Stable Diffusion [8] and initialized with the public weights [^2] https://huggingface.co/CompVis/stable-diffusion-v1-4. During sequence LDM finetuning, we updated the parameters of SAM, SA layers, and motion encoder. We further refined the text- and image prior-sequence alignment by keeping the query projection matrices in cross-attention layers trainable. During generator training, we employed the multimodal conditions joint training strategy [76]. This allows users to flexibly select any single condition or combination during sampling, without finetuning the model for each specific combination. Motivated by [46], we adopted three paradigms to explore the impacts of the synthetic samples for downstream classifiers. Baseline: merely trained with real clips. Real-finetune: initially pretrained with synthetic clips and then finetuned with real ones. Joint-train: real clips were first over-sampled to match the number of synthetic clips, the classifier was then trained on the real-synthetic mixed set. All classification experiments were conducted on MMAction2 [77] except for FTC [78] and CSG-3DCT [1].
[…]
Method Backbone Training Paradigm Carotid TUSC [69] ACDC [70]
Acc. $\Delta$ AUROC $\Delta$ Acc. $\Delta$ AUROC $\Delta$ Acc. $\Delta$ AUROC $\Delta$
I3D [79] CNN A 79.59 0.737 60.72 0.559 83.49 0.733
B 85.03 $\uparrow$ 5.44 0.813 $\uparrow$ 0.076 65.41 $\uparrow$ 4.69 0.611 $\uparrow$ 0.052 87.16 $\uparrow$ 3.67 0.795 $\uparrow$ 0.062
C 84.35 $\uparrow$ 4.76 0.811 $\uparrow$ 0.074 62.88 $\uparrow$ 2.16 0.578 $\uparrow$ 0.019 85.32 $\uparrow$ 1.83 0.762 $\uparrow$ 0.029
R(2+1)D [80] CNN A 82.99 0.781 63.96 0.595 83.85 0.737
B 83.67 $\uparrow$ 0.68 0.797 $\uparrow$ 0.016 64.68 $\uparrow$ 0.72 0.601 $\uparrow$ 0.006 85.69 $\uparrow$ 1.84 0.769 $\uparrow$ 0.032
C 85.04 $\uparrow$ 2.05 0.804 $\uparrow$ 0.023 65.41 $\uparrow$ 1.45 0.610 $\uparrow$ 0.015 85.69 $\uparrow$ 1.84 0.771 $\uparrow$ 0.034
SlowFast [81] CNN A 80.95 0.742 62.88 0.579 81.65 0.702
B 81.63 $\uparrow$ 0.68 0.770 $\uparrow$ 0.028 65.76 $\uparrow$ 2.88 0.609 $\uparrow$ 0.030 86.05 $\uparrow$ 4.40 0.772 $\uparrow$ 0.070
C 84.35 $\uparrow$ 3.40 0.811 $\uparrow$ 0.069 65.41 $\uparrow$ 2.53 0.609 $\uparrow$ 0.030 83.85 $\uparrow$ 2.20 0.742 $\uparrow$ 0.040
CSN [82] CNN A 80.27 0.757 61.80 0.570 85.69 0.765
B 84.35 $\uparrow$ 4.08 0.784 $\uparrow$ 0.027 67.21 $\uparrow$ 5.41 0.631 $\uparrow$ 0.061 87.15 $\uparrow$ 1.46 0.794 $\uparrow$ 0.029
C 87.07 $\uparrow$ 6.80 0.843 $\uparrow$ 0.086 65.41 $\uparrow$ 3.61 0.610 $\uparrow$ 0.040 88.62 $\uparrow$ 2.93 0.814 $\uparrow$ 0.049
TPN [83] CNN A 82.99 0.784 64.32 0.602 83.85 0.735
B 87.76 $\uparrow$ 4.77 0.856 $\uparrow$ 0.072 66.85 $\uparrow$ 2.53 0.633 $\uparrow$ 0.031 85.32 $\uparrow$ 1.47 0.764 $\uparrow$ 0.029
C 85.71 $\uparrow$ 2.72 0.823 $\uparrow$ 0.039 65.41 $\uparrow$ 1.09 0.609 $\uparrow$ 0.007 84.95 $\uparrow$ 1.10 0.756 $\uparrow$ 0.021
TimeSformer [84] Transformer A 77.55 0.721 65.41 0.614 72.48 0.562
B 80.95 $\uparrow$ 3.40 0.753 $\uparrow$ 0.032 66.49 $\uparrow$ 1.08 0.626 $\uparrow$ 0.012 76.51 $\uparrow$ 4.03 0.626 $\uparrow$ 0.064
C 79.59 $\uparrow$ 2.04 0.730 $\uparrow$ 0.009 66.49 $\uparrow$ 1.08 0.618 $\uparrow$ 0.004 77.62 $\uparrow$ 5.14 0.642 $\uparrow$ 0.080
MViTv2 [85] Transformer A 74.15 0.663 59.64 0.543 77.98 0.642
B 78.91 $\uparrow$ 4.76 0.740 $\uparrow$ 0.077 63.96 $\uparrow$ 4.32 0.596 $\uparrow$ 0.053 78.72 $\uparrow$ 0.74 0.653 $\uparrow$ 0.011
C 76.87 $\uparrow$ 2.72 0.701 $\uparrow$ 0.038 62.17 $\uparrow$ 2.53 0.568 $\uparrow$ 0.025 80.18 $\uparrow$ 2.20 0.680 $\uparrow$ 0.038
VideoSwin [86] Transformer A 79.59 0.732 58.92 0.535 74.31 0.572
B 83.67 $\uparrow$ 4.08 0.789 $\uparrow$ 0.057 61.08 $\uparrow$ 2.16 0.551 $\uparrow$ 0.016 79.45 $\uparrow$ 5.14 0.663 $\uparrow$ 0.091
C 83.67 $\uparrow$ 4.08 0.776 $\uparrow$ 0.044 61.44 $\uparrow$ 2.52 0.562 $\uparrow$ 0.027 77.62 $\uparrow$ 3.31 0.632 $\uparrow$ 0.060
UniFormerV2 [87] Hybrid A 76.87 0.710 65.41 0.611 75.78 0.614
B 80.27 $\uparrow$ 3.40 0.743 $\uparrow$ 0.033 65.77 $\uparrow$ 0.36 0.618 $\uparrow$ 0.007 77.25 $\uparrow$ 1.47 0.645 $\uparrow$ 0.031
C 81.63 $\uparrow$ 4.76 0.760 $\uparrow$ 0.050 66.13 $\uparrow$ 0.72 0.617 $\uparrow$ 0.006 77.61 $\uparrow$ 1.83 0.637 $\uparrow$ 0.023
FTC [78] Hybrid A 76.99 0.723 62.09 0.584 80.95 0.689
B 80.91 $\uparrow$ 3.92 0.750 $\uparrow$ 0.027 64.09 $\uparrow$ 2.00 0.593 $\uparrow$ 0.009 82.38 $\uparrow$ 1.43 0.703 $\uparrow$ 0.014
C 81.80 $\uparrow$ 4.81 0.758 $\uparrow$ 0.035 65.66 $\uparrow$ 3.57 0.614 $\uparrow$ 0.030 84.38 $\uparrow$ 3.43 0.718 $\uparrow$ 0.029
CSG-3DCT [1] Hybrid A 84.35 0.812 61.44 0.571 87.16 0.797
B 87.07 $\uparrow$ 2.72 0.831 $\uparrow$ 0.019 63.96 $\uparrow$ 2.52 0.590 $\uparrow$ 0.019 88.62 $\uparrow$ 1.46 0.819 $\uparrow$ 0.022
C 87.76 $\uparrow$ 3.41 0.837 $\uparrow$ 0.025 65.05 $\uparrow$ 3.61 0.609 $\uparrow$ 0.038 88.63 $\uparrow$ 1.47 0.820 $\uparrow$ 0.023
IV-B Generative Augmentation-based Sequence Classification
We utilized Real-finetune and Joint-train paradigms to integrate real and synthetic clips for downstream multi-class sequence classification training. The efficacy of our framework was fully validated across three medical datasets using eleven common classifiers. To demonstrate the generalizability of Ctrl-GenAug, the selected networks involve 3D CNN, transformer, and CNN-Transformer hybrid designs, which specialize in capturing local, global, and mixed features, respectively.
Table II compares the average performance (i.e., accuracy and AUROC) on different training paradigms across a wide range of sequence classifiers, including five CNN-based classic models, three transformer-based models, and three CNN-Transformer hybrid-based methods. It is evident that the performance ceiling of existing popular models, from simple CNN-based to complex CNN-Transformer hybrids, can be further elevated using synthetic samples obtained from our proposed framework, under Real-finetune and Joint-train paradigms. On the carotid dataset, the accuracy/AUROC can be most boosted by 5.44%/0.077 and 6.80%/0.086 through Real-finetune and Joint-train paradigms, respectively. Due to high intra-class variations in TUSC, the baseline exhibits visibly inferior results, with an average accuracy of 62.42% compared to other datasets. However, it can be seen from Table II that samples synthesized by Ctrl-GenAug make significant contributions and enhance the classification performance to an acceptance level. Among all the classifiers, CSN harvests the largest performance gain, which is non-trivially enhanced by 5.41%/0.061 and 3.61%/0.040 accuracy/AUROC under Real-finetune and Joint-train paradigms, respectively. Notably, for ACDC, the overall performance improvement remains significant even with a high baseline. For instance, CSG-3DCT is boosted from 87.16$\rightarrow$88.63(+1.47%) accuracy and from 0.797$\rightarrow$0.820 (+0.023) AUROC under the Joint-train practice. Extensive experimental results show that synthetic sample sets produced by Ctrl-GenAug can be powerful data engines to enhance various diagnostic tasks in medical sequence analysis.
IV-C Improved Diagnosis in Underrepresented High-risk Sets
We tested the effectiveness of Ctrl-GenAug by evaluating the diagnostic performance in underrepresented high-risk situations using our carotid dataset. In our study, underrepresented represents the tail data with small amounts (i.e., moderate&severe), and overrepresented defines mild data with sufficient amounts. To build a class-biased dataset, we skewed the real training dataset by randomly sampling 25% clips from each high-risk class (moderate&severe) in the underrepresented sets. We leveraged Ctrl-GenAug to expand the clips of each high-risk class to match the number of the low-risk class (mild), resulting in an augmented dataset. Four downstream networks were trained on the augmented dataset with Joint-train paradigm, and six metrics were used for class-level evaluation, including sensitivity, specificity, accuracy, F1-score, precision, and AUROC.
Fig. 6 quantitatively compares the carotid stenosis diagnostic performance of underrepresented high-risk sets on Baseline and Joint-train paradigms. It shows that Baseline performs poorly on underrepresented high-risk sets, especially in terms of sensitivity (average 18.52% in moderate and 36.25% in severe sets). In comparison, the classifiers can be greatly enhanced by jointly training with synthetic samples (e.g., SlowFast: 11.11$\rightarrow$44.44sensitivity on the moderate set).
We also provided the category-level analysis, and summarized the average F1 scores across four classifiers for a global observation and conclusion. Specifically, Baseline reached 69.92%, 25.92%, and 43.44% average F1 scores for mild, moderate, and severe cases, respectively. The F1 performance of Joint-train in these three categories was 82.44%, 52.73%, and 60.40%. We found that well-designed classifiers jointly trained on synthetic samples from Ctrl-GenAug and real ones, can effectively close diagnostic performance gaps between overrepresented (i.e., mild) and underrepresented sets (i.e., moderate and severe) while improving the former ($\sim$13$\uparrow$).
Figure 6: Carotid performance radar map, using 4 models trained on Baseline (green) and Joint-train (blue) paradigms in underrepresented high-risk sets.
IV-D Classification Robustness in Out-domain Conditions
We further explored the impact of synthetic samples from Ctrl-GenAug on diagnostic robustness in out-domain settings using the carotid dataset. Training data of the sequence generator consisted of labeled ID and unlabeled OD subsets. We added the hospital identifier in the descriptive text of the ID data to condition the domain distribution. For the OD data, which did not contain the diagnostic class label, we padded the corresponding conditioning vector with zeros, while solely preserving hospital IDs to form descriptive texts. As shown in Table III, the average accuracy and AUROC of three classifiers display better results on the Real-finetune and Joint-train paradigms compared to the Baseline. It validates that diagnostic robustness can be enhanced with the aid of Ctrl-GenAug in scenarios where we only have access to unlabeled cases from additional medical centers due to limited resources.
[…]
Method Backbone Training Paradigm Carotid
Acc. $\Delta$ AUROC $\Delta$
I3D [79] CNN A 70.00 0.609
B 71.67 $\uparrow$ 1.67 0.648 $\uparrow$ 0.039
C 75.00 $\uparrow$ 5.00 0.682 $\uparrow$ 0.073
TPN [83] CNN A 75.00 0.667
B 76.67 $\uparrow$ 1.67 0.731 $\uparrow$ 0.064
C 76.67 $\uparrow$ 1.67 0.678 $\uparrow$ 0.011
CSG-3DCT [1] Hybrid A 68.33 0.599
B 73.33 $\uparrow$ 5.00 0.689 $\uparrow$ 0.090
C 70.00 $\uparrow$ 1.67 0.651 $\uparrow$ 0.052
IV-E Analysis of Multimodal Conditions Guidance
To illustrate the role of different conditional guidance, we conduct comparative experiments with generators that resort to various banks of conditions for training and sampling. Table IV validates that all proposed conditions are effective for enhancing synthesis and diagnosis tasks. Specifically, without text control, the downstream diagnostic performance is comparatively poor. Fig. 5(b-c) compares typical synthetic thyroid nodule sequences generated under condition banks with and without text guidance. The former is visibly high-fidelity and faithful to the given text prompt (e.g., smooth margin), while the latter exhibits less distinguishable features for diagnosis (e.g., blurry margin). That is to say, the proposed Ctrl-GenAug can create more diagnosis-reliable samples by enhancing semantic steerability in the generation process. Besides, by incorporating image prior knowledge in the sampling process, the average accuracy and AUROC of three classifiers are improved by 4.20% and 0.040, respectively. This proves the domain gap between synthetic and real samples is mitigated by introducing the image prior. Moreover, as shown in Table IV, conditioning the generator on the motion field produces more content-consistent and smoother samples, leading to continuous improvements in diagnostic performance. It can also be observed that the FVD results show a limited correlation with downstream evaluation metrics, confirming the finding in [7].
TABLE IV: Ablation study for different conditional controls in TUSC [69], including class label (C), text (T), image prior (I), and motion field (MF). SlowFast [81] (C1), CSN [82] (C2), and FTC [78] (C3) trained on Joint-train paradigm were used for downstream diagnosis. “Hybrid”, the classifier with CNN-Transformer design.
Dataset Synthesis Downstream Diagnosis
Control Metric Backbone $\Delta$Accuracy $\Delta$AUROC
[…]
TUSC [69] 4.18 90.40% 95.03% C1 CNN $\downarrow$ 1.80% $\downarrow$ 0.015
C2 CNN $\uparrow$ 1.44% $\uparrow$ 0.022
C3 Hybrid $\uparrow$ 1.08% $\uparrow$ 0.030
7.01 90.51% 96.91% C1 CNN $\downarrow$ 1.08% $\downarrow$ 0.001
C2 CNN $\downarrow$ 0.36% $\downarrow$ 0.008
C3 Hybrid $\downarrow$ 1.44% $\downarrow$ 0.011
5.59 91.56% 94.66% C1 CNN $\uparrow$ 1.80% $\uparrow$ 0.017
C2 CNN $\uparrow$ 1.08% $\uparrow$ 0.014
C3 Hybrid $\uparrow$ 2.16% $\uparrow$ 0.016
3.69 92.14% 95.59% C1 CNN $\uparrow$ 2.53% $\uparrow$ 0.030
C2 CNN $\uparrow$ 3.61% $\uparrow$ 0.040
C3 Hybrid $\uparrow$ 3.57% $\uparrow$ 0.030
IV-F Impact of the Sequential Augmentation Module
To verify the impact of the proposed sequential augmentation module, as shown in Table V, we compared the performance of three generator variants on the synthesis and downstream tasks. Ours-S, Ours-SK, and Ours-SKM denote our ablation studies, including gradually adding sequential attention (’-S’), key-frame/slice attention (’-K’), and motion field attention (’-M’) to the generator with spatial inflation only. Table V shows that the downstream accuracy and AUROC improvements of Ours-SKM are significantly higher than those of Ours-S and Ours-SK in both datasets. In terms of quantitative assessment of sequential coherence, Ours-SKM achieves better results overall than other variants. These prove that this module can help the generator synthesize cross-frame/slice consistent and dynamic-smooth clips, which are diagnosis-promotive for downstream diagnostic tasks. Moreover, it can be observed that the FVD results across these three generators show no significant difference, revealing its limited value for evaluating diagnosis-oriented medical sequence synthesis tasks.
TABLE V: Ablation study for the sequential augmentation module of the generator in TUSC [69] and ACDC [70]. SlowFast [81] (C1), CSN [82] (C2), and FTC [78] (C3) trained on Joint-train paradigm were used for downstream diagnosis.
“Hybrid”, the classifier with CNN-Transformer design.
Dataset Synthesis Downstream Diagnosis
[…]
TUSC [69] Ours-S 3.76 87.44% 91.07% C1 CNN $\downarrow$ 0.36% $\downarrow$ 0.009
C2 CNN $\uparrow$ 1.08% $\uparrow$ 0.011
C3 Hybrid $\uparrow$ 1.44% $\uparrow$ 0.007
Ours-SK 4.53 91.10% 95.02% C1 CNN $\uparrow$ 0.72% $\uparrow$ 0.007
C2 CNN $\uparrow$ 2.88% $\uparrow$ 0.032
C3 Hybrid $\uparrow$ 3.24% $\uparrow$ 0.021
Ours-SKM 3.69 92.14% 95.59% C1 CNN $\uparrow$ 2.53% $\uparrow$ 0.030
C2 CNN $\uparrow$ 3.61% $\uparrow$ 0.040
C3 Hybrid $\uparrow$ 3.57% $\uparrow$ 0.030
ACDC [70] Ours-S 8.75 70.11% 87.62% C1 CNN $\uparrow$ 0.73% $\uparrow$ 0.015
C2 CNN $\downarrow$ 0.37% $\uparrow$ 0.000
C3 Hybrid $\downarrow$ 0.73% $\downarrow$ 0.005
Ours-SK 8.52 77.29% 92.60% C1 CNN $\uparrow$ 1.47% $\uparrow$ 0.021
C2 CNN $\uparrow$ 1.83% $\uparrow$ 0.032
C3 Hybrid $\uparrow$ 1.83% $\uparrow$ 0.014
Ours-SKM 8.18 77.08% 92.77% C1 CNN $\uparrow$ 2.20% $\uparrow$ 0.040
C2 CNN $\uparrow$ 2.93% $\uparrow$ 0.049
C3 Hybrid $\uparrow$ 3.43% $\uparrow$ 0.029
Figure 7: Carotid stenosis and heart disease diagnostic accuracy (%) comparison using two classifiers on our filtering strategies under Real-finetune and Joint-train paradigms. The accuracy gains over the Baseline are depicted.
[…]
Generative Training Paradigm Augmentation Method Carotid TUSC [69] ACDC [70]
Acc. $\Delta$ AUROC $\Delta$ Acc. $\Delta$ AUROC $\Delta$ Acc. $\Delta$ AUROC $\Delta$
/ None 75.51 0.718 59.28 0.532 80.55 0.681
/ TraAug 79.59 $\uparrow$ 4.08 0.737 $\uparrow$ 0.019 60.72 $\uparrow$ 1.44 0.559 $\uparrow$ 0.027 83.49 $\uparrow$ 2.94 0.733 $\uparrow$ 0.052
Real-finetune TraAug + GenAug 85.03 $\uparrow$ 9.52 0.813 $\uparrow$ 0.095 65.41 $\uparrow$ 6.13 0.611 $\uparrow$ 0.079 87.16 $\uparrow$ 6.61 0.795 $\uparrow$ 0.114
/ None 78.91 0.735 60.36 0.552 80.92 0.688
/ TraAug 80.95 $\uparrow$ 2.04 0.751 $\uparrow$ 0.016 61.08 $\uparrow$ 0.72 0.562 $\uparrow$ 0.010 81.28 $\uparrow$ 0.36 0.696 $\uparrow$ 0.008
Joint-train TraAug + GenAug 84.35 $\uparrow$ 5.44 0.811 $\uparrow$ 0.076 62.88 $\uparrow$ 2.52 0.578 $\uparrow$ 0.026 85.32 $\uparrow$ 4.40 0.762 $\uparrow$ 0.074
IV-G Effectiveness of the Noisy Synthetic Data Filter
As shown in Fig. 7, we validated the indispensable role of our proposed noisy synthetic data filter by assessing its impact on downstream diagnostic accuracy, using two classifiers trained on Real-finetune and Joint-train paradigms across the carotid and ACDC datasets. For instance, on the carotid dataset, the Joint-train CSN resorting to non-quality controlled synthetic clips (i.e., ’None’), displays no significant difference with its Baseline (80.95% vs. 80.27%), even an accuracy degradation for its VideoSwin counterpart (78.91% vs. 79.59%). Then, by implementing our CF and SF strategy, diagnostic accuracy is significantly improved by 2.04% and 3.40%, respectively. Equipped with both, it achieves an accuracy of 87.07%, outperforming the ’None’ by 6.12%.
IV-H Impacts of Traditional and Generative Augmentations
We tested the contributions of different data augmentations in three downstream diagnostic tasks. The traditional augmentations used in this experiment can be found in Table I. We report the performance gains of the two variants with data augmentations compared to the augmentation-free setup. For experimental fairness, we kept the training datasets for the None and TraAug variants the same size as the method combining traditional and generative augmentations. As presented in Table VI, the diagnostic performance shows noticeable improvements using the traditional augmentation method. Further coupled with the generative augmentation, the results show more substantial improvements than TraAug. We also observe that in the case of using traditional augmentation, adding extra real training samples yields limited performance gains, even a drop for ACDC. These prove that generative augmentation can effectively supplement traditional data augmentation and provide intrinsically diverse data characteristics for enhancing deep classifier learning, which support the viewpoints of [7].
V Conclusion and discussion
In this study, we present a new and general diffusion-based generative augmentation framework, named Ctrl-GenAug, to facilitate medical sequence classification by leveraging customized and diagnosis-reliable synthetic sequences. To the best of our knowledge, this is the first comprehensive study to investigate the impact of the controllable generative augmentation scheme on aiding medical sequence classification. To improve the quality of synthetic data for promoting downstream classification, we propose a multimodal conditions-guided medical sequence generator that ensures flexibly controllable synthesis across semantic, sequential, and data distribution aspects. Moreover, we propose a highly effective noisy synthetic data filter to better learn from synthetic data, including adaptively filtering diagnosis-inhibitive synthetic sequences at class semantics and sequential levels. This can better connect the synthesis task and the downstream one, thus further enhancing our sequence classification performance. Extensive experiments on 3 medical datasets across multiple organs and modalities, with 11 popular networks trained on 3 paradigms, show the effectiveness and generality of Ctrl-GenAug. Furthermore, we conduct extensive empirical analysis demonstrating that Ctrl-GenAug can be effectively leveraged to improve diagnostic performance in underrepresented high-risk populations and out-domain robustness. We believe that Ctrl-GenAug can serve as a robust and practical data augmentation tool for various clinical scenarios.
We then discuss the limitations of Ctrl-GenAug. First, the denoising diffusion implicit model (DDIM) sampling process is time-consuming due to the large inference steps. This limits the clinical practicality of Ctrl-GenAug for fast and large-scale production. Second, in out-domain conditions, the proposed framework requires target domain data during generator training, which can be tough to obtain in advance in clinical practice. In future work, we will adopt fast-sampling strategies (e.g., AMED-Solver [88]) to better balance the trade-off between sampling time and sample quality. Besides, we will incorporate test-time adaptation methods [89, 90] to provide a more user-friendly tool for out-domain applications in clinical settings. Last, we will attempt to further extend the proposed method to segmentation and detection tasks.
Supplementary Material
Figure S1: Additional synthetic results on Carotid (a-d), TUSC [69] (e-h), and ACDC [70] (i-l). We used the multimodal conditions joint training strategy [76] during training and flexibly dropped several during sampling.
TABLE S1: A brief review of studies in the medical field utilizing diffusion-based generative models to promote downstream tasks. We consider six aspects: the conditional control, the downstream task, the underlying architecture of the downstream network, the number of the downstream network, and the modality and type of the medical dataset on which the surveyed approaches were applied. Arch., architecture. Num., number. OCT, optical coherence tomography.
Study Conditional Control Downstream Task Downstream Arch. Downstream Net Num. Medical Modality Data Type
Shang et al. [49] disease class,
image readability
classification CNN-based,
Transformer-based
2 Fundus image
Farooq et al. [50] text classification CNN-based,
Transformer-based
2 Dermoscopy image
Zhang et al. [51] text, edge segmentation CNN-Transformer hybrid-based 1 US, CT, MRI image
Sagers et al. [91] text classification CNN-based 1 Dermoscopy image
Khosravi et al. [92] attribute, 14 pathology labels classification CNN-based 1 X-ray image
Yu et al. [93] pathology labels,
lesion bounding boxes,
device types,
pathology-specific knowledge (NCM/CAL labels)
classification Transformer-based 1 US image
Ktena et al. [25] disease class, attribute classification CNN-based 1 Histopathology,
X-ray,
Dermoscopy
image
Luo et al. [7] disease class classification CNN-based 10 Fundus,
Dermoscopy,
X-ray,
US
image
Li et al. [24] multi-scale latent
representation
classification Transformer-based 1 Endoscopy video
Gong et al. [94] mask segmentation CNN-based 1 MRI volume
Peng et al. [95] mask segmentation CNN-based, CNN-Transformer hybrid-based 3 CT volume
Kebaili et al. [96] tumor attribute, mask segmentation CNN-based 1 MRI volume
Yu et al. [97] mask segmentation CNN-based 1 CT volume
Chen et al. [98] mask segmentation CNN-based,
CNN-Transformer hybrid-based
3 CT volume
Dorjsembe et al. [99] mask segmentation CNN-based 1 MRI volume
Huang et al. [100] mask segmentation CNN-based 2 OCT volume
Ours disease class, text,
image prior, motion field
classification CNN-based, Transformer-based, CNN-Transformer hybrid-based 11 US, MRI video,
volume

### v2

@@
 A sequential augmentation module is integrated to enhance the temporal/stereoscopic coherence of generated samples.
-Then, we propose a noisy synthetic data filter to suppress unreliable cases at semantic and sequential levels.
+Then, we propose a noisy synthetic data filter to suppress unreliable cases at the semantic and sequential levels.
 Extensive experiments on 3 medical datasets, using 11 networks trained on 3 paradigms, comprehensively analyze the effectiveness and generality of Ctrl-GenAug, particularly in underrepresented high-risk populations and out-domain conditions.
-Index Terms:
-Classification, Controllable Generative Augmentation, Medical Sequence Synthesis, Diffusion Model.
-publicationid: pubid: This work has been submitted to the IEEE for possible publication. Copyright may be transferred without notice, after which this version may no longer be accessible.
-I Introduction
+keywords:
+Classification, Controllable Generative Augmentation, Medical Sequence Synthesis, Diffusion Model
+journal: ***
+1 Introduction
 Dynamic information in medical imaging (e.g., video sequences) plays a vital role in clinical diagnosis.
-Recent deep learning-based classifiers have shown the ability to improve the diagnostic accuracy of different diseases [1, 2].
-Despite rapid advancements, the performance of current advanced solutions is still limited by several issues: 1) The scarcity of dynamic clinical cases, coupled with the high annotation cost, limits data availability; 2) Imbalanced data distribution, driven by the rarity of high-risk positive cases, skews model training; 3) Deep models are prone to brittle performance degradation when tested on out-domain data, e.g., cases from different medical centers [3, 4].
+Recent deep learning-based classifiers have shown the ability to improve the diagnostic accuracy of different diseases (Zhou et al., 2023; Wang et al., 2024c).
+Despite rapid advancements, the performance of current advanced solutions is still limited by several issues: 1) The scarcity of dynamic clinical cases, coupled with the high annotation cost, limits data availability; 2) Imbalanced data distribution, driven by the rarity of high-risk positive cases, skews model training; 3) Deep models are prone to brittle performance degradation when tested on out-domain data, e.g., cases from different medical centers (Varoquaux and Cheplygina, 2022; Huang et al., 2022).
 This risk presents a significant obstacle to deploying models in real-world medical settings.
-To address these issues, there is growing interest in generative augmentation paradigms, which employ advanced generative models to synthesize medical samples, thereby augmenting relevant diagnostic tasks.
-Denoising diffusion probabilistic models (DDPMs), which utilize explicit likelihood estimation and a progressive sampling process for data synthesis, have more well-established mathematical explanations and abilities to achieve stable, controllable, and diverse synthesis than previous generative methods like generative adversarial networks (GANs) [5, 6, 7].
-Diffusion models have garnered remarkable success in several natural image fields, including static image generation [8, 9], video synthesis [10, 11, 12, 13, 14, 15, 16, 17], video editing [18, 19, 20], and image animation [21, 22].
-Recently, researchers have explored high-fidelity medical sequence synthesis by applying the above approaches within the medical field [23, 24].
-Most recently, pioneers deeply investigated the impact of the diffusion-based generative augmentation scheme on solving data scarcity [7, 25] and domain generalization [25] issues in diagnostic tasks.
+To address these issues, generative augmentation paradigms have employed advanced generative models to synthesize medical samples, thereby augmenting relevant diagnostic tasks.
+Denoising diffusion probabilistic models (DDPMs), which utilize explicit likelihood estimation and progressive sampling, have more well-established mathematical explanations and abilities to achieve stable, controllable, and diverse data synthesis than previous generative methods like generative adversarial networks (GANs) (Croitoru et al., 2023; Dhariwal and Nichol, 2021; Luo et al., 2024).
+Diffusion models have garnered remarkable success in several natural image fields, including static image generation (Rombach et al., 2022; Zhan et al., 2023), video synthesis (Singer et al., 2022; Zhou et al., 2022; Ge et al., 2023; Ho et al., 2022b; Wu et al., 2023a; Wang et al., 2024a; Ren et al., 2024; Han et al., 2024), video editing (Khachatryan et al., 2023; Zhang et al., 2023; Qi et al., 2023), and image animation (Guo et al., 2024; Uzolas et al., 2024).
+Recently, researchers have explored high-fidelity medical sequence synthesis by applying the above approaches within the medical field (Zhou et al., 2024a; Li et al., 2024a).
+Most recently, pioneers deeply investigated the impact of the diffusion-based generative augmentation scheme on solving data scarcity (Luo et al., 2024; Ktena et al., 2024) and domain generalization (Ktena et al., 2024) issues in diagnostic tasks.
 However, they failed to fully ensure the reliability of synthetic medical images for classification due to limited semantic steerability and absent quality control over the generated images.
-Moreover, their techniques focus on image-level synthesis rather than sequence-level synthesis, which is crucial for medical modalities like MRI, ultrasound (US), etc [7].
+Moreover, their techniques focused on image-level synthesis rather than sequence-level synthesis, which is crucial for medical modalities like MRI, ultrasound (US), etc (Luo et al., 2024).
 Thus, our study plans to devise a diagnosis-reliable controllable generative augmentation framework to facilitate accurate and robust medical sequence classification.
@@
 1(c)) of the same disease category exhibit large visual variances (e.g., shapes, intensities, positions, etc).
-This may confuse diffusion learning seriously, thus causing uncontrollable and unreliable generation.
+This may seriously confuse diffusion learning, thus causing uncontrollable and unreliable generation.
 Second, artifacts and noises (e.g., US speckle noise) may prevent models from accurately perceiving vital anatomical targets, hindering the synthesis of high-fidelity sequences.
@@
 1(c)) challenge the coherence of synthesized sequences.
-The fourth challenge lies in the domain gap between synthetic and real samples, where achieving real-domain customization is crucial for effectively utilizing synthetic samples in downstream learning [26].
+The fourth challenge lies in the domain gap between synthetic and real samples, where achieving real-domain customization is crucial for effectively utilizing synthetic samples in downstream learning (He et al., 2023).
 Lastly, even a well-designed sequence generator cannot always guarantee high-quality synthesis due to random sampling, and unsatisfactory synthetic samples may negatively impact subsequent classifier learning.
@@
 Our contributions are three-fold:
+  • 1.
     We propose a multimodal conditions-guided sequence generator to ensure controllable and high-fidelity synthesis.
@@
 Moreover, we design a sequential augmentation module to promote dynamics modeling, thus improving the temporal or stereoscopic coherence of generated data.
+  • 2.
     We introduce a noisy synthetic data filter to suppress harmful synthetic sequences at class semantics and sequential levels.
 By effectively reducing noise and enhancing the reliability of the synthetic databases, we can associate the synthetic samples with downstream tasks, thus better improving the classification performance.
+  • 3.
     Comprehensive experiments on 3 medical datasets, with 11 downstream networks trained on 3 paradigms, show that our proposed framework can improve diagnostic performance, especially within underrepresented high-risk sets.
 Ctrl-GenAug also increases out-domain robustness, further indicating its practicality in clinical scenarios.
-It is noted that the sequence generator extends our prior work presented at MICCAI [23], with significant improvements in controllability at both semantic and sequential levels.
+It is noted that the sequence generator extends our prior work presented at MICCAI (Zhou et al., 2024a), with significant improvements in controllability at both semantic and sequential levels.
 We highlight that customized semantics and consistent sequential dynamics are crucial for enhancing the learning capability of downstream classifiers.
@@
 Besides, we propose a sequential augmentation module to further ensure the sequential consistency and smoothness of generated samples.
-II Related Works
-In this section, we briefly review the approaches for controllable sequence synthesis using diffusion models.
-Given the strong inter-slice association in 3D medical volumes [27], we treat them as video sequence data in our study and summarize related video-based methods.
+2 Related Works
+In this part, we briefly review the controllable sequence synthesis approaches using diffusion models.
+Given the strong inter-slice association in 3D medical volumes (Zhu et al., 2024), we treat them as video sequence data in our study and summarize related video-based methods.
 Finally, we deeply involve existing diffusion-based generative augmentation schemes.
-II-A Controllable Video Synthesis with Diffusion Models
-Diffusion models, a class of generative models, have recently attracted significant attention, especially text-to-video (T2V) synthesis [28, 29, 10, 11].
+2.1 Controllable Video Synthesis with Diffusion Models
+Diffusion models, a class of generative models, have recently attracted significant attention, especially text-to-video (T2V) synthesis (Blattmann et al., 2023; Ho et al., 2022a; Singer et al., 2022; Zhou et al., 2022).
 However, these methods did not achieve precise control over the visual appearance and dynamics of the synthesized video, resulting in limited practicality.
 To mitigate this issue, researchers have concentrated on incorporating multimodal conditions into T2V frameworks to guide controllable synthesis.
-Specifically, AnimateDiff [30] and MoonShot [31] introduced text and image conditions to facilitate the control of visual appearance.
-Most recently, several studies [19, 32, 33] achieved fine-grained spatial compositional control by integrating additional diverse conditions (i.e., sketch, mask, and depth sequences).
-Meanwhile, two spatial-sequential (S2) modeling strategies (fully attention layers [19] and S2 condition encoder [33]) were introduced to capture the dynamics of sequential conditions, promoting the inter-frame features interaction.
-Besides, dense optical flow [34] was adopted to control sequential consistency [35, 36].
+Specifically, AnimateDiff (Guo et al., 2023) and MoonShot (Zhang et al., 2024a) introduced text and image conditions to control the visual appearance.
+Most recently, several studies (Zhang et al., 2023; Zhao et al., 2023; Wang et al., 2024b) achieved fine-grained spatial compositional control by integrating additional diverse conditions (i.e., sketch, mask, and depth sequences).
+Meanwhile, two spatial-sequential (S2) modeling strategies (fully attention layers (Zhang et al., 2023) and S2 condition encoder (Wang et al., 2024b)) were introduced to capture the dynamics of sequential conditions, promoting the inter-frame features interaction.
+Besides, dense optical flow (Zach et al., 2007) was adopted to control sequential consistency (Liang et al., 2023; Ni et al., 2023).
 The above methods have shown promising controllability in natural video synthesis.
@@
 Hence, they may not suit the medical video synthesis tasks.
-In the medical field, several diffusion-based methods have been proposed to generate photorealistic echocardiography videos using the conditional guidance of a single frame [37], ejection fractions [37], single semantic mask [38, 39], and single sketch image [23].
+In the medical field, several diffusion-based methods have been proposed to generate photorealistic echocardiography videos using the conditional guidance of a single frame (Reynaud et al., 2023), ejection fractions (Reynaud et al., 2023), single semantic mask (Van Phi et al., 2024; Nguyen et al., 2024), and single sketch image (Zhou et al., 2024a).
 Zhou et al.
-[23] additionally leveraged mitral valve (MV) skeletons to control the complex motion trajectories of MV in the generated echocardiograms.
+(Zhou et al., 2024a) additionally leveraged mitral valve (MV) skeletons to control the complex motion trajectories of MV in the generated echocardiograms.
 Besides, Li et al.
-[24] proposed an approach to generate endoscopy videos that simulate clinical scenes by integrating a spatial-temporal transformer with 2D vision foundation model priors.
+(Li et al., 2024a) proposed to generate endoscopy videos that simulate clinical scenes by integrating a spatial-temporal transformer with 2D vision foundation model priors.
 However, the above methods may have the following drawbacks: unsteerable semantics due to the lack of attribute guidance (e.g., category or target shape), poor control over sequences, etc.
 Thus, they may be unsuitable for direct adoption in the generative augmentation tasks.
-II-B Generative Augmentation with Diffusion Models
-Synthetic data can effectively supplement traditional data augmentation (e.g., rotation) and provide diverse samples to help deep model training [7, 40].
-Previous works have succeeded in utilizing synthetic data generated with diffusion models to enhance natural image classification [26, 41, 42, 43], segmentation [44, 45, 46], and detection [47].
+Figure 2: t-SNE visualization of features of synthetic and real training data by a pre-trained I3D (Carreira and Zisserman, 2017) model on our three datasets.
+2.2 Generative Augmentation with Diffusion Models
+Synthetic data can provide intrinsically diverse characteristics and visual appearances (Luo et al., 2024; Pan et al., 2023), thus enhancing model learning.
+This aligns with the observation shown in Fig.
+2.
+Previous works have succeeded in utilizing synthetic data by diffusion models to enhance natural image classification (He et al., 2023; Sarıyıldız et al., 2023; Trabucco et al., 2023; Azizi et al., 2023), segmentation (Sankaranarayanan et al., 2018; Wu et al., 2023b; Yang et al., 2024), and detection (Feng et al., 2024).
 Recently, Singh et al.
-[48] performed an in-depth analysis of models trained with synthetic data across various robustness measures, and verified that they achieved good suitability in real-world settings.
-To address data scarcity and promote diagnostic performance, the intelligent medical field has raised increasing interest in exploring diffusion-based generative augmentation for aiding different downstream diagnostic tasks [49, 50, 25, 51, 52, 53].
+(Singh et al., 2024) performed an in-depth analysis of models trained with synthetic data across various robustness measures, and verified that they achieved good suitability in real-world settings.
+To address data scarcity and promote diagnostic performance, the intelligent medical field has raised increasing interest in exploring diffusion-based generative augmentation for aiding different downstream diagnostic tasks (Shang et al., 2023; Farooq et al., 2024; Ktena et al., 2024; Zhang et al., 2024c; Dorjsembe et al., 2024b; Li et al., 2024b).
 Luo et al.
-[7] marked the first focus on comprehensively analyzing the impact of their designed uncertainty-guided diffusion models on downstream diagnostic tasks.
+(Luo et al., 2024) marked the first focus on comprehensively analyzing the impact of their designed uncertainty-guided diffusion models on downstream diagnostic tasks.
 Nevertheless, current studies have two main limitations:
 On the one hand, due to the lack of adequate semantic and sequential guidance during sampling, coupled with the absence of post-sampling quality control for generated samples, existing works could hardly fully ensure the reliability of synthetic databases for downstream disease recognition.
-For instance, it is desirable to embed descriptive text (e.g., lesion morphology) to provide more informative guidance associated with disease grading.
+For instance, descriptive text (e.g., lesion morphology) can provide informative guidance related to disease grading.
 Unlike previous solutions, we aim to develop a generative augmentation framework that enables highly semantic- and sequential-customized sequence synthesis and suppresses incorrectly synthesized samples to better facilitate downstream classifier learning.
@@
 A brief review of studies in the medical field utilizing diffusion models to promote downstream tasks can be found in the Supplementary Material.
-III Method
+3 Method
 Figure 3: Pipeline of using our proposed framework to facilitate medical sequence recognition, which can be worked with a variety of classifiers.
 Here, we use the carotid plaque US video sequence as an example to demonstrate the overall process.
-III-A Model Overview
+3.1 Model Overview
 Fig.
@@
 To formulate customized, high-fidelity, and coherent synthetic databases to boost classification, we design the whole framework from model-centric, i.e., sequence generator (Secs.
-III-BIII-C, and III-D) and data-centric, i.e., data filter (Sec.
-III-E) perspectives.
-III-B Basic Architecture of Sequence Generator
+3.23.3, and 3.4) and data-centric, i.e., data filter (Sec.
+3.5) perspectives.
+3.2 Basic Architecture of Sequence Generator
 Fig.
 4 shows the pipeline of our proposed sequence generator.
-It supports customized and high-quality medical sequence generation via multimodal conditions.
-The generator is implemented with a two-stage training scheme.
+It supports customized and high-quality medical sequence generation via multimodal conditions, using a two-stage training scheme.
 In the pretraining stage, it attends to high-fidelity visual features learning for controllable image synthesis.
@@
 We then provide a short background of video diffusion models in Sec.
-III-B1, followed by a detailed description of the basic architecture of our generator in Secs.
-III-B2 and III-B3.
-III-B1 Preliminaries of Video DDPMs
-They are trained to learn complex video distribution by iteratively denoising corrupted inputs using pixel-space diffusion methods [54, 55].
-Given a training video $\mathit{x}_{0}\sim\mathit{q}$ and let $\mathit{T}$ the timestep, DDPMs first produce a sequence of noisy inputs via a diffusion process $\mathit{q}(\mathit{x}_{t}|\mathit{x}_{0},\mathit{t}),\mathit{t}\in{1,2,...,\mathit{T}}$, which progressively adds Gaussian noise $\epsilon\in\mathcal{N}(\mathbf{0},\mathbf{I})$ to $\mathit{x}_{0}$.
-To ease the computational burden due to pixel-space training, video latent diffusion models [28] perform the diffusion process in latent space of a variational autoencoder [56] (VAE).
+3.2.1, followed by a detailed description of the basic architecture of our generator in Secs.
+3.2.2 and 3.2.3.
+3.2.1 Preliminaries of Video DDPMs
+Video DDPMs aim to learn complex distributions by iteratively denoising corrupted inputs using diffusion methods (Sohl-Dickstein et al., 2015; Ho et al., 2020).
+Given a video $\mathit{x}_{0}\sim\mathit{q}$ and timestep $\mathit{T}$, DDPMs first produce a sequence of noisy inputs via diffusion process $\mathit{q}(\mathit{x}_{t}|\mathit{x}_{0},\mathit{t}),\mathit{t}\in{1,2,...,\mathit{T}}$, which progressively adds Gaussian noise $\epsilon\in\mathcal{N}(\mathbf{0},\mathbf{I})$ to $\mathit{x}_{0}$.
+To ease the computational burden of traditional pixel-space training, video latent diffusion models (Blattmann et al., 2023) perform the diffusion process in latent space of a variational autoencoder (Esser et al., 2021) (VAE).
 The model is then trained to estimate the parameterized Gaussian transition $\mathit{p}(\mathit{x}_{t-1}|\mathit{x}_{t})$ via a denoising network $\theta$.
@@
 $\epsilon_{\theta}$ and $\epsilon$ represent the predicted and target noise, respectively.
-Here, $\mathbf{c}$ denotes the (optional) control signal that the model can be conditioned on.
+$\mathbf{c}$ denotes the control signal (optional).
 For our generator, $\mathbf{c}$ involves multimodal conditions including semantic and sequential ones.
-III-B2 Factorized Learning of Visual Features and Sequential Patterns
-Developing a generator with pure 3D architecture design is an intuitive scheme for sequence synthesis [28].
-However, it is very challenging to simultaneously learn visual features and sequential patterns due to the scarcity and complexity of medical data, while causing high training costs.
+3.2.2 Factorized Learning of Visual Features and Sequential Patterns
+Developing a generator with pure 3D architecture design is an intuitive scheme for sequence synthesis (Blattmann et al., 2023).
+However, due to the scarcity and complexity of medical data, learning visual features and sequential patterns at the same time is very challenging and imposes high training costs.
 To this end, we propose to factorize both learning by first pretraining a latent diffusion model (LDM) and then finetuning its sequence counterpart.
 In this way, the generator enables realistic sequence synthesis, while easing the training burden.
-III-B3 2D-to-3D Model Inflation
-In the pretraining phase, we develop a 2D UNet [57] in LDM to predict noises for image synthesis.
-Referring to [13], we extend the 2D UNet to a 3D version via an inflation scheme to build the sequence LDM.
+3.2.3 2D-to-3D Model Inflation
+During pretraining, we develop a 2D UNet (Ronneberger et al., 2015) in LDM to predict noises for image synthesis.
+Referring to (Ho et al., 2022b), we extend the 2D UNet to a 3D version via an inflation scheme to build the sequence LDM.
 Specifically, all the spatial convolution layers are inflated to pseudo-3D counterparts by expanding the kernels at the sequential dimension (e.g., $3\times 3\rightarrow 1\times 3\times 3$ kernel).
 Besides, we perform sequential insertion by adding sequential attention (SA) layers (see Fig.
-3).
+4).
 Fig.
-4(a) shows the principle of the SA mechanism, where each patch queries to those at the same spatial position and across frames/slices.
-This design encourages the generator to model the sequential patterns, while not significantly altering visual feature distribution baked in the LDM [31].
+5(a) shows the principle of the SA mechanism, where each patch queries to those at the same spatial position and across frames/slices.
+This design encourages the generator to model the sequential patterns, while not significantly altering visual feature distribution baked in the LDM (Zhang et al., 2024a).
 Overall, sequence LDM enables inheriting the rich visual concepts preserved in LDM and focusing on sequential pattern aggregation, making the model learning efficient.
-Figure 3: Pipeline of our proposed sequence generator. The MedSAM image encoder [58] is adopted for domain-specific image prior feature extraction.
-III-C Multimodal Conditions Guidance
+Figure 4: Pipeline of our proposed sequence generator. The MedSAM image encoder (Huang et al., 2024b) is used for domain-specific image prior feature extraction.
+3.3 Multimodal Conditions Guidance
 Sequence generators guided by a single condition (e.g., text) have the following drawbacks: 1) poor controllability and 2) sequential inconsistency.
@@
 We consider four multimodal conditions to ensure comprehensive and accurate control over the sequence synthesis procedure (see Figs.
-2-3).
+3-4).
 Specifically, these conditions are divided into semantic and sequential ones for visual appearance control and serial guidance, respectively.
@@
 Details are described below.
-III-C1 Semantic Conditions for Visual Appearance Control
+3.3.1 Semantic Conditions for Visual Appearance Control
 As shown in Fig. 4, the sequence generator exploits three semantic conditions to perform visual appearance control, thus achieving controllable and high-fidelity synthesis.
 - Descriptive Text: It provides an intuitive indication of the coarse-grained semantic concepts of sequences.
-In our implementation, we pre-align the visual and textual features by finetuning the widely used CLIP model [59] using our medical datasets to mitigate the large gap between them.
+In our implementation, we pre-align the visual and textual features by finetuning the widely used CLIP model (Radford et al., 2021) using our medical datasets to mitigate the large gap between them.
 Then, semantic information of texts can feasibly query relevant medical visual patterns, thus easing subsequent training.
-Similar to [8], the text guides the generator via cross-attention layers.
+Similar to (Rombach et al., 2022), the text guides the generator via cross-attention layers.
 - Class Label: Generating medical data with the expected disease class is crucial for enhancing downstream classification, particularly improving the diagnostic accuracy of underrepresented high-risk sets.
@@
 We further propose a dedicated insertion method for this purpose.
-First, similar to the timestep embedding process [8] in diffusion models, we adopt a class label encoder composed of a label discretization layer and an embedding layer to obtain its embeddings.
+First, similar to the timestep embedding process (Rombach et al., 2022) in diffusion models, we adopt a class label encoder composed of a label discretization layer and an embedding layer to obtain its embeddings.
 The addition between the timestep and class label embeddings is then input to the spatial convolution layer in each UNet block for visual appearance control.
@@
 As image prior offers global semantics like texts, we align image prior features $f_{i}$ extracted from image prior encoder (see Fig.
-3) with text embeddings $f_{t}$ generated by our medical data-specific CLIP text encoder for joint guidance.
-Specifically, motivated by [60], we replace the text-guided cross-attention layers in original UNet blocks with decoupled counterparts that handle texts and image priors in parallel and then merge the results by addition.
+4) with text embeddings $f_{t}$ generated by our medical data-specific CLIP text encoder for joint guidance.
+Specifically, motivated by (Ye et al., 2023), we replace the text-guided cross-attention layers in original UNet blocks with decoupled counterparts that handle texts and image priors in parallel and then merge the results by addition.
 The decoupled attention can be formulated as:
@@
 $\sigma(\cdot)$, softmax function.
-III-C2 Sequential Condition for Serial Guidance
-Apart from perceiving visual concepts, modeling sequential knowledge is also important in sequence synthesis, with sequential cues playing a vital role in this process.
-Most existing methods [35, 36] adopted dense optical flow [34] to promote sequential dynamics modeling.
-However, optical flow extraction requires high computational demands [33].
+3.3.2 Sequential Condition for Serial Guidance
+Apart from perceiving visual concepts, modeling sequential knowledge is also important in sequence synthesis, with sequential cues playing a vital role.
+Most existing methods (Liang et al., 2023; Ni et al., 2023) adopted dense optical flow (Zach et al., 2007) to promote sequential dynamics modeling.
+However, optical flow extraction requires high computational demands (Wang et al., 2024b).
 Thus, we instead introduce computation-efficient motion field as a sequential condition in sequence LDM.
-It explicitly showcases the pixel-wise motions between adjacent frames/slices (see Figs.
-2-3).
-In our approach, we first extract the motion fields of real sequences using a Python package [61].
-Then, a motion encoder (see Fig.
-3) receives the motion fields and produces motion features.
+It explicitly showcases the pixel-wise motions between adjacent frames/slices (Figs.
+3-4).
+In our approach, we first extract the motion fields of real sequences using a Python package (Bommes et al., 2020).
+Then, a motion encoder (Fig.
+4) receives the motion fields and produces motion features.
 Last, the features are concatenated with latent representations $\mathit{z}$ from VAE along the channel dimension for serial guidance.
 Figure 5: Schematics of three attention mechanisms for sequential modeling (a-c) and our proposed sequential augmentation module in this study (d).
-III-D Sequential Augmentation Module
+3.4 Sequential Augmentation Module
 Solely equipping with SA layers and sequential cues in sequence LDM may present inadequate consistency and coherence across synthetic frames/slices.
@@
 As shown in Fig.
-4(d), SAM integrates two attention mechanisms in cascade for sequential augmentation.
-III-D1 Key-frame/slice Attention
+5(d), SAM integrates two attention mechanisms in cascade for sequential augmentation.
+3.4.1 Key-frame/slice Attention
 The common spatial self-attention layers can lead to sequential inconsistency due to the lack of interaction across frames/slices.
@@
 […]
-III-D2 Motion Field Attention
+3.4.2 Motion Field Attention
 To further boost the sequential coherence of the generated sequences, we propose to reuse the motion field condition (refer to Sec.
-III-C2) by introducing a motion field attention (MFA) mechanism after KA.
+3.3.2) by introducing a motion field attention (MFA) mechanism after KA.
 As shown in Fig.
-4(c), MFA requires the patches to communicate with those in the same motion field-based pathway including itself, which eliminates flickers of the generated sequences to make the contents visually smooth.
-Inspired by [62], MFA is implemented in two steps: a) motion field-based patch pathway sampling and b) attention calculation.
+5(c), MFA requires the patches to communicate with those in the same motion field-based pathway including itself, which eliminates flickers of the generated sequences to make the contents visually smooth.
+Inspired by (Cong et al., 2024), MFA is implemented in two steps: a) motion field-based patch pathway sampling and b) attention calculation.
 In step 1, we sample the patch pathways based on the m-scaled downsampled motion fields.
@@
 43:return $\mathrm{S_{3}}$triangleright final synthetic database for downstream training
-III-E Noisy Synthetic Data Filter
+3.5 Noisy Synthetic Data Filter
 We use the proposed sequence generator to constitute the synthetic samples set.
@@
 Although visually realistic and smooth (good cases, Fig.
-5), sequence synthesis may still suffer from class semantics misalignment, cross-frame/slice inconsistency or over-consistency (i.e., almost static clip), and inter-clip similarity.
+6), sequence synthesis may still suffer from class semantics misalignment, cross-frame/slice inconsistency or over-consistency (i.e., almost static clip), and inter-clip similarity.
 For instance, in Fig.
-5(e), the synthesized carotid clip category is wrong, which should be moderate rather than mild.
+6(e), the synthesized carotid clip category is wrong, which should be moderate rather than mild.
 In Fig.
-5(f), the generated clip includes abrupt changes in anatomical structures.
+6(f), the generated clip includes abrupt changes in anatomical structures.
 Hence, blindly using all synthetic clips for classifier learning will significantly cause a performance drop due to noisy samples.
@@
 This can also link the synthetic data to downstream tasks, thus potentially achieving a higher performance upper bound.
-III-E1 Class Semantics Misalignment Filtering
+3.5.1 Class Semantics Misalignment Filtering
 This scheme is designed to filter the generated clips whose visual contents mismatch the corresponding class labels.
-Motivated by anomaly detection [63], we consider that such noisy synthetic samples will produce large losses if fed into a powerful classifier trained on real ones.
+Motivated by anomaly detection (Kwon et al., 2020), we consider that such noisy synthetic samples will produce large losses if fed into a powerful classifier trained on real ones.
 Hence, a loss threshold can be set to identify and filter the anomalistic clips.
@@
 If the loss of a synthetic clip with $c$ exceeds the threshold $L_{c}$, it will be viewed as a noisy sample and excluded from the synthetic pool.
-III-E2 Sequential Filtering
+3.5.2 Sequential Filtering
 This strategy aims to filter noisy samples at a) inner-sequence and b) inter-sequence levels.
@@
 Specifically, we retain a synthetic clip whose cross-frame/slice consistency falls in a pre-computed range.
-It is determined by K-means clustering [64] based on the cross-frame/slice consistency of all synthetic clips.
-To assess the video sequence consistency, [33] tends to use CLIP [59] image embeddings to compute the average cosine similarity across consecutive frames.
+It is determined by K-means clustering (Macqueen, 1967) based on the cross-frame/slice consistency of all synthetic clips.
+To assess the video sequence consistency, (Wang et al., 2024b) tends to use CLIP (Radford et al., 2021) image embeddings to compute the average cosine similarity across consecutive frames.
 However, this method leads to inadequate evaluation due to the usage of limited informative CLIP embeddings.
@@
 We quantify the inter-clip similarity by calculating the cosine distance of VAE latent embeddings between frame/slice pairs in two clips and averaging the obtained distance values.
-III-F Evaluation
-As a common practice, existing sequence synthesis works evaluate the synthetic quality by mostly using Fréchet Inception Distance (FID) [65] and Fréchet Video Distance (FVD) [66].
-However, researchers have verified that these metrics do not consistently correlate with performance metrics on downstream tasks [46, 7].
-To exhaustively evaluate the values of synthetic samples to downstream tasks, eleven popular classifiers are deployed for comparison with respect to diagnostic accuracy and area under the receiver operating characteristic curve (AUROC).
-Besides, we adopt the proposed VAE-Seq to assess the cross-frame/slice consistency (refer to Sec.
-III-E2).
-Moreover, we consider it crucial to evaluate the smoothness of generated clips.
-Motivated by [67], we quantify it using the proposed metric (termed Dynamic Smoothness), which is the mean absolute error between the reconstructed frames/slices from the interpolation model [68] and the original real ones.
+3.6 Evaluation
+Existing sequence synthesis works evaluate the synthetic quality by mostly using Fréchet Inception Distance (FID) (Heusel et al., 2017) and Fréchet Video Distance (FVD) (Unterthiner et al., 2018).
+However, researchers have verified that these metrics do not consistently correlate with performance metrics on downstream tasks (Yang et al., 2024; Luo et al., 2024).
+To exhaustively evaluate the values of synthetic cases to downstream tasks, 11 popular classifiers are deployed for comparison with respect to diagnostic accuracy and area under the receiver operating characteristic curve (AUROC).
+Besides, we adopt the proposed VAE-Seq to assess the cross-frame/slice consistency (more details refer to Sec.
+3.5.2).
+Moreover, motivated by (Huang et al., 2024c), we quantify the smoothness of generated clips using the metric termed Dynamic Smoothness, which is the mean absolute error between the reconstructed frames/slices from the interpolation model (Li et al., 2023) and the original real ones.
 Figure 6: Typical good and bad synthetic samples of three datasets using different banks of conditions.
@@
 See more synthetic results in Supplementary Material.
-IV Experiments
-In this section, we first comprehensively validated the effectiveness and generality of our method on three medical datasets, using eleven popular networks trained on three paradigms.
-We then tested the performance of our method in underrepresented high-risk sets and out-domain conditions.
-Last, we analyzed the contributions of our proposed multimodal conditions guidance, sequential augmentation module, and noisy synthetic data filter respectively, as well as the impacts of traditional and generative augmentations.
-TABLE I: Experimental datasets and settings.
-Carotid TUSC [69] ACDC [70]
+4 Experiments
+We first comprehensively validated the effectiveness and generality of our Ctrl-GenAug on 3 medical datasets, using 11 popular networks trained on 3 paradigms.
+We then tested the performance of Ctrl-GenAug in underrepresented high-risk sets and out-domain conditions.
+Last, we analyzed the contributions of our multimodal conditions guidance, sequential augmentation module, and noisy synthetic data filter, respectively; as well as the impacts of traditional and generative augmentations.
+4.1 Datasets and Settings
+Carotid was collected by three medical centers with approval from local institutional review boards, including a) The Third Affiliated Hospital of Sun Yat-sen University, b) Shenzhen Longgang District People’s Hospital, and c) The Eighth Affiliated Hospital of Sun Yat-sen University.
+It consists of 273 US video sequences from 231 patients with carotid plaque-induced stenosis.
+The severity of carotid stenosis is graded into three stages, i.e., mild, moderate, and severe.
+We uniformly sampled several non-overlapping 8-frame clips from each sequence.
+The final dataset involved an in-domain (ID) subset (407 clips from 193 patients collected by Center (a)) and an out-domain (OD) subset (79 clips from 38 patients acquired from Center (b)-(c)).
+The ID/OD subset was randomly split into 309/39 and 98/40 clips for training and testing at the patient level.
+Each clip has up to two plaques and was labeled with stenosis grading and descriptive text by experienced sonographers using Pair annotation software package (Liang et al., 2022).
+The text annotations indicated the plaque characteristics of echogenicity and location (see Fig.
+3).
+We cropped the region of interest based on the pretrained carotid vessel detector (Liu et al., 2023) for easing the model learning.
+TUSC (Stanford-AIMI, 2021) contains 192 thyroid US video sequences from 167 patients collected by the Stanford University Medical Center, with each including one nodule.
+Each sequence provided a TI-RADS level and nodule descriptors (Tessler et al., 2017) (e.g., shape), the latter of which was used to form descriptive text.
+The TI-RADS level is ranging from 1 to 5.
+Considering the imbalanced distribution of TI-RADS levels, as suggested by the experienced sonographers, we rebuilt the dataset by removing the TI-RADS1 video sequence and combining the levels 2-3 sequences into a unified category named TI-RADS2-3.
+We uniformly sampled the sequences, obtaining 633 8-frame clips, with 448 for training and 185 for testing.
+Each sampled clip inherited the same TI-RADS level and text as the original sequence.
+The bounding box of the nodule region for each frame was extracted by radiologists, similar to (Yamashita et al., 2022).
+ACDC (Bernard et al., 2018) includes 4D cardiac cine-MRI from 150 patients acquired at the University Hospital of Dijon.
+They were evenly divided into five classes: normal patients (NOR), myocardial infarction (MINF), dilated cardiomyopathy (DCM), hypertrophic cardiomyopathy (HCM), and abnormal right ventricle (ARV).
+We first extracted volume sequences at end-systole and diastole phases from each case and assigned corresponding classes.
+Then, 8-slice clips were uniformly sampled from each sequence along the slice axis.
+Finally, 205 and 109 clips were obtained for training and testing.
+Text prompts were set to empty strings during generator training.
+Table 1 presents a detailed description of experimental datasets and settings.
+Using all available conditions, we synthesized 300 clips per category for the ACDC dataset and 500 clips per class for others.
+All frames/slices were resized to 256$\times$256.
+By default, we set m=8 for the MFA mechanism.
+For the sequence generator, the pretraining LDM was built upon Stable Diffusion (Rombach et al., 2022) and initialized with the public weights [^4] https://huggingface.co/CompVis/stable-diffusion-v1-4.
+During sequence LDM finetuning, we updated the parameters of SAM, SA layers, and motion encoder.
+We further refined the text- and image prior-sequence alignment by keeping the query projection matrices in cross-attention layers trainable.
+During generator training, we employed the multimodal conditions joint training strategy (Huang et al., 2023a).
+This allows users to flexibly select any single condition or combination during sampling, without finetuning the model for each specific combination.
+Motivated by (Yang et al., 2024), we adopted three paradigms to explore the impacts of the synthetic samples for downstream classifiers.
+Baseline: merely trained with real clips.
+Real-finetune: initially pretrained with synthetic clips and then finetuned with real ones.
+Joint-train: real clips were first over-sampled to match the number of synthetic clips, the classifier was then trained on the real-synthetic mixed set.
+All classification experiments were conducted on MMAction2 (Contributors, 2020) except for FTC (Ahmadi et al., 2023) and CSG-3DCT (Zhou et al., 2023).
+Table 1: Experimental datasets and settings. CF, Classifier-free.
+Carotid TUSC ACDC
 Patient 231 167 150
@@
 Inference
-Classifier-free guidance [71] factor(7.5)   Sampling steps(200)
+CF guidance (Ho and Salimans, 2022) factor(7.5)   Sampling steps(200)
 Eleven classifiers Adam(lr=1e-4)   lr scheduler(Cosine)   Epochs(100)   Batch size(8)
 Traditional augmentations(Random Color/Move/Gaussian/Rotation/Flip)
-IV-A Datasets and Settings
-Carotid dataset was collected by three medical centers with approval from local institutional review boards, including a) The Third Affiliated Hospital of Sun Yat-sen University, b) Shenzhen Longgang District People’s Hospital, and c) The Eighth Affiliated Hospital of Sun Yat-sen University.
-It consists of 273 US video sequences from 231 patients with carotid plaque-induced stenosis.
-The severity of carotid stenosis is graded into three stages, i.e., mild, moderate, and severe.
-We uniformly sampled several non-overlapping 8-frame clips from each sequence.
-The final dataset involved an in-domain (ID) subset (407 clips from 193 patients collected by Center (a)) and an out-domain (OD) subset (79 clips from 38 patients acquired from Center (b)-(c)).
-The ID/OD subset was randomly split into 309/39 and 98/40 clips for training and testing at the patient level.
-Each clip has up to two plaques and was labeled with stenosis grading and descriptive text by experienced sonographers using Pair annotation software package [72].
-The text annotations indicated the plaque characteristics of echogenicity and location (see Fig.
-2).
-We cropped the region of interest based on the pretrained carotid vessel detector [73] for easing the model learning.
-TUSC [69] contains 192 thyroid US video sequences from 167 patients collected by the Stanford University Medical Center, with each including one nodule.
-Each sequence provided a TI-RADS level and nodule descriptors [74] (e.g., shape), the latter of which was used to form descriptive text.
-The TI-RADS level is ranging from 1 to 5.
-Considering the imbalanced distribution of TI-RADS levels, as suggested by the experienced sonographers, we rebuilt the dataset by removing the TI-RADS1 video sequence and combining the levels 2-3 sequences into a unified category named TI-RADS2-3.
-We uniformly sampled the sequences, obtaining 633 8-frame clips, with 448 for training and 185 for testing.
-Each sampled clip inherited the same TI-RADS level and text as the original sequence.
-The bounding box of the nodule region for each frame was extracted by radiologists, similar to [75].
-ACDC [70] includes 4D cardiac cine-MRI from 150 patients acquired at the University Hospital of Dijon.
-The patients were evenly divided into five classes: normal patients (NOR), myocardial infarction (MINF), dilated cardiomyopathy (DCM), hypertrophic cardiomyopathy (HCM), and abnormal right ventricle (ARV).
-We first extracted 3D volume sequences at end-systole and diastole phases from each case and assigned corresponding classes.
-Then, the 8-slice clips were uniformly sampled from each sequence along the slice axis.
-Finally, 205 and 109 clips were constructed for training and testing.
-Text prompts were set to empty strings during generator training.
-Table I presents a detailed description of experimental datasets and settings.
-Using all available conditions, we synthesized 300 clips per category for the ACDC dataset and 500 clips per class for others.
-All frames/slices were resized to 256$\times$256.
-By default, we set m=8 for the MFA mechanism.
-For the sequence generator, the pretraining LDM was built upon Stable Diffusion [8] and initialized with the public weights [^2] https://huggingface.co/CompVis/stable-diffusion-v1-4.
-During sequence LDM finetuning, we updated the parameters of SAM, SA layers, and motion encoder.
-We further refined the text- and image prior-sequence alignment by keeping the query projection matrices in cross-attention layers trainable.
-During generator training, we employed the multimodal conditions joint training strategy [76].
-This allows users to flexibly select any single condition or combination during sampling, without finetuning the model for each specific combination.
-Motivated by [46], we adopted three paradigms to explore the impacts of the synthetic samples for downstream classifiers.
-Baseline: merely trained with real clips.
-Real-finetune: initially pretrained with synthetic clips and then finetuned with real ones.
-Joint-train: real clips were first over-sampled to match the number of synthetic clips, the classifier was then trained on the real-synthetic mixed set.
-All classification experiments were conducted on MMAction2 [77] except for FTC [78] and CSG-3DCT [1].
 […]
-Method Backbone Training Paradigm Carotid TUSC [69] ACDC [70]
+Method Backbone Training Paradigm Carotid TUSC (Stanford-AIMI, 2021) ACDC (Bernard et al., 2018)
 Acc. $\Delta$ AUROC $\Delta$ Acc. $\Delta$ AUROC $\Delta$ Acc. $\Delta$ AUROC $\Delta$
-I3D [79] CNN A 79.59 0.737 60.72 0.559 83.49 0.733
+I3D  (Carreira and Zisserman, 2017) CNN A 79.59 0.737 60.72 0.559 83.49 0.733
 B 85.03 $\uparrow$ 5.44 0.813 $\uparrow$ 0.076 65.41 $\uparrow$ 4.69 0.611 $\uparrow$ 0.052 87.16 $\uparrow$ 3.67 0.795 $\uparrow$ 0.062
 C 84.35 $\uparrow$ 4.76 0.811 $\uparrow$ 0.074 62.88 $\uparrow$ 2.16 0.578 $\uparrow$ 0.019 85.32 $\uparrow$ 1.83 0.762 $\uparrow$ 0.029
-R(2+1)D [80] CNN A 82.99 0.781 63.96 0.595 83.85 0.737
+R(2+1)D  (Tran et al., 2018) CNN A 82.99 0.781 63.96 0.595 83.85 0.737
 B 83.67 $\uparrow$ 0.68 0.797 $\uparrow$ 0.016 64.68 $\uparrow$ 0.72 0.601 $\uparrow$ 0.006 85.69 $\uparrow$ 1.84 0.769 $\uparrow$ 0.032
 C 85.04 $\uparrow$ 2.05 0.804 $\uparrow$ 0.023 65.41 $\uparrow$ 1.45 0.610 $\uparrow$ 0.015 85.69 $\uparrow$ 1.84 0.771 $\uparrow$ 0.034
-SlowFast [81] CNN A 80.95 0.742 62.88 0.579 81.65 0.702
+SlowFast  (Feichtenhofer et al., 2019) CNN A 80.95 0.742 62.88 0.579 81.65 0.702
 B 81.63 $\uparrow$ 0.68 0.770 $\uparrow$ 0.028 65.76 $\uparrow$ 2.88 0.609 $\uparrow$ 0.030 86.05 $\uparrow$ 4.40 0.772 $\uparrow$ 0.070
 C 84.35 $\uparrow$ 3.40 0.811 $\uparrow$ 0.069 65.41 $\uparrow$ 2.53 0.609 $\uparrow$ 0.030 83.85 $\uparrow$ 2.20 0.742 $\uparrow$ 0.040
-CSN [82] CNN A 80.27 0.757 61.80 0.570 85.69 0.765
+CSN  (Tran et al., 2019) CNN A 80.27 0.757 61.80 0.570 85.69 0.765
 B 84.35 $\uparrow$ 4.08 0.784 $\uparrow$ 0.027 67.21 $\uparrow$ 5.41 0.631 $\uparrow$ 0.061 87.15 $\uparrow$ 1.46 0.794 $\uparrow$ 0.029
 C 87.07 $\uparrow$ 6.80 0.843 $\uparrow$ 0.086 65.41 $\uparrow$ 3.61 0.610 $\uparrow$ 0.040 88.62 $\uparrow$ 2.93 0.814 $\uparrow$ 0.049
-TPN [83] CNN A 82.99 0.784 64.32 0.602 83.85 0.735
+TPN  (Yang et al., 2020) CNN A 82.99 0.784 64.32 0.602 83.85 0.735
 B 87.76 $\uparrow$ 4.77 0.856 $\uparrow$ 0.072 66.85 $\uparrow$ 2.53 0.633 $\uparrow$ 0.031 85.32 $\uparrow$ 1.47 0.764 $\uparrow$ 0.029
 C 85.71 $\uparrow$ 2.72 0.823 $\uparrow$ 0.039 65.41 $\uparrow$ 1.09 0.609 $\uparrow$ 0.007 84.95 $\uparrow$ 1.10 0.756 $\uparrow$ 0.021
-TimeSformer [84] Transformer A 77.55 0.721 65.41 0.614 72.48 0.562
+TimeSformer  (Bertasius et al., 2021) Transformer A 77.55 0.721 65.41 0.614 72.48 0.562
 B 80.95 $\uparrow$ 3.40 0.753 $\uparrow$ 0.032 66.49 $\uparrow$ 1.08 0.626 $\uparrow$ 0.012 76.51 $\uparrow$ 4.03 0.626 $\uparrow$ 0.064
 C 79.59 $\uparrow$ 2.04 0.730 $\uparrow$ 0.009 66.49 $\uparrow$ 1.08 0.618 $\uparrow$ 0.004 77.62 $\uparrow$ 5.14 0.642 $\uparrow$ 0.080
-MViTv2 [85] Transformer A 74.15 0.663 59.64 0.543 77.98 0.642
+MViTv2  (Li et al., 2022b) Transformer A 74.15 0.663 59.64 0.543 77.98 0.642
 B 78.91 $\uparrow$ 4.76 0.740 $\uparrow$ 0.077 63.96 $\uparrow$ 4.32 0.596 $\uparrow$ 0.053 78.72 $\uparrow$ 0.74 0.653 $\uparrow$ 0.011
 C 76.87 $\uparrow$ 2.72 0.701 $\uparrow$ 0.038 62.17 $\uparrow$ 2.53 0.568 $\uparrow$ 0.025 80.18 $\uparrow$ 2.20 0.680 $\uparrow$ 0.038
-VideoSwin [86] Transformer A 79.59 0.732 58.92 0.535 74.31 0.572
+VideoSwin  (Liu et al., 2022) Transformer A 79.59 0.732 58.92 0.535 74.31 0.572
 B 83.67 $\uparrow$ 4.08 0.789 $\uparrow$ 0.057 61.08 $\uparrow$ 2.16 0.551 $\uparrow$ 0.016 79.45 $\uparrow$ 5.14 0.663 $\uparrow$ 0.091
 C 83.67 $\uparrow$ 4.08 0.776 $\uparrow$ 0.044 61.44 $\uparrow$ 2.52 0.562 $\uparrow$ 0.027 77.62 $\uparrow$ 3.31 0.632 $\uparrow$ 0.060
-UniFormerV2 [87] Hybrid A 76.87 0.710 65.41 0.611 75.78 0.614
+UniFormerV2  (Li et al., 2022a) Hybrid A 76.87 0.710 65.41 0.611 75.78 0.614
 B 80.27 $\uparrow$ 3.40 0.743 $\uparrow$ 0.033 65.77 $\uparrow$ 0.36 0.618 $\uparrow$ 0.007 77.25 $\uparrow$ 1.47 0.645 $\uparrow$ 0.031
 C 81.63 $\uparrow$ 4.76 0.760 $\uparrow$ 0.050 66.13 $\uparrow$ 0.72 0.617 $\uparrow$ 0.006 77.61 $\uparrow$ 1.83 0.637 $\uparrow$ 0.023
-FTC [78] Hybrid A 76.99 0.723 62.09 0.584 80.95 0.689
+FTC  (Ahmadi et al., 2023) Hybrid A 76.99 0.723 62.09 0.584 80.95 0.689
 B 80.91 $\uparrow$ 3.92 0.750 $\uparrow$ 0.027 64.09 $\uparrow$ 2.00 0.593 $\uparrow$ 0.009 82.38 $\uparrow$ 1.43 0.703 $\uparrow$ 0.014
 C 81.80 $\uparrow$ 4.81 0.758 $\uparrow$ 0.035 65.66 $\uparrow$ 3.57 0.614 $\uparrow$ 0.030 84.38 $\uparrow$ 3.43 0.718 $\uparrow$ 0.029
-CSG-3DCT [1] Hybrid A 84.35 0.812 61.44 0.571 87.16 0.797
+CSG-3DCT  (Zhou et al., 2023) Hybrid A 84.35 0.812 61.44 0.571 87.16 0.797
 B 87.07 $\uparrow$ 2.72 0.831 $\uparrow$ 0.019 63.96 $\uparrow$ 2.52 0.590 $\uparrow$ 0.019 88.62 $\uparrow$ 1.46 0.819 $\uparrow$ 0.022
 C 87.76 $\uparrow$ 3.41 0.837 $\uparrow$ 0.025 65.05 $\uparrow$ 3.61 0.609 $\uparrow$ 0.038 88.63 $\uparrow$ 1.47 0.820 $\uparrow$ 0.023
-IV-B Generative Augmentation-based Sequence Classification
+4.2 Generative Augmentation-based Sequence Classification
 We utilized Real-finetune and Joint-train paradigms to integrate real and synthetic clips for downstream multi-class sequence classification training.
@@
 To demonstrate the generalizability of Ctrl-GenAug, the selected networks involve 3D CNN, transformer, and CNN-Transformer hybrid designs, which specialize in capturing local, global, and mixed features, respectively.
-Table II compares the average performance (i.e., accuracy and AUROC) on different training paradigms across a wide range of sequence classifiers, including five CNN-based classic models, three transformer-based models, and three CNN-Transformer hybrid-based methods.
+Table 2 compares the average performance (i.e., accuracy and AUROC) on different training paradigms across a wide range of sequence classifiers, including five CNN-based classic models, three transformer-based models, and three CNN-Transformer hybrid-based methods.
 It is evident that the performance ceiling of existing popular models, from simple CNN-based to complex CNN-Transformer hybrids, can be further elevated using synthetic samples obtained from our proposed framework, under Real-finetune and Joint-train paradigms.
@@
 Due to high intra-class variations in TUSC, the baseline exhibits visibly inferior results, with an average accuracy of 62.42% compared to other datasets.
-However, it can be seen from Table II that samples synthesized by Ctrl-GenAug make significant contributions and enhance the classification performance to an acceptance level.
+However, it can be seen from Table 2 that samples synthesized by Ctrl-GenAug make significant contributions and enhance the classification performance to an acceptance level.
 Among all the classifiers, CSN harvests the largest performance gain, which is non-trivially enhanced by 5.41%/0.061 and 3.61%/0.040 accuracy/AUROC under Real-finetune and Joint-train paradigms, respectively.
@@
 Extensive experimental results show that synthetic sample sets produced by Ctrl-GenAug can be powerful data engines to enhance various diagnostic tasks in medical sequence analysis.
-IV-C Improved Diagnosis in Underrepresented High-risk Sets
+4.3 Improved Diagnosis in Underrepresented High-risk Sets
 We tested the effectiveness of Ctrl-GenAug by evaluating the diagnostic performance in underrepresented high-risk situations using our carotid dataset.
@@
 Figure 7: Carotid performance radar map, using 4 models trained on Baseline (green) and Joint-train (blue) paradigms in underrepresented high-risk sets.
-IV-D Classification Robustness in Out-domain Conditions
+4.4 Classification Robustness in Out-domain Conditions
 We further explored the impact of synthetic samples from Ctrl-GenAug on diagnostic robustness in out-domain settings using the carotid dataset.
@@
 For the OD data, which did not contain the diagnostic class label, we padded the corresponding conditioning vector with zeros, while solely preserving hospital IDs to form descriptive texts.
-As shown in Table III, the average accuracy and AUROC of three classifiers display better results on the Real-finetune and Joint-train paradigms compared to the Baseline.
+As shown in Table 3, the average accuracy and AUROC of three classifiers display better results on the Real-finetune and Joint-train paradigms compared to the Baseline.
 It validates that diagnostic robustness can be enhanced with the aid of Ctrl-GenAug in scenarios where we only have access to unlabeled cases from additional medical centers due to limited resources.
@@
 Acc. $\Delta$ AUROC $\Delta$
-I3D [79] CNN A 70.00 0.609
+I3D  (Carreira and Zisserman, 2017) CNN A 70.00 0.609
 B 71.67 $\uparrow$ 1.67 0.648 $\uparrow$ 0.039
 C 75.00 $\uparrow$ 5.00 0.682 $\uparrow$ 0.073
-TPN [83] CNN A 75.00 0.667
+TPN  (Yang et al., 2020) CNN A 75.00 0.667
 B 76.67 $\uparrow$ 1.67 0.731 $\uparrow$ 0.064
 C 76.67 $\uparrow$ 1.67 0.678 $\uparrow$ 0.011
-CSG-3DCT [1] Hybrid A 68.33 0.599
+CSG-3DCT  (Zhou et al., 2023) Hybrid A 68.33 0.599
 B 73.33 $\uparrow$ 5.00 0.689 $\uparrow$ 0.090
 C 70.00 $\uparrow$ 1.67 0.651 $\uparrow$ 0.052
-IV-E Analysis of Multimodal Conditions Guidance
-To illustrate the role of different conditional guidance, we conduct comparative experiments with generators that resort to various banks of conditions for training and sampling.
-Table IV validates that all proposed conditions are effective for enhancing synthesis and diagnosis tasks.
-Specifically, without text control, the downstream diagnostic performance is comparatively poor.
-Fig.
-5(b-c) compares typical synthetic thyroid nodule sequences generated under condition banks with and without text guidance.
-The former is visibly high-fidelity and faithful to the given text prompt (e.g., smooth margin), while the latter exhibits less distinguishable features for diagnosis (e.g., blurry margin).
-That is to say, the proposed Ctrl-GenAug can create more diagnosis-reliable samples by enhancing semantic steerability in the generation process.
-Besides, by incorporating image prior knowledge in the sampling process, the average accuracy and AUROC of three classifiers are improved by 4.20% and 0.040, respectively.
-This proves the domain gap between synthetic and real samples is mitigated by introducing the image prior.
-Moreover, as shown in Table IV, conditioning the generator on the motion field produces more content-consistent and smoother samples, leading to continuous improvements in diagnostic performance.
-It can also be observed that the FVD results show a limited correlation with downstream evaluation metrics, confirming the finding in [7].
-TABLE IV: Ablation study for different conditional controls in TUSC [69], including class label (C), text (T), image prior (I), and motion field (MF).
-SlowFast [81] (C1), CSN [82] (C2), and FTC [78] (C3) trained on Joint-train paradigm were used for downstream diagnosis.
+Table 4: Ablation study for different conditional controls in TUSC (Stanford-AIMI, 2021), including class label (C), text (T), image prior (I), and motion field (MF).
+SlowFast (Feichtenhofer et al., 2019) (C1), CSN (Tran et al., 2019) (C2), and FTC (Ahmadi et al., 2023) (C3) trained on Joint-train paradigm were used for downstream diagnosis.
 “Hybrid”, the classifier with CNN-Transformer design.
@@
 […]
-TUSC [69] 4.18 90.40% 95.03% C1 CNN $\downarrow$ 1.80% $\downarrow$ 0.015
+TUSC (Stanford-AIMI, 2021) 4.18 90.40% 95.03% C1 CNN $\downarrow$ 1.80% $\downarrow$ 0.015
 C2 CNN $\uparrow$ 1.44% $\uparrow$ 0.022
@@
 C3 Hybrid $\uparrow$ 3.57% $\uparrow$ 0.030
-IV-F Impact of the Sequential Augmentation Module
-To verify the impact of the proposed sequential augmentation module, as shown in Table V, we compared the performance of three generator variants on the synthesis and downstream tasks.
-Ours-S, Ours-SK, and Ours-SKM denote our ablation studies, including gradually adding sequential attention (’-S’), key-frame/slice attention (’-K’), and motion field attention (’-M’) to the generator with spatial inflation only.
-Table V shows that the downstream accuracy and AUROC improvements of Ours-SKM are significantly higher than those of Ours-S and Ours-SK in both datasets.
-In terms of quantitative assessment of sequential coherence, Ours-SKM achieves better results overall than other variants.
-These prove that this module can help the generator synthesize cross-frame/slice consistent and dynamic-smooth clips, which are diagnosis-promotive for downstream diagnostic tasks.
-Moreover, it can be observed that the FVD results across these three generators show no significant difference, revealing its limited value for evaluating diagnosis-oriented medical sequence synthesis tasks.
-TABLE V: Ablation study for the sequential augmentation module of the generator in TUSC [69] and ACDC [70].
-SlowFast [81] (C1), CSN [82] (C2), and FTC [78] (C3) trained on Joint-train paradigm were used for downstream diagnosis.
+4.5 Analysis of Multimodal Conditions Guidance
+To illustrate the role of different conditional guidance, we conduct comparative experiments with generators that resort to various banks of conditions for training and sampling.
+Table 4 validates that all proposed conditions are effective for enhancing synthesis and diagnosis tasks.
+Specifically, without text control, the downstream diagnostic performance is comparatively poor.
+Fig.
+6(b-c) compares typical synthetic thyroid nodule sequences generated under condition banks with and without text guidance.
+The former is visibly high-fidelity and faithful to the given text prompt (e.g., smooth margin), while the latter exhibits less distinguishable features for diagnosis (e.g., blurry margin).
+That is to say, the proposed Ctrl-GenAug can create more diagnosis-reliable samples by enhancing semantic steerability in the generation process.
+Besides, by incorporating image prior knowledge in the sampling process, the average accuracy and AUROC of three classifiers are improved by 4.20% and 0.040, respectively.
+This proves the domain gap between synthetic and real samples is mitigated by introducing the image prior.
+Moreover, as shown in Table 4, conditioning the generator on the motion field produces more content-consistent and smoother samples, leading to continuous improvements in diagnostic performance.
+It can also be observed that the FVD results show a limited correlation with downstream evaluation metrics, confirming the finding in (Luo et al., 2024).
+Table 5: Ablation study for the sequential augmentation module of the generator in TUSC (Stanford-AIMI, 2021) and ACDC (Bernard et al., 2018).
+SlowFast (Feichtenhofer et al., 2019) (C1), CSN (Tran et al., 2019) (C2), and FTC (Ahmadi et al., 2023) (C3) trained on Joint-train paradigm were used for downstream diagnosis.
 “Hybrid”, the classifier with CNN-Transformer design.
@@
 […]
-TUSC [69] Ours-S 3.76 87.44% 91.07% C1 CNN $\downarrow$ 0.36% $\downarrow$ 0.009
+TUSC (Stanford-AIMI, 2021) Ours-S 3.76 87.44% 91.07% C1 CNN $\downarrow$ 0.36% $\downarrow$ 0.009
 C2 CNN $\uparrow$ 1.08% $\uparrow$ 0.011
@@
 C3 Hybrid $\uparrow$ 3.57% $\uparrow$ 0.030
-ACDC [70] Ours-S 8.75 70.11% 87.62% C1 CNN $\uparrow$ 0.73% $\uparrow$ 0.015
+ACDC (Bernard et al., 2018) Ours-S 8.75 70.11% 87.62% C1 CNN $\uparrow$ 0.73% $\uparrow$ 0.015
 C2 CNN $\downarrow$ 0.37% $\uparrow$ 0.000
@@
 C3 Hybrid $\uparrow$ 3.43% $\uparrow$ 0.029
+4.6 Impact of the Sequential Augmentation Module
+To verify the impact of the proposed sequential augmentation module, as shown in Table 5, we compared the performance of three generator variants on the synthesis and downstream tasks.
+Ours-S, Ours-SK, and Ours-SKM denote our ablation studies, including gradually adding sequential attention (’-S’), key-frame/slice attention (’-K’), and motion field attention (’-M’) to the generator with spatial inflation only.
+Table 5 shows that the downstream accuracy and AUROC improvements of Ours-SKM are significantly higher than those of Ours-S and Ours-SK in both datasets.
+In terms of quantitative assessment of sequential coherence, Ours-SKM achieves better results overall than other variants.
+These prove that this module can help the generator synthesize cross-frame/slice consistent and dynamic-smooth clips, which are diagnosis-promotive for downstream diagnostic tasks.
+Moreover, it can be observed that the FVD results across these three generators show no significant difference, revealing its limited value for evaluating diagnosis-oriented medical sequence synthesis tasks.
+4.7 Effectiveness of the Noisy Synthetic Data Filter
+As shown in Fig.
+8, we validated the indispensable role of our proposed noisy synthetic data filter by assessing its impact on downstream diagnostic accuracy, using two classifiers trained on Real-finetune and Joint-train paradigms across the carotid and ACDC datasets.
+For instance, on the carotid dataset, the Joint-train CSN resorting to non-quality controlled synthetic clips (i.e., ’None’), displays no significant difference with its Baseline (80.95% vs.
+80.27%), even an accuracy degradation for its VideoSwin counterpart (78.91% vs.
+79.59%).
+Then, by implementing our CF and SF strategy, diagnostic accuracy is significantly improved by 2.04% and 3.40%, respectively.
+Equipped with both, it achieves an accuracy of 87.07%, outperforming the ’None’ by 6.12%.
 Figure 8: Carotid stenosis and heart disease diagnostic accuracy (%) comparison using two classifiers on our filtering strategies under Real-finetune and Joint-train paradigms.
@@
 […]
-Generative Training Paradigm Augmentation Method Carotid TUSC [69] ACDC [70]
+Generative Training Paradigm Augmentation Method Carotid TUSC (Stanford-AIMI, 2021) ACDC (Bernard et al., 2018)
 Acc. $\Delta$ AUROC $\Delta$ Acc. $\Delta$ AUROC $\Delta$ Acc. $\Delta$ AUROC $\Delta$
@@
 Joint-train TraAug + GenAug 84.35 $\uparrow$ 5.44 0.811 $\uparrow$ 0.076 62.88 $\uparrow$ 2.52 0.578 $\uparrow$ 0.026 85.32 $\uparrow$ 4.40 0.762 $\uparrow$ 0.074
-IV-G Effectiveness of the Noisy Synthetic Data Filter
-As shown in Fig.
-7, we validated the indispensable role of our proposed noisy synthetic data filter by assessing its impact on downstream diagnostic accuracy, using two classifiers trained on Real-finetune and Joint-train paradigms across the carotid and ACDC datasets.
-For instance, on the carotid dataset, the Joint-train CSN resorting to non-quality controlled synthetic clips (i.e., ’None’), displays no significant difference with its Baseline (80.95% vs.
-80.27%), even an accuracy degradation for its VideoSwin counterpart (78.91% vs.
-79.59%).
-Then, by implementing our CF and SF strategy, diagnostic accuracy is significantly improved by 2.04% and 3.40%, respectively.
-Equipped with both, it achieves an accuracy of 87.07%, outperforming the ’None’ by 6.12%.
-IV-H Impacts of Traditional and Generative Augmentations
+4.8 Impacts of Traditional and Generative Augmentations
 We tested the contributions of different data augmentations in three downstream diagnostic tasks.
-The traditional augmentations used in this experiment can be found in Table I.
+The traditional augmentations used in this experiment can be found in Table 1.
 We report the performance gains of the two variants with data augmentations compared to the augmentation-free setup.
 For experimental fairness, we kept the training datasets for the None and TraAug variants the same size as the method combining traditional and generative augmentations.
-As presented in Table VI, the diagnostic performance shows noticeable improvements using the traditional augmentation method.
+As presented in Table 6, the diagnostic performance shows noticeable improvements using the traditional augmentation method.
 Further coupled with the generative augmentation, the results show more substantial improvements than TraAug.
 We also observe that in the case of using traditional augmentation, adding extra real training samples yields limited performance gains, even a drop for ACDC.
-These prove that generative augmentation can effectively supplement traditional data augmentation and provide intrinsically diverse data characteristics for enhancing deep classifier learning, which support the viewpoints of [7].
-V Conclusion and discussion
+These prove that generative augmentation can effectively supplement traditional data augmentation and provide intrinsically diverse data characteristics for enhancing deep classifier learning, which support the viewpoints of (Luo et al., 2024).
+5 Conclusion and discussion
 In this study, we present a new and general diffusion-based generative augmentation framework, named Ctrl-GenAug, to facilitate medical sequence classification by leveraging customized and diagnosis-reliable synthetic sequences.
@@
 This can better connect the synthesis task and the downstream one, thus further enhancing our sequence classification performance.
-Extensive experiments on 3 medical datasets across multiple organs and modalities, with 11 popular networks trained on 3 paradigms, show the effectiveness and generality of Ctrl-GenAug.
-Furthermore, we conduct extensive empirical analysis demonstrating that Ctrl-GenAug can be effectively leveraged to improve diagnostic performance in underrepresented high-risk populations and out-domain robustness.
+Extensive experiments on 3 datasets across multiple organs and modalities, with 11 popular networks trained on 3 paradigms, show the effectiveness and generality of Ctrl-GenAug.
+Furthermore, the empirical analysis validates that Ctrl-GenAug can be effectively leveraged to improve diagnostic performance in underrepresented high-risk populations and out-domain robustness.
 We believe that Ctrl-GenAug can serve as a robust and practical data augmentation tool for various clinical scenarios.
@@
 Second, in out-domain conditions, the proposed framework requires target domain data during generator training, which can be tough to obtain in advance in clinical practice.
-In future work, we will adopt fast-sampling strategies (e.g., AMED-Solver [88]) to better balance the trade-off between sampling time and sample quality.
-Besides, we will incorporate test-time adaptation methods [89, 90] to provide a more user-friendly tool for out-domain applications in clinical settings.
-Last, we will attempt to further extend the proposed method to segmentation and detection tasks.
+In future work, we will adopt fast-sampling strategies (e.g., AMED-Solver (Zhou et al., 2024b)) to better balance the trade-off between sampling time and sample quality.
+Besides, we will incorporate test-time adaptation methods (Huang et al., 2023b; Luo et al., 2023; Huang et al., 2023c; Zhang et al., 2024b) to provide a more user-friendly tool for out-domain applications in clinical settings.
+Last, we will attempt to further extend the proposed method to segmentation and detection tasks (Huang et al., 2025; Tao et al., 2025).
 Supplementary Material
-Figure S1: Additional synthetic results on Carotid (a-d), TUSC [69] (e-h), and ACDC [70] (i-l).
-We used the multimodal conditions joint training strategy [76] during training and flexibly dropped several during sampling.
+Figure S1: Additional synthetic results on Carotid (a-d), TUSC (Stanford-AIMI, 2021) (e-h), and ACDC (Bernard et al., 2018) (i-l).
+We used the multimodal conditions joint training strategy (Huang et al., 2023a) during training and flexibly dropped several during sampling.
 Table S1: A brief review of studies in the medical field utilizing diffusion-based generative models to promote downstream tasks.
@@
 Study Conditional Control Downstream Task Downstream Arch. Downstream Net Num. Medical Modality Data Type
-Shang et al. [49] disease class,
+(Shang et al., 2023) disease class,
 image readability
@@
 2 Fundus image
-Farooq et al. [50] text classification CNN-based,
+(Farooq et al., 2024) text classification CNN-based,
 Transformer-based
 2 Dermoscopy image
-Zhang et al. [51] text, edge segmentation CNN-Transformer hybrid-based 1 US, CT, MRI image
-Sagers et al. [91] text classification CNN-based 1 Dermoscopy image
-Khosravi et al. [92] attribute, 14 pathology labels classification CNN-based 1 X-ray image
-Yu et al. [93] pathology labels,
+(Zhang et al., 2024c) text, edge segmentation CNN-Transformer hybrid-based 1 US, CT, MRI image
+(Sagers et al., 2023) text classification CNN-based 1 Dermoscopy image
+(Khosravi et al., 2024) attribute, 14 pathology labels classification CNN-based 1 X-ray image
+(Yu et al., 2024a) pathology labels,
 lesion bounding boxes,
@@
 classification Transformer-based 1 US image
-Ktena et al. [25] disease class, attribute classification CNN-based 1 Histopathology,
+(Ktena et al., 2024) disease class, attribute classification CNN-based 1 Histopathology,
 X-ray,
@@
 image
-Luo et al. [7] disease class classification CNN-based 10 Fundus,
+(Luo et al., 2024) disease class classification CNN-based 10 Fundus,
 Dermoscopy,
@@
 image
-Li et al. [24] multi-scale latent
+(Li et al., 2024a) multi-scale latent
 representation
 classification Transformer-based 1 Endoscopy video
-Gong et al. [94] mask segmentation CNN-based 1 MRI volume
-Peng et al. [95] mask segmentation CNN-based, CNN-Transformer hybrid-based 3 CT volume
-Kebaili et al. [96] tumor attribute, mask segmentation CNN-based 1 MRI volume
-Yu et al. [97] mask segmentation CNN-based 1 CT volume
-Chen et al. [98] mask segmentation CNN-based,
+(Gong et al., 2024) mask segmentation CNN-based 1 MRI volume
+(Peng et al., 2024) mask segmentation CNN-based, CNN-Transformer hybrid-based 3 CT volume
+(Kebaili et al., 2024) tumor attribute, mask segmentation CNN-based 1 MRI volume
+(Yu et al., 2024b) mask segmentation CNN-based 1 CT volume
+(Chen et al., 2024) mask segmentation CNN-based,
 CNN-Transformer hybrid-based
 3 CT volume
-Dorjsembe et al. [99] mask segmentation CNN-based 1 MRI volume
-Huang et al. [100] mask segmentation CNN-based 2 OCT volume
+(Dorjsembe et al., 2024a) mask segmentation CNN-based 1 MRI volume
+(Huang et al., 2024a) mask segmentation CNN-based 2 OCT volume
 Ours disease class, text,

### v3

@@
-Ctrl-GenAug: Controllable Generative Augmentation for Medical Sequence Classification
+
 Abstract
@@
 Then, we propose a noisy synthetic data filter to suppress unreliable cases at the semantic and sequential levels.
-Extensive experiments on 3 medical datasets, using 11 networks trained on 3 paradigms, comprehensively analyze the effectiveness and generality of Ctrl-GenAug, particularly in underrepresented high-risk populations and out-domain conditions.
-keywords:
-Classification, Controllable Generative Augmentation, Medical Sequence Synthesis, Diffusion Model
-journal: ***
+Extensive experiments on 5 medical datasets with 4 different modalities, including comparisons with 15 augmentation methods and evaluations using 11 networks trained on 3 paradigms, comprehensively demonstrate the effectiveness and generality of Ctrl-GenAug, particularly with pronounced performance gains in underrepresented high-risk populations and out-domain conditions.
+Codes, models, and synthetic databases are available at https://github.com/XinRuiZhou0106/Ctrl-GenAug.
 1 Introduction
 Dynamic information in medical imaging (e.g., video sequences) plays a vital role in clinical diagnosis.
-Recent deep learning-based classifiers have shown the ability to improve the diagnostic accuracy of different diseases (Zhou et al., 2023; Wang et al., 2024c).
+Recent deep learning-based classifiers have shown the ability to improve the diagnostic accuracy of different diseases (Zhou et al., 2023; Wang et al., 2024b).
 Despite rapid advancements, the performance of current advanced solutions is still limited by several issues: 1) The scarcity of dynamic clinical cases, coupled with the high annotation cost, limits data availability; 2) Imbalanced data distribution, driven by the rarity of high-risk positive cases, skews model training; 3) Deep models are prone to brittle performance degradation when tested on out-domain data, e.g., cases from different medical centers (Varoquaux and Cheplygina, 2022; Huang et al., 2022).
@@
 (b) Thyroid US videos with different TI-RADS levels, where blue boxes indicate nodules.
-(c) Cardiac MRI volumes with distinct diseases.
+(c) Cardiac Cine-MRI volumes with distinct diseases.
 Three key anatomical structures associated with diagnosis are highlighted with masks, including the left ventricle (blue), myocardium (green), and right ventricle (red).
+(d) Lung CT volumes exhibiting varying degrees of tissue abnormalities with COVID-19.
+(e) Knee T2-weighted MRI volumes with and without anterior cruciate ligament (ACL) tears.
 Intuitively, building customized, high-fidelity, and sequentially coherent synthetic medical databases, and effectively utilizing them, is essential for enhancing sequence recognition.
@@
 1(a-b)) or structures (Fig.
-1(c)) of the same disease category exhibit large visual variances (e.g., shapes, intensities, positions, etc).
+1(c-e)) of the same disease category exhibit large visual variances (e.g., shapes, intensities, positions, etc).
 This may seriously confuse diffusion learning, thus causing uncontrollable and unreliable generation.
@@
 1(a-b)) and the varying sizes of anatomical regions (Fig.
-1(c)) challenge the coherence of synthesized sequences.
+1(c-e)) challenge the coherence of synthesized sequences.
 The fourth challenge lies in the domain gap between synthetic and real samples, where achieving real-domain customization is crucial for effectively utilizing synthetic samples in downstream learning (He et al., 2023).
@@
 This is a general method for enhancing medical sequence classification.
-We believe this is the first comprehensive study to analyze the impact of the controllable generative augmentation scheme on promoting multi-modal medical sequence classification.
+We believe this is the first comprehensive study to analyze the impact of controllable generative augmentation on multi-modal medical sequence classification.
 Our contributions are three-fold:
-  • 1.
   • We propose a multimodal conditions-guided sequence generator to ensure controllable and high-fidelity synthesis.
@@
 Moreover, we design a sequential augmentation module to promote dynamics modeling, thus improving the temporal or stereoscopic coherence of generated data.
-  • 2.
   • We introduce a noisy synthetic data filter to suppress harmful synthetic sequences at class semantics and sequential levels.
 By effectively reducing noise and enhancing the reliability of the synthetic databases, we can associate the synthetic samples with downstream tasks, thus better improving the classification performance.
-  • 3.
-    Comprehensive experiments on 3 medical datasets, with 11 downstream networks trained on 3 paradigms, show that our proposed framework can improve diagnostic performance, especially within underrepresented high-risk sets.
-Ctrl-GenAug also increases out-domain robustness, further indicating its practicality in clinical scenarios.
+  • Comprehensive experiments on 5 medical datasets with 4 modalities (US, CT, Cine/T2-weighted MRI), including benchmarking against 15 augmentation methods and evaluations using 11 downstream classifiers trained on 3 paradigms, show that our approach consistently achieves superior performance across diverse diagnostic tasks, while maintaining robust effectiveness with multiple classifier architectures.
+Ctrl-GenAug also improves diagnostic performance in underrepresented high-risk sets and enhances out-domain robustness, underscoring its practicality in real clinical scenarios.
 It is noted that the sequence generator extends our prior work presented at MICCAI (Zhou et al., 2024a), with significant improvements in controllability at both semantic and sequential levels.
@@
 Specifically, AnimateDiff (Guo et al., 2023) and MoonShot (Zhang et al., 2024a) introduced text and image conditions to control the visual appearance.
-Most recently, several studies (Zhang et al., 2023; Zhao et al., 2023; Wang et al., 2024b) achieved fine-grained spatial compositional control by integrating additional diverse conditions (i.e., sketch, mask, and depth sequences).
-Meanwhile, two spatial-sequential (S2) modeling strategies (fully attention layers (Zhang et al., 2023) and S2 condition encoder (Wang et al., 2024b)) were introduced to capture the dynamics of sequential conditions, promoting the inter-frame features interaction.
+Most recently, several studies (Zhang et al., 2023; Zhao et al., 2023; Wang et al., 2023) achieved fine-grained spatial compositional control by integrating additional diverse conditions (i.e., sketch, mask, and depth sequences).
+Meanwhile, two spatial-sequential (S2) modeling strategies (fully attention layers (Zhang et al., 2023) and S2 condition encoder (Wang et al., 2023)) were introduced to capture the dynamics of sequential conditions, promoting the inter-frame features interaction.
 Besides, dense optical flow (Zach et al., 2007) was adopted to control sequential consistency (Liang et al., 2023; Ni et al., 2023).
@@
 Hence, they may not suit the medical video synthesis tasks.
+Figure 2: t-SNE visualization of features of synthetic and real training data by a pre-trained I3D (Carreira and Zisserman, 2017) model on our five datasets.
+Our Ctrl-GenAug produces synthetic samples with diverse feature patterns that enrich the training distribution.
 In the medical field, several diffusion-based methods have been proposed to generate photorealistic echocardiography videos using the conditional guidance of a single frame (Reynaud et al., 2023), ejection fractions (Reynaud et al., 2023), single semantic mask (Van Phi et al., 2024; Nguyen et al., 2024), and single sketch image (Zhou et al., 2024a).
 Zhou et al.
-(Zhou et al., 2024a) additionally leveraged mitral valve (MV) skeletons to control the complex motion trajectories of MV in the generated echocardiograms.
+(2024a) additionally leveraged mitral valve (MV) skeletons to control the complex motion trajectories of MV in the generated echocardiograms.
 Besides, Li et al.
-(Li et al., 2024a) proposed to generate endoscopy videos that simulate clinical scenes by integrating a spatial-temporal transformer with 2D vision foundation model priors.
+(2024a) proposed to generate endoscopy videos that simulate clinical scenes by integrating a spatial-temporal transformer with 2D vision foundation model priors.
 However, the above methods may have the following drawbacks: unsteerable semantics due to the lack of attribute guidance (e.g., category or target shape), poor control over sequences, etc.
 Thus, they may be unsuitable for direct adoption in the generative augmentation tasks.
-Figure 2: t-SNE visualization of features of synthetic and real training data by a pre-trained I3D (Carreira and Zisserman, 2017) model on our three datasets.
 2.2 Generative Augmentation with Diffusion Models
@@
 Recently, Singh et al.
-(Singh et al., 2024) performed an in-depth analysis of models trained with synthetic data across various robustness measures, and verified that they achieved good suitability in real-world settings.
+(2024) performed an in-depth analysis of models trained with synthetic data across various robustness measures, and verified that they achieved good suitability in real-world settings.
 To address data scarcity and promote diagnostic performance, the intelligent medical field has raised increasing interest in exploring diffusion-based generative augmentation for aiding different downstream diagnostic tasks (Shang et al., 2023; Farooq et al., 2024; Ktena et al., 2024; Zhang et al., 2024c; Dorjsembe et al., 2024b; Li et al., 2024b).
 Luo et al.
-(Luo et al., 2024) marked the first focus on comprehensively analyzing the impact of their designed uncertainty-guided diffusion models on downstream diagnostic tasks.
+(2024) marked the first focus on comprehensively analyzing the impact of their designed uncertainty-guided diffusion models on downstream diagnostic tasks.
 Nevertheless, current studies have two main limitations:
@@
 3 shows the pipeline of using our proposed controllable generative augmentation framework to aid medical sequence classification.
-First, we propose a sequence generator that enables perceiving multiple semantic and sequential conditions to guide controllable and diverse generation.
-Then, an efficient noisy data filter is introduced to suppress unsatisfactory synthetic sequences.
-Last, the quality-controlled synthetic sequences and real ones will work together to improve the performance of arbitrary classifiers.
+Our augmentation system is composed of two key components, including a sequence generator and a noisy synthetic data filter.
+First, the sequence generator perceives multiple semantic (including class label, descriptive text, and image prior) and sequential conditions (i.e., motion field) to produce customized and diverse synthetic databases.
+Then, the noisy synthetic data filter is employed to effectively suppress unsatisfactory generated sequences, resulting in reliability-enhanced and coherent synthetic databases for downstream diagnosis.
+In the application stage, the quality-controlled synthetic sequences and real ones work together to improve the performance of arbitrary classifiers.
 To formulate customized, high-fidelity, and coherent synthetic databases to boost classification, we design the whole framework from model-centric, i.e., sequence generator (Secs.
-3.23.3, and 3.4) and data-centric, i.e., data filter (Sec.
+3.2, 3.3, and 3.4) and data-centric, i.e., data filter (Sec.
 3.5) perspectives.
@@
 During inference, given a single or combination of multimodal conditions as control signal inputs, high-quality and steerable sequence generation from Gaussian noise can work.
-We then provide a short background of video diffusion models in Sec.
-3.2.1, followed by a detailed description of the basic architecture of our generator in Secs.
-3.2.2 and 3.2.3.
-3.2.1 Preliminaries of Video DDPMs
+We then provide a short background of video diffusion models, followed by a detailed description of the basic architecture of our generator.
+Preliminaries of Video DDPMs.
 Video DDPMs aim to learn complex distributions by iteratively denoising corrupted inputs using diffusion methods (Sohl-Dickstein et al., 2015; Ho et al., 2020).
@@
 Mathematically, the optimized objective can be a simplified variant of the variational lower bound:
-$\begin{split}\min_{\theta}\mathbb{E}_{\mathit{z}_{0},\epsilon\sim\mathcal{N}(\mathbf{0},\mathbf{I}),\mathbf{c},\mathit{t}}||\epsilon-\epsilon_{\theta}(\mathit{z}_{t},\mathbf{c},\mathit{t})||^{2},\end{split}$ (1)
+[…]
 where $\mathit{z}_{0}$ is the latent code of $\mathit{x}_{0}$.
@@
 For our generator, $\mathbf{c}$ involves multimodal conditions including semantic and sequential ones.
-3.2.2 Factorized Learning of Visual Features and Sequential Patterns
+Factorized Learning of Visual Features and Sequential Patterns.
 Developing a generator with pure 3D architecture design is an intuitive scheme for sequence synthesis (Blattmann et al., 2023).
@@
 In this way, the generator enables realistic sequence synthesis, while easing the training burden.
-3.2.3 2D-to-3D Model Inflation
+2D-to-3D Model Inflation.
 During pretraining, we develop a 2D UNet (Ronneberger et al., 2015) in LDM to predict noises for image synthesis.
@@
 Overall, sequence LDM enables inheriting the rich visual concepts preserved in LDM and focusing on sequential pattern aggregation, making the model learning efficient.
-Figure 4: Pipeline of our proposed sequence generator. The MedSAM image encoder (Huang et al., 2024b) is used for domain-specific image prior feature extraction.
+Figure 4: Pipeline of our proposed sequence generator.
+The MedSAM image encoder (Huang et al., 2024b) is used for domain-specific image prior feature extraction.
+During inference, given a bank of customized conditions (e.g., descriptive text with underlined carotid plaque attributes, the class label mild, image prior, and motion field), our generator produces a real-domain style sequence that faithfully adheres to all specified conditions.
 3.3 Multimodal Conditions Guidance
@@
 Details are described below.
-3.3.1 Semantic Conditions for Visual Appearance Control
-As shown in Fig. 4, the sequence generator exploits three semantic conditions to perform visual appearance control, thus achieving controllable and high-fidelity synthesis.
+Semantic Conditions for Visual Appearance Control.
+As shown in Fig.
+4, the sequence generator exploits three semantic conditions to perform visual appearance control, thus achieving controllable and high-fidelity synthesis.
 - Descriptive Text: It provides an intuitive indication of the coarse-grained semantic concepts of sequences.
-In our implementation, we pre-align the visual and textual features by finetuning the widely used CLIP model (Radford et al., 2021) using our medical datasets to mitigate the large gap between them.
+In our implementation, we pre-align the visual and textual features by finetuning the widely used CLIP model (Radford and others, 2021) using our medical datasets to mitigate the large gap between them.
 Then, semantic information of texts can feasibly query relevant medical visual patterns, thus easing subsequent training.
@@
 The decoupled attention can be formulated as:
-[…]
+$\left\{\begin{aligned} &\mathbf{Q_{g}}=W^{Q_{g}}f_{g},\mathbf{K_{g}}=W^{K_{g}}f_{g},\mathbf{V_{g}}=W^{V_{g}}f_{g},g\in\left\{\mathbf{t},\mathbf{i}\right\},\\ &Attention\left(\mathbf{Q_{g}},\mathbf{K_{g}},\mathbf{V_{g}}\right)=\sum_{g}\left[\sigma(\frac{\mathbf{Q_{g}K_{g}^{\top}}}{\sqrt{d}}\mathbf{V_{g}})\right],\end{aligned}\right$.
+(2)
 where $W^{Q},W^{K},W^{V}$ denote the trainable linear projection matrices.
@@
 $\sigma(\cdot)$, softmax function.
-3.3.2 Sequential Condition for Serial Guidance
+Sequential Condition for Serial Guidance.
 Apart from perceiving visual concepts, modeling sequential knowledge is also important in sequence synthesis, with sequential cues playing a vital role.
 Most existing methods (Liang et al., 2023; Ni et al., 2023) adopted dense optical flow (Zach et al., 2007) to promote sequential dynamics modeling.
-However, optical flow extraction requires high computational demands (Wang et al., 2024b).
+However, optical flow extraction requires high computational demands (Wang et al., 2023).
 Thus, we instead introduce computation-efficient motion field as a sequential condition in sequence LDM.
@@
 5(d), SAM integrates two attention mechanisms in cascade for sequential augmentation.
-3.4.1 Key-frame/slice Attention
+Key-frame/slice Attention.
 The common spatial self-attention layers can lead to sequential inconsistency due to the lack of interaction across frames/slices.
@@
 $\begin{split}\mathbf{Q}=W^{Q}z_{l},\mathbf{K}=W^{K}\left\{z_{1},z_{l-1}\right\},\mathbf{V}=W^{V}\left\{z_{1},z_{l-1}\right\},\end{split}$ (3)
-[…]
-3.4.2 Motion Field Attention
+where $W^{Q},W^{K},W^{V}$ are initialized on the original spatial self-attention weights for inheriting the semantic perception capability of LDM in the finetuning stage.
+$\left\{\cdot\right\}$ represents concatenation operation.
+It is highlighted that KA retains low computational complexity compared with full attention (Zhang et al., 2023).
+Please refer to Fig.
+5(b) for a visual illustration.
+Motion Field Attention.
 To further boost the sequential coherence of the generated sequences, we propose to reuse the motion field condition (refer to Sec.
-3.3.2) by introducing a motion field attention (MFA) mechanism after KA.
+3.3) by introducing a motion field attention (MFA) mechanism after KA.
 As shown in Fig.
@@
 In step 1, we sample the patch pathways based on the m-scaled downsampled motion fields.
-Take a video sequence as an example, for a patch $p_{l}$ on the l-th frame of a f-frame video, the $path$ can be derived from the motion field.
+Take a video sequence as an example, for a patch $p_{l}$ on the l-th frame of a f-frame video, the pathpath can be derived from the motion field.
 Since the sampling procedure inevitably generates multiple pathways for the same patch, we randomly sample a pathway to ensure its uniqueness to which each patch belongs.
@@
 $\begin{split}\mathbf{Q}=z_{p_{l}},\mathbf{K}=\mathbf{V}=z_{path},\end{split}$ (4)
-where $z_{path}$ represents the latent features of patches on the $path$, as follows:
+where $z_{path}$ represents the latent features of patches on the pathpath, as follows:
 $\begin{split}z_{path}=\left[z_{p_{0}},z_{p_{1}},\dots,z_{p_{l}},\dots,z_{p_{f-1}},z_{p_{f}}\right].\end{split}$ (5)
@@
 Algorithm 1 Noisy synthetic data filtering.
-1:synthetic samples set $\mathrm{S}$, real data-trained classifier $p$.
+1:synthetic samples set $\mathrm{S}$, real data-trained classifier pp.
 2:# Stage 1: Class semantics misalignment filtering
-3:$\mathrm{S_{1}}=\varnothing$triangleright synthetic samples set obtained after stage 1
+[…]
 4:for $i\in\left\{1,...,n\right\}$ do
 5:  $c\leftarrow$ class in the conditions bank of the group $\mathrm{S}[i]$
-6:    $l_{x}=-\sum c\mathrm{log}\sigma(p(\mathrm{S}[i][x]))$triangleright loss of the sample indexed $x$
-7:    $L_{c}=\frac{1}{M}\sum_{x=1}^{M}l_{x}$triangleright noise identification threshold
-8$\triangleright$ $M$ clips in the group
+6:  $l_{x}=-\sum c\mathrm{log}\sigma(p(\mathrm{S}[i][x]))$ $\triangleright$ loss of the sample indexed $x$
+7:  $L_{c}=\frac{1}{M}\sum_{x=1}^{M}l_{x}$ $\triangleright$ noise identification threshold
+[…]
 9:  for $x\in\left\{1,...,M\right\}$ do
 10:    if $l_{x}>L_{c}$ then
-11:             $continue$
+11:     continuecontinue
 12:    else
-13:             $\mathrm{S_{1}}\leftarrow\mathrm{S_{1}}\cup\{\mathrm{S}[i][x]\}$triangleright save less noisy sample
+13:     $\mathrm{S_{1}}\leftarrow\mathrm{S_{1}}\cup\{\mathrm{S}[i][x]\}$ $\triangleright$ save less noisy sample
 14:    end if
@@
 16:end for
-17:return $\mathrm{S_{1}}$triangleright with $n_{1}(<n)$ groups totaling $N_{1}(<N)$ clips
+17:return $\mathrm{S_{1}}$ $\triangleright$ with $n_{1}(<n)$ groups totaling $N_{1}(<N)$ clips
 18:# Stage 2: Inner-sequence filtering
-19:$\mathrm{S_{2}}=\varnothing$triangleright synthetic samples set obtained after stage 2
-20:$\mathrm{A}=\left\{A_{1},...,A_{N}\right\}$triangleright VAE-Seq values set of $\mathrm{S}$
-21:$\mathrm{B}=\left\{B_{1},...,B_{N_{1}}\right\}$triangleright VAE-Seq values set of $\mathrm{S_{1}}$
-22:$t_{l},t_{h}\leftarrow\mathbf{KMeans}(\mathrm{A},\mathbf{K}=4)$triangleright boundary thresholds
+[…]
+[…]
+[…]
+[…]
 23:for $j\in\left\{1,...,N_{1}\right\}$ do
 24:  if $B_{j}\geq t_{l}\And B_{j}\leq t_{h}$ then
-25:         $\mathrm{S_{2}}\leftarrow\mathrm{S_{2}}\cup\{y_{j}|y_{j}\in\mathrm{S_{1}}\}$triangleright save gentle dynamic sample
+25:    $\mathrm{S_{2}}\leftarrow\mathrm{S_{2}}\cup\{y_{j}|y_{j}\in\mathrm{S_{1}}\}$ $\triangleright$ save gentle dynamic sample
 26:  else
-27:         $continue$
+27:    continuecontinue
 28:  end if
 29:end for
-30:return $\mathrm{S_{2}}$triangleright with $n_{2}(<n_{1})$ groups totaling $N_{2}(<N_{1})$ clips
+30:return $\mathrm{S_{2}}$ $\triangleright$ with $n_{2}(<n_{1})$ groups totaling $N_{2}(<N_{1})$ clips
 31:# Stage 3: Inter-sequence filtering
-32:$\mathrm{S_{3}}=\varnothing$triangleright synthetic samples set obtained after stage 3
+[…]
 33:for $k\in\left\{1,...,n_{2}\right\}$ do
@@
 […]
-37:             $\mathrm{S_{3}}\leftarrow\mathrm{S_{3}}\cup\mathrm{S_{2}}[k][q]$triangleright save diverse sample
+37:     $\mathrm{S_{3}}\leftarrow\mathrm{S_{3}}\cup\mathrm{S_{2}}[k][q]$ $\triangleright$ save diverse sample
 38:    else
-39:             $continue$
+39:     continuecontinue
 40:    end if
@@
 42:end for
-43:return $\mathrm{S_{3}}$triangleright final synthetic database for downstream training
+43:return $\mathrm{S_{3}}$ $\triangleright$ final synthetic database for downstream training
 3.5 Noisy Synthetic Data Filter
@@
 For instance, in Fig.
-6(e), the synthesized carotid clip category is wrong, which should be moderate rather than mild.
+6(b), the synthesized carotid clip category is wrong, which should be moderate rather than mild.
 In Fig.
-6(f), the generated clip includes abrupt changes in anatomical structures.
+6(f, h, j), the generated clip includes abrupt changes in anatomical structures.
 Hence, blindly using all synthetic clips for classifier learning will significantly cause a performance drop due to noisy samples.
@@
 This can also link the synthetic data to downstream tasks, thus potentially achieving a higher performance upper bound.
-3.5.1 Class Semantics Misalignment Filtering
+Figure 6: Typical good and bad synthetic samples of five datasets using different banks of conditions, with representative failure patterns highlighted.
+Conditional controls include class label (C), text (T), image prior (I), and motion field (MF).
+See more synthetic results in Supplementary Material.
+Class Semantics Misalignment Filtering.
 This scheme is designed to filter the generated clips whose visual contents mismatch the corresponding class labels.
@@
 If the loss of a synthetic clip with $c$ exceeds the threshold $L_{c}$, it will be viewed as a noisy sample and excluded from the synthetic pool.
-3.5.2 Sequential Filtering
-This strategy aims to filter noisy samples at a) inner-sequence and b) inter-sequence levels.
-For a), we screen out synthetic clips with gentle dynamic, avoiding those that are either inconsistent or over-consistent affecting downstream learning.
-Specifically, we retain a synthetic clip whose cross-frame/slice consistency falls in a pre-computed range.
-It is determined by K-means clustering (Macqueen, 1967) based on the cross-frame/slice consistency of all synthetic clips.
-To assess the video sequence consistency, (Wang et al., 2024b) tends to use CLIP (Radford et al., 2021) image embeddings to compute the average cosine similarity across consecutive frames.
-However, this method leads to inadequate evaluation due to the usage of limited informative CLIP embeddings.
-To resolve the problem, we propose a metric termed VAE-Seq that utilizes latent embeddings from pretrained VAE, instead of deriving embeddings from CLIP space, to assess cross-frame/slice consistency.
-Therefore, with the higher dimension of latents (e.g., 4$\times$32$\times$32 for input size 256$\times$256 compared to 768 in CLIP embeddings), VAE-Seq reflects an accurate evaluation using more fine-grained and informative features.
-For b), we seek to diversify the samples in the synthetic set to prevent overfitting and avoid wasting computational resources in downstream training.
-To this end, we perform the inter-sequence filtering based on inter-clip similarity in each group, as shown in Algorithm 1.
-We quantify the inter-clip similarity by calculating the cosine distance of VAE latent embeddings between frame/slice pairs in two clips and averaging the obtained distance values.
-3.6 Evaluation
-Existing sequence synthesis works evaluate the synthetic quality by mostly using Fréchet Inception Distance (FID) (Heusel et al., 2017) and Fréchet Video Distance (FVD) (Unterthiner et al., 2018).
-However, researchers have verified that these metrics do not consistently correlate with performance metrics on downstream tasks (Yang et al., 2024; Luo et al., 2024).
-To exhaustively evaluate the values of synthetic cases to downstream tasks, 11 popular classifiers are deployed for comparison with respect to diagnostic accuracy and area under the receiver operating characteristic curve (AUROC).
-Besides, we adopt the proposed VAE-Seq to assess the cross-frame/slice consistency (more details refer to Sec.
-3.5.2).
-Moreover, motivated by (Huang et al., 2024c), we quantify the smoothness of generated clips using the metric termed Dynamic Smoothness, which is the mean absolute error between the reconstructed frames/slices from the interpolation model (Li et al., 2023) and the original real ones.
-Figure 6: Typical good and bad synthetic samples of three datasets using different banks of conditions.
-Conditional controls include class label (C), text (T), image prior (I), and motion field (MF).
-See more synthetic results in Supplementary Material.
+In our experiments, the classifier employed for semantic filtering is identical to the downstream classifier by default.
+[…]
 4 Experiments
-We first comprehensively validated the effectiveness and generality of our Ctrl-GenAug on 3 medical datasets, using 11 popular networks trained on 3 paradigms.
-We then tested the performance of Ctrl-GenAug in underrepresented high-risk sets and out-domain conditions.
-Last, we analyzed the contributions of our multimodal conditions guidance, sequential augmentation module, and noisy synthetic data filter, respectively; as well as the impacts of traditional and generative augmentations.
+We first benchmarked Ctrl-GenAug against 15 augmentation methods across 5 datasets with 4 medical modalities to comprehensively validate its effectiveness and generality.
+We further systematically assessed its downstream robustness using 11 popular classifiers trained under 3 paradigms.
+We then evaluated its performance in underrepresented high-risk sets and out-domain conditions.
+Finally, we conducted extensive ablation studies to examine the contributions of each component and the behavior of Ctrl-GenAug under different diffusion settings, backbone selections in the semantics misalignment filter, and varying annotation requirements.
+Table 1: Experimental datasets and settings.
+Carotid TUSC ACDC MosMedData MRNet
+Modality US US Cine-MRI CT T2-weighted MRI
+Organ Carotid artery Thyroid Heart Lung Knee
+Patient 231 167 150 1,110 1,199
+Clip 486 633 314 2,573 1,958
+Mild/Moderate/Severe TI-RADS2-3/4/5 NOR/MINF/DCM/HCM/ARV Normal/Mild/Severe Negative/Positive
+Class (train) 141/104/64 146/189/113 40/40/40/40/45 512/1,226/317 1,449/320
+Class (test) 51/27/20 64/73/48 21/20/20/20/28 127/314/77 109/80
+Class (OD) 37/24/18 - - - -
+Sequence generator Training
+Downsampling rate of VAE(8)   $\mathit{T}$(1,000)   AdamW(lr=1e-4)   lr scheduler(Cosine)
+Warm(500 iterations)   Epochs(200)   Pretraining batch size(64)   Finetuning batch size(8)
+Inference
+Classifier-free guidance (Ho and Salimans, 2022) factor(7.5)   Sampling steps(200)
+Classifiers Adam(default lr=1e-4, unless specified)   Batch size(default=8, unless specified)   lr scheduler(Cosine)
+Epochs(100)   Standard augmentation practices(Random Color/Move/Gaussian/Rotation/Flip)
 4.1 Datasets and Settings
@@
 Text prompts were set to empty strings during generator training.
+MosMedData (Vladzymyrskyy et al., 2020) contains 3D lung CT sequences from 1,110 patients collected by municipal hospitals in Moscow.
+Each sequence was labeled according to the severity of lung tissue abnormalities with COVID-19, originally graded from CT-0 to CT-4 (CT-0: normal, CT-1: mild, CT-2/CT-3/CT-4: increasing severity).
+For analysis, CT-2, CT-3, and CT-4 were merged into a single severe category to alleviate class imbalance.
+We uniformly sampled the sequences to obtain 2,573 8-slice clips, of which 2,055 were used for training and 518 for testing.
+MRNet (Bien et al., 2018) publicly releases 1,250 knee T2-weighted MRI exams, each comprising three orthogonal planes, from 1,199 patients acquired at the Stanford University Medical Center.
+Given the complexity of handling multiple imaging planes simultaneously, as suggested in (Khandelwal et al., 2018), we only included sagittal-plane sequences in the analysis for diagnosing anterior cruciate ligament (ACL) tears.
+Each sequence was then uniformly sampled into several 8-slice clips.
+The final dataset was split following the official protocol, resulting in 1,769 training clips and 189 testing clips.
 Table 1 presents a detailed description of experimental datasets and settings.
-Using all available conditions, we synthesized 300 clips per category for the ACDC dataset and 500 clips per class for others.
-All frames/slices were resized to 256$\times$256.
+Using all available conditions, we synthesized 300 clips per category for ACDC, 500 clips per class for TUSC and Carotid, 1,500 clips per category for MosMedData, and 2,000 clips per class for MRNet.
+All frames/slices were resized to $\times256$.
 By default, we set m=8 for the MFA mechanism.
@@
 This allows users to flexibly select any single condition or combination during sampling, without finetuning the model for each specific combination.
+All datasets were employed to benchmark the proposed Ctrl-GenAug against a wide range of augmentation methods and to comprehensively examine its effectiveness across multiple classifiers.
 Motivated by (Yang et al., 2024), we adopted three paradigms to explore the impacts of the synthetic samples for downstream classifiers.
@@
 Joint-train: real clips were first over-sampled to match the number of synthetic clips, the classifier was then trained on the real-synthetic mixed set.
-All classification experiments were conducted on MMAction2 (Contributors, 2020) except for FTC (Ahmadi et al., 2023) and CSG-3DCT (Zhou et al., 2023).
-Table 1: Experimental datasets and settings. CF, Classifier-free.
-Carotid TUSC ACDC
-Patient 231 167 150
-Clip 486 633 314
-Mild/Moderate/Severe TI-RADS2-3/4/5 NOR/MINF/DCM/HCM/ARV
-Class (train) 141/104/64 146/189/113 40/40/40/40/45
-Class (test) 51/27/20 64/73/48 21/20/20/20/28
-Class (OD) 37/24/18 - -
-Sequence generator Training
-Downsampling rate of VAE(8)   $\mathit{T}$(1,000)
-AdamW(lr=1e-4)   lr scheduler(Cosine)   Warm(500 iterations)
-Epochs(200)   Pretraining batch size(64)   Finetuning batch size(8)
-Inference
-CF guidance (Ho and Salimans, 2022) factor(7.5)   Sampling steps(200)
-Eleven classifiers Adam(lr=1e-4)   lr scheduler(Cosine)   Epochs(100)   Batch size(8)
-Traditional augmentations(Random Color/Move/Gaussian/Rotation/Flip)
-[…]
-Method Backbone Training Paradigm Carotid TUSC (Stanford-AIMI, 2021) ACDC (Bernard et al., 2018)
+For classification experiments on the MosMedData and MRNet datasets, we set the batch size to 16 and the initial learning rate to 2e-4.
+All diagnostic experiments were conducted on MMAction2 (Contributors, 2020) except for FTC (Ahmadi et al., 2023) and CSG-3DCT (Zhou et al., 2023).
+Table 2: Diagnostic performance comparison of diverse augmentation methods using I3D (Carreira and Zisserman, 2017) on multi-organ and multi-modal datasets.
+Classifiers with GAN- and diffusion-based augmentations were trained on Real-finetune paradigm for diagnosis.
+Underlined results show no significant difference compared with Ours, all others indicate significant differences ($p<0.05$).
+95% confidence intervals for both metrics are reported.
+The best results are shown in bold.
+Acc., accuracy ($\%$).
+Augmentation Method Carotid TUSC ACDC MosMedData MRNet
+Acc. AUROC Acc. AUROC Acc. AUROC Acc. AUROC Acc. AUROC
+Traditional Augmentation
+Basic 79.59 [73.47, 85.71] 0.737 [0.660, 0.815] 60.72 [56.04, 65.41] 0.559 [0.506, 0.613] 83.49 [79.82, 87.17] 0.733 [0.678, 0.791] 67.18 [64.35, 70.01] 0.566 [0.535, 0.597] 63.49 [56.08, 70.37] 0.592 [0.532, 0.647]
+Random Elastic Deformation 80.27 [74.15, 85.71] 0.747 [0.672, 0.817] 60.72 [56.04, 65.41] 0.554 [0.502, 0.610] 83.12 [79.45, 86.79] 0.725 [0.671, 0.781] 68.08 [65.25, 70.91] 0.584 [0.553, 0.615] 66.67 [59.26, 73.54] 0.623 [0.563, 0.680]
+Average 79.93 0.742 60.72 0.557 83.31 0.729 67.63 0.575 65.08 0.608
+Composition-based Augmentation
+AutoAugment (Cubuk et al., 2019) 81.63 [75.62, 86.37] 0.740 [0.682, 0.823] 61.08 [57.84, 66.52] 0.567 [0.512, 0.624] 85.68 [82.26, 89.74] 0.756 [0.702, 0.800] 68.72 [65.84, 71.13] 0.598 [0.569, 0.630] 66.13 [58.25, 72.10] 0.611 [0.553, 0.621]
+AdaAug (Cheung and Yeung, 2021) 82.99 [76.54, 87.21] 0.763 [0.712, 0.841] 60.36 [56.31, 65.83] 0.549 [0.489, 0.603] 84.59 [81.35, 89.05] 0.741 [0.685, 0.792] 69.37 [66.20, 71.85] 0.590 [0.558, 0.624] 66.67 [59.26, 72.43] 0.604 [0.540, 0.619]
+GRA-Video (An et al., 2022) 80.95 [76.19, 85.71] 0.722 [0.663, 0.798] 61.44 [58.00, 67.34] 0.556 [0.508, 0.618] 84.22 [80.64, 88.12] 0.747 [0.689, 0.784] 70.79 [67.59, 72.20] 0.584 [0.551, 0.619] 66.13 [58.87, 72.58] 0.617 [0.569, 0.630]
+Sample-aware RandAugment (Xiao et al., 2025) 82.31 [76.19, 87.76] 0.758 [0.700, 0.834] 61.80 [58.85, 67.76] 0.587 [0.529, 0.631] 83.85 [79.98, 87.25] 0.739 [0.689, 0.790] 70.27 [67.18, 71.94] 0.593 [0.562, 0.627] 66.67 [59.26, 72.43] 0.625 [0.559, 0.674]
+Average 81.97 0.746 61.17 0.565 84.59 0.746 69.79 0.591 66.40 0.614
+Mixing-based Augmentation
+Mixup (Zhang et al., 2017) 83.67 [77.51, 88.20] 0.772 [0.715, 0.839] 62.52 [57.84, 67.57] 0.591 [0.534, 0.648] 84.95 [81.10, 89.23] 0.759 [0.707, 0.805] 71.17 [68.05, 72.83] 0.589 [0.551, 0.614] 67.72 [59.99, 74.06] 0.631 [0.575, 0.685]
+CutMix (Yun et al., 2019) 82.31 [76.19, 87.76] 0.765 [0.695, 0.833] 62.52 [57.84, 67.57] 0.581 [0.526, 0.638] 83.12 [79.45, 86.79] 0.726 [0.676, 0.778] 69.24 [66.41, 72.20] 0.555 [0.524, 0.587] 68.25 [61.38, 74.60] 0.635 [0.579, 0.688]
+VideoMix (Yun et al., 2020) 83.67 [77.51, 88.20] 0.779 [0.723, 0.860] 63.24 [60.00, 68.00] 0.571 [0.514, 0.626] 85.32 [82.00, 89.51] 0.766 [0.718, 0.820] 71.04 [68.00, 73.58] 0.573 [0.536, 0.607] 69.31 [63.02, 75.73] 0.657 [0.595, 0.719]
+LayerMix (Ahmad et al., 2025) 84.35 [78.23, 89.80] 0.792 [0.720, 0.858] 61.44 [56.76, 66.13] 0.570 [0.514, 0.621] 83.85 [80.18, 87.52] 0.740 [0.686, 0.796] 71.17 [68.21, 74.00] 0.567 [0.536, 0.600] 69.84 [63.48, 76.19] 0.662 [0.601, 0.721]
+Average 83.50 0.777 62.43 0.578 84.31 0.748 70.66 0.571 68.78 0.646
+GAN-based Augmentation
+TGANv2 (Saito et al., 2020) 80.95 [74.83, 87.07] 0.728 [0.675, 0.803] 60.72 [56.28, 64.90] 0.564 [0.508, 0.621] 85.68 [82.26, 89.74] 0.760 [0.703, 0.805] 71.30 [68.84, 73.12] 0.587 [0.553, 0.619] 68.25 [61.29, 74.65] 0.626 [0.566, 0.681]
+StyleGAN-V (Skorokhodov et al., 2022) 82.31 [76.19, 87.76] 0.749 [0.694, 0.822] 61.80 [58.85, 67.34] 0.593 [0.536, 0.645] 85.32 [81.35, 89.74] 0.768 [0.721, 0.823] 71.43 [69.22, 73.65] 0.592 [0.555, 0.612] 68.25 [61.19, 74.62] 0.633 [0.578, 0.688]
+Average 81.63 0.739 61.26 0.579 85.50 0.764 71.37 0.590 68.25 0.630
+Diffusion-based Augmentation
+VideoComposer (Wang et al., 2023) 80.95 [74.83, 87.07] 0.764 [0.697, 0.833] 63.60 [58.92, 68.29] 0.590 [0.538, 0.645] 85.32 [81.65, 88.99] 0.761 [0.710, 0.815] 71.56 [68.73, 74.39] 0.591 [0.559, 0.626] 68.78 [61.90, 75.13] 0.653 [0.588, 0.712]
+Endora (Li et al., 2024a) 82.31 [76.19, 87.76] 0.772 [0.700, 0.842] 64.68 [60.00, 69.73] 0.601 [0.546, 0.658] 84.22 [80.55, 87.53] 0.747 [0.693, 0.803] 69.24 [66.41, 72.08] 0.581 [0.547, 0.617] 69.31 [62.43, 75.66] 0.657 [0.596, 0.717]
+3D MedDiff (Wang et al., 2025) 82.99 [77.51, 87.21] 0.786 [0.739, 0.867] 61.80 [57.94, 66.55] 0.573 [0.510, 0.632] 84.58 [81.15, 89.34] 0.755 [0.700, 0.796] 71.68 [69.55, 74.89] 0.601 [0.568, 0.638] 67.72 [60.15, 74.00] 0.638 [0.581, 0.694]
+Average 82.08 0.774 63.36 0.588 84.71 0.754 70.83 0.591 68.60 0.649
+Ours 85.03 [79.59, 90.48] 0.813 [0.744, 0.880] 65.41 [60.36, 70.09] 0.611 [0.556, 0.664] 87.16 [83.49, 90.46] 0.795 [0.742, 0.843] 72.59 [69.88, 75.42] 0.609 [0.575, 0.646] 71.43 [65.08, 77.78] 0.674 [0.616, 0.733]
+Table 3: Diagnostic performance comparison using eleven classifiers trained on Baseline (A), Real-finetune (B), and Joint-train (C) paradigms in Carotid, TUSC (Stanford-AIMI, 2021), and ACDC (Bernard et al., 2018) datasets.
+Underlined results show no significant difference compared with Baseline (A), all others indicate significant differences ($p<0.05$).
+95% confidence intervals for both metrics are reported.
+“Hybrid”, the classifier with CNN-Transformer design.
+Acc., accuracy ($\%$).
+Classifier Backbone Training Paradigm Carotid TUSC ACDC
 Acc. $\Delta$ AUROC $\Delta$ Acc. $\Delta$ AUROC $\Delta$ Acc. $\Delta$ AUROC $\Delta$
-I3D  (Carreira and Zisserman, 2017) CNN A 79.59 0.737 60.72 0.559 83.49 0.733
-B 85.03 $\uparrow$ 5.44 0.813 $\uparrow$ 0.076 65.41 $\uparrow$ 4.69 0.611 $\uparrow$ 0.052 87.16 $\uparrow$ 3.67 0.795 $\uparrow$ 0.062
-C 84.35 $\uparrow$ 4.76 0.811 $\uparrow$ 0.074 62.88 $\uparrow$ 2.16 0.578 $\uparrow$ 0.019 85.32 $\uparrow$ 1.83 0.762 $\uparrow$ 0.029
-R(2+1)D  (Tran et al., 2018) CNN A 82.99 0.781 63.96 0.595 83.85 0.737
-B 83.67 $\uparrow$ 0.68 0.797 $\uparrow$ 0.016 64.68 $\uparrow$ 0.72 0.601 $\uparrow$ 0.006 85.69 $\uparrow$ 1.84 0.769 $\uparrow$ 0.032
-C 85.04 $\uparrow$ 2.05 0.804 $\uparrow$ 0.023 65.41 $\uparrow$ 1.45 0.610 $\uparrow$ 0.015 85.69 $\uparrow$ 1.84 0.771 $\uparrow$ 0.034
-SlowFast  (Feichtenhofer et al., 2019) CNN A 80.95 0.742 62.88 0.579 81.65 0.702
-B 81.63 $\uparrow$ 0.68 0.770 $\uparrow$ 0.028 65.76 $\uparrow$ 2.88 0.609 $\uparrow$ 0.030 86.05 $\uparrow$ 4.40 0.772 $\uparrow$ 0.070
-C 84.35 $\uparrow$ 3.40 0.811 $\uparrow$ 0.069 65.41 $\uparrow$ 2.53 0.609 $\uparrow$ 0.030 83.85 $\uparrow$ 2.20 0.742 $\uparrow$ 0.040
-CSN  (Tran et al., 2019) CNN A 80.27 0.757 61.80 0.570 85.69 0.765
-B 84.35 $\uparrow$ 4.08 0.784 $\uparrow$ 0.027 67.21 $\uparrow$ 5.41 0.631 $\uparrow$ 0.061 87.15 $\uparrow$ 1.46 0.794 $\uparrow$ 0.029
-C 87.07 $\uparrow$ 6.80 0.843 $\uparrow$ 0.086 65.41 $\uparrow$ 3.61 0.610 $\uparrow$ 0.040 88.62 $\uparrow$ 2.93 0.814 $\uparrow$ 0.049
-TPN  (Yang et al., 2020) CNN A 82.99 0.784 64.32 0.602 83.85 0.735
-B 87.76 $\uparrow$ 4.77 0.856 $\uparrow$ 0.072 66.85 $\uparrow$ 2.53 0.633 $\uparrow$ 0.031 85.32 $\uparrow$ 1.47 0.764 $\uparrow$ 0.029
-C 85.71 $\uparrow$ 2.72 0.823 $\uparrow$ 0.039 65.41 $\uparrow$ 1.09 0.609 $\uparrow$ 0.007 84.95 $\uparrow$ 1.10 0.756 $\uparrow$ 0.021
-TimeSformer  (Bertasius et al., 2021) Transformer A 77.55 0.721 65.41 0.614 72.48 0.562
-B 80.95 $\uparrow$ 3.40 0.753 $\uparrow$ 0.032 66.49 $\uparrow$ 1.08 0.626 $\uparrow$ 0.012 76.51 $\uparrow$ 4.03 0.626 $\uparrow$ 0.064
-C 79.59 $\uparrow$ 2.04 0.730 $\uparrow$ 0.009 66.49 $\uparrow$ 1.08 0.618 $\uparrow$ 0.004 77.62 $\uparrow$ 5.14 0.642 $\uparrow$ 0.080
-MViTv2  (Li et al., 2022b) Transformer A 74.15 0.663 59.64 0.543 77.98 0.642
-B 78.91 $\uparrow$ 4.76 0.740 $\uparrow$ 0.077 63.96 $\uparrow$ 4.32 0.596 $\uparrow$ 0.053 78.72 $\uparrow$ 0.74 0.653 $\uparrow$ 0.011
-C 76.87 $\uparrow$ 2.72 0.701 $\uparrow$ 0.038 62.17 $\uparrow$ 2.53 0.568 $\uparrow$ 0.025 80.18 $\uparrow$ 2.20 0.680 $\uparrow$ 0.038
-VideoSwin  (Liu et al., 2022) Transformer A 79.59 0.732 58.92 0.535 74.31 0.572
-B 83.67 $\uparrow$ 4.08 0.789 $\uparrow$ 0.057 61.08 $\uparrow$ 2.16 0.551 $\uparrow$ 0.016 79.45 $\uparrow$ 5.14 0.663 $\uparrow$ 0.091
-C 83.67 $\uparrow$ 4.08 0.776 $\uparrow$ 0.044 61.44 $\uparrow$ 2.52 0.562 $\uparrow$ 0.027 77.62 $\uparrow$ 3.31 0.632 $\uparrow$ 0.060
-UniFormerV2  (Li et al., 2022a) Hybrid A 76.87 0.710 65.41 0.611 75.78 0.614
-B 80.27 $\uparrow$ 3.40 0.743 $\uparrow$ 0.033 65.77 $\uparrow$ 0.36 0.618 $\uparrow$ 0.007 77.25 $\uparrow$ 1.47 0.645 $\uparrow$ 0.031
-C 81.63 $\uparrow$ 4.76 0.760 $\uparrow$ 0.050 66.13 $\uparrow$ 0.72 0.617 $\uparrow$ 0.006 77.61 $\uparrow$ 1.83 0.637 $\uparrow$ 0.023
-FTC  (Ahmadi et al., 2023) Hybrid A 76.99 0.723 62.09 0.584 80.95 0.689
-B 80.91 $\uparrow$ 3.92 0.750 $\uparrow$ 0.027 64.09 $\uparrow$ 2.00 0.593 $\uparrow$ 0.009 82.38 $\uparrow$ 1.43 0.703 $\uparrow$ 0.014
-C 81.80 $\uparrow$ 4.81 0.758 $\uparrow$ 0.035 65.66 $\uparrow$ 3.57 0.614 $\uparrow$ 0.030 84.38 $\uparrow$ 3.43 0.718 $\uparrow$ 0.029
-CSG-3DCT  (Zhou et al., 2023) Hybrid A 84.35 0.812 61.44 0.571 87.16 0.797
-B 87.07 $\uparrow$ 2.72 0.831 $\uparrow$ 0.019 63.96 $\uparrow$ 2.52 0.590 $\uparrow$ 0.019 88.62 $\uparrow$ 1.46 0.819 $\uparrow$ 0.022
-C 87.76 $\uparrow$ 3.41 0.837 $\uparrow$ 0.025 65.05 $\uparrow$ 3.61 0.609 $\uparrow$ 0.038 88.63 $\uparrow$ 1.47 0.820 $\uparrow$ 0.023
-4.2 Generative Augmentation-based Sequence Classification
+I3D (Carreira and Zisserman, 2017) CNN A 79.59 [73.47, 85.71] 0.737 [0.660, 0.815] 60.72 [56.04, 65.41] 0.559 [0.506, 0.613] 83.49 [79.82, 87.17] 0.733 [0.678, 0.791]
+B 85.03 [79.59, 90.48] $\uparrow$ 5.44 0.813 [0.744, 0.880] $\uparrow$ 0.076 65.41 [60.36, 70.09] $\uparrow$ 4.69 0.611 [0.556, 0.664] $\uparrow$ 0.052 87.16 [83.49, 90.46] $\uparrow$ 3.67 0.795 [0.742, 0.843] $\uparrow$ 0.062
+C 84.35 [78.91, 89.80] $\uparrow$ 4.76 0.811 [0.741, 0.874] $\uparrow$ 0.074 62.88 [58.20, 67.57] $\uparrow$ 2.16 0.578 [0.525, 0.634] $\uparrow$ 0.019 85.32 [81.65, 88.99] $\uparrow$ 1.83 0.762 [0.706, 0.818] $\uparrow$ 0.029
+R(2+1)D (Tran et al., 2018) CNN A 82.99 [76.87, 89.12] 0.781 [0.707, 0.854] 63.96 [59.27, 68.65] 0.595 [0.541, 0.649] 83.85 [80.18, 87.52] 0.737 [0.681, 0.792]
+B 83.67 [77.55, 89.12] $\uparrow$ 0.68 0.797 [0.728, 0.866] $\uparrow$ 0.016 64.68 [60.00, 69.73] $\uparrow$ 0.72 0.601 [0.548, 0.659] $\uparrow$ 0.006 85.69 [82.39, 88.99] $\uparrow$ 1.84 0.769 [0.714, 0.823] $\uparrow$ 0.032
+C 85.04 [78.91, 90.48] $\uparrow$ 2.05 0.804 [0.737, 0.868] $\uparrow$ 0.023 65.41 [60.72, 70.45] $\uparrow$ 1.45 0.610 [0.554, 0.665] $\uparrow$ 0.015 85.69 [82.02, 89.36] $\uparrow$ 1.84 0.771 [0.723, 0.821] $\uparrow$ 0.034
+SlowFast (Feichtenhofer et al., 2019) CNN A 80.95 [74.83, 86.41] 0.742 [0.669, 0.814] 62.88 [57.84, 67.93] 0.579 [0.524, 0.637] 81.65 [77.61, 85.32] 0.702 [0.650, 0.753]
+B 81.63 [75.51, 87.07] $\uparrow$ 0.68 0.770 [0.697, 0.838] $\uparrow$ 0.028 65.76 [61.08, 70.45] $\uparrow$ 2.88 0.609 [0.555, 0.664] $\uparrow$ 0.030 86.05 [82.39, 89.36] $\uparrow$ 4.40 0.772 [0.719, 0.823] $\uparrow$ 0.070
+C 84.35 [78.91, 89.80] $\uparrow$ 3.40 0.811 [0.740, 0.876] $\uparrow$ 0.069 65.41 [60.36, 70.09] $\uparrow$ 2.53 0.609 [0.552, 0.662] $\uparrow$ 0.030 83.85 [80.54, 87.52] $\uparrow$ 2.20 0.742 [0.689, 0.794] $\uparrow$ 0.040
+CSN (Tran et al., 2019) CNN A 80.27 [74.15, 86.39] 0.757 [0.683, 0.828] 61.80 [57.12, 66.85] 0.570 [0.514, 0.627] 85.69 [82.02, 89.36] 0.765 [0.710, 0.821]
+B 84.35 [78.91, 89.80] $\uparrow$ 4.08 0.784 [0.710, 0.853] $\uparrow$ 0.027 67.21 [62.16, 72.25] $\uparrow$ 5.41 0.631 [0.575, 0.688] $\uparrow$ 0.061 87.15 [83.49, 90.83] $\uparrow$ 1.46 0.794 [0.743, 0.845] $\uparrow$ 0.029
+C 87.07 [81.63, 91.84] $\uparrow$ 6.80 0.843 [0.774, 0.905] $\uparrow$ 0.086 65.41 [60.72, 70.10] $\uparrow$ 3.61 0.610 [0.554, 0.665] $\uparrow$ 0.040 88.62 [84.95, 91.93] $\uparrow$ 2.93 0.814 [0.761, 0.865] $\uparrow$ 0.049
+TPN (Yang et al., 2020) CNN A 82.99 [76.87, 88.44] 0.784 [0.711, 0.853] 64.32 [59.64, 69.37] 0.602 [0.546, 0.657] 83.85 [80.18, 87.52] 0.735 [0.681, 0.787]
+B 87.76 [82.99, 92.52] $\uparrow$ 4.77 0.856 [0.795, 0.912] $\uparrow$ 0.072 66.85 [62.16, 71.89] $\uparrow$ 2.53 0.633 [0.580, 0.688] $\uparrow$ 0.031 85.32 [81.65, 88.62] $\uparrow$ 1.47 0.764 [0.708, 0.816] $\uparrow$ 0.029
+C 85.71 [79.59, 90.48] $\uparrow$ 2.72 0.823 [0.751, 0.887] $\uparrow$ 0.039 65.41 [60.72, 70.09] $\uparrow$ 1.09 0.609 [0.556, 0.663] $\uparrow$ 0.007 84.95 [81.28, 88.62] $\uparrow$ 1.10 0.756 [0.702, 0.812] $\uparrow$ 0.021
+TimeSformer (Bertasius et al., 2021) Transformer A 77.55 [70.75, 83.67] 0.721 [0.646, 0.794] 65.41 [60.36, 70.45] 0.614 [0.558, 0.674] 72.48 [69.17, 76.15] 0.562 [0.513, 0.615]
+B 80.95 [74.83, 86.39] $\uparrow$ 3.40 0.753 [0.686, 0.819] $\uparrow$ 0.032 66.49 [61.80, 71.17] $\uparrow$ 1.08 0.626 [0.574, 0.678] $\uparrow$ 0.012 76.51 [72.84, 80.18] $\uparrow$ 4.03 0.626 [0.575, 0.680] $\uparrow$ 0.064
+C 79.59 [72.79, 85.71] $\uparrow$ 2.04 0.730 [0.650, 0.805] $\uparrow$ 0.009 66.49 [61.80, 71.53] $\uparrow$ 1.08 0.618 [0.564, 0.675] $\uparrow$ 0.004 77.62 [73.94, 81.28] $\uparrow$ 5.14 0.642 [0.589, 0.698] $\uparrow$ 0.080
+MViTv2 (Li et al., 2022b) Transformer A 74.15 [68.03, 80.27] 0.663 [0.593, 0.735] 59.64 [54.95, 64.32] 0.543 [0.490, 0.597] 77.98 [74.31, 81.65] 0.642 [0.590, 0.695]
+B 78.91 [72.79, 85.03] $\uparrow$ 4.76 0.740 [0.665, 0.810] $\uparrow$ 0.077 63.96 [59.28, 68.65] $\uparrow$ 4.32 0.596 [0.540, 0.647] $\uparrow$ 0.053 78.72 [75.05, 82.39] $\uparrow$ 0.74 0.653 [0.601, 0.710] $\uparrow$ 0.011
+C 76.87 [70.75, 82.99] $\uparrow$ 2.72 0.701 [0.625, 0.772] $\uparrow$ 0.038 62.17 [57.12, 67.21] $\uparrow$ 2.53 0.568 [0.512, 0.624] $\uparrow$ 0.025 80.18 [76.51, 84.22] $\uparrow$ 2.20 0.680 [0.626, 0.738] $\uparrow$ 0.038
+VideoSwin (Liu et al., 2022) Transformer A 79.59 [73.47, 85.71] 0.732 [0.658, 0.807] 58.92 [54.23, 63.61] 0.535 [0.480, 0.591] 74.31 [70.64, 77.98] 0.572 [0.543, 0.604]
+B 83.67 [78.23, 89.12] $\uparrow$ 4.08 0.789 [0.722, 0.858] $\uparrow$ 0.057 61.08 [56.40, 65.77] $\uparrow$ 2.16 0.551 [0.497, 0.605] $\uparrow$ 0.016 79.45 [75.78, 83.12] $\uparrow$ 5.14 0.663 [0.612, 0.718] $\uparrow$ 0.091
+C 83.67 [77.55, 89.12] $\uparrow$ 4.08 0.776 [0.701, 0.846] $\uparrow$ 0.044 61.44 [56.76, 66.49] $\uparrow$ 2.52 0.562 [0.508, 0.618] $\uparrow$ 0.027 77.62 [73.94, 81.28] $\uparrow$ 3.31 0.632 [0.588, 0.674] $\uparrow$ 0.060
+UniFormerV2 (Li et al., 2022a) Hybrid A 76.87 [70.75, 82.31] 0.710 [0.638, 0.782] 65.41 [60.71, 70.09] 0.611 [0.554, 0.664] 75.78 [72.11, 79.45] 0.614 [0.561, 0.665]
+B 80.27 [74.15, 85.73] $\uparrow$ 3.40 0.743 [0.672, 0.815] $\uparrow$ 0.033 65.77 [61.07, 70.81] $\uparrow$ 0.36 0.618 [0.562, 0.674] $\uparrow$ 0.007 77.25 [73.58, 80.92] $\uparrow$ 1.47 0.645 [0.592, 0.699] $\uparrow$ 0.031
+C 81.63 [76.17, 87.07] $\uparrow$ 4.76 0.760 [0.689, 0.830] $\uparrow$ 0.050 66.13 [61.43, 70.81] $\uparrow$ 0.72 0.617 [0.562, 0.671] $\uparrow$ 0.006 77.61 [73.94, 81.28] $\uparrow$ 1.83 0.637 [0.591, 0.686] $\uparrow$ 0.023
+FTC (Ahmadi et al., 2023) Hybrid A 76.99 [70.86, 80.12] 0.723 [0.669, 0.776] 62.09 [57.71, 64.98] 0.584 [0.523, 0.625] 80.95 [77.93, 83.20] 0.689 [0.618, 0.713]
+B 80.91 [76.44, 82.97] $\uparrow$ 3.92 0.750 [0.685, 0.782] $\uparrow$ 0.027 64.09 [58.56, 67.31] $\uparrow$ 2.00 0.593 [0.541, 0.630] $\uparrow$ 0.009 82.38 [79.30, 85.24] $\uparrow$ 1.43 0.703 [0.638, 0.732] $\uparrow$ 0.014
+C 81.80 [77.23, 83.56] $\uparrow$ 4.81 0.758 [0.702, 0.798] $\uparrow$ 0.035 65.66 [60.28, 68.53] $\uparrow$ 3.57 0.614 [0.567, 0.659] $\uparrow$ 0.030 84.38 [80.21, 86.10] $\uparrow$ 3.43 0.718 [0.651, 0.743] $\uparrow$ 0.029
+CSG-3DCT (Zhou et al., 2023) Hybrid A 84.35 [78.23, 89.80] 0.812 [0.740, 0.879] 61.44 [56.76, 66.49] 0.571 [0.515, 0.626] 87.16 [83.49, 90.46] 0.797 [0.747, 0.846]
+B 87.07 [81.63, 91.84] $\uparrow$ 2.72 0.831 [0.762, 0.897] $\uparrow$ 0.019 63.96 [59.28, 68.65] $\uparrow$ 2.52 0.590 [0.537, 0.644] $\uparrow$ 0.019 88.62 [85.32, 91.93] $\uparrow$ 1.46 0.819 [0.764, 0.872] $\uparrow$ 0.022
+C 87.76 [82.31, 92.52] $\uparrow$ 3.41 0.837 [0.769, 0.902] $\uparrow$ 0.025 65.05 [60.00, 69.73] $\uparrow$ 3.61 0.609 [0.551, 0.664] $\uparrow$ 0.038 88.63 [85.32, 91.93] $\uparrow$ 1.47 0.820 [0.768, 0.872] $\uparrow$ 0.023
+Table 4: Diagnostic performance comparison using eleven classifiers trained on Baseline (A), Real-finetune (B), and Joint-train (C) paradigms in MosMedData (Vladzymyrskyy et al., 2020) and MRNet (Bien et al., 2018) datasets.
+Underlined results show no significant difference compared with Baseline (A), all others indicate significant differences ($p<0.05$).
+95% confidence intervals for both metrics are reported.
+“Hybrid”, the classifier with CNN-Transformer design.
+Acc., accuracy ($\%$).
+Classifier Backbone Training Paradigm MosMedData MRNet
+Acc. $\Delta$ AUROC $\Delta$ Acc. $\Delta$ AUROC $\Delta$
+I3D (Carreira and Zisserman, 2017) CNN A 67.18 [64.35, 70.01] 0.566 [0.535, 0.597] 63.49 [56.08, 70.37] 0.592 [0.532, 0.647]
+B 72.59 [69.88, 75.42] $\uparrow$ 5.41 0.609 [0.575, 0.646] $\uparrow$ 0.043 71.43 [65.08, 77.78] $\uparrow$ 7.94 0.674 [0.616, 0.733] $\uparrow$ 0.082
+C 69.50 [66.67, 72.33] $\uparrow$ 2.32 0.585 [0.551, 0.620] $\uparrow$ 0.019 70.37 [64.02, 76.72] $\uparrow$ 6.88 0.665 [0.606, 0.723] $\uparrow$ 0.073
+R(2+1)D (Tran et al., 2018) CNN A 71.30 [68.34, 74.26] 0.572 [0.541, 0.603] 67.72 [60.83, 74.07] 0.639 [0.579, 0.698]
+B 72.33 [69.50, 75.16] $\uparrow$ 1.03 0.610 [0.577, 0.644] $\uparrow$ 0.038 69.31 [62.43, 76.19] $\uparrow$ 1.59 0.651 [0.592, 0.710] $\uparrow$ 0.012
+C 74.00 [71.30, 76.71] $\uparrow$ 2.70 0.601 [0.569, 0.633] $\uparrow$ 0.029 68.78 [61.90, 75.15] $\uparrow$ 1.06 0.643 [0.584, 0.701] $\uparrow$ 0.004
+SlowFast (Feichtenhofer et al., 2019) CNN A 65.38 [62.55, 68.21] 0.576 [0.541, 0.614] 67.72 [60.85, 74.60] 0.632 [0.574, 0.690]
+B 73.75 [70.91, 76.45] $\uparrow$ 8.37 0.622 [0.586, 0.657] $\uparrow$ 0.046 73.02 [66.67, 79.37] $\uparrow$ 5.30 0.688 [0.630, 0.744] $\uparrow$ 0.056
+C 69.11 [66.28, 72.07] $\uparrow$ 3.73 0.604 [0.568, 0.641] $\uparrow$ 0.028 69.84 [63.49, 76.19] $\uparrow$ 2.12 0.662 [0.600, 0.720] $\uparrow$ 0.030
+CSN (Tran et al., 2019) CNN A 70.79 [67.95, 73.49] 0.636 [0.602, 0.671] 71.96 [65.08, 78.31] 0.685 [0.623, 0.744]
+B 74.13 [71.17, 76.83] $\uparrow$ 3.34 0.643 [0.607, 0.679] $\uparrow$ 0.007 75.66 [69.31, 81.48] $\uparrow$ 3.70 0.719 [0.663, 0.779] $\uparrow$ 0.034
+C 73.36 [70.52, 76.06] $\uparrow$ 2.57 0.641 [0.606, 0.679] $\uparrow$ 0.005 73.54 [66.67, 79.38] $\uparrow$ 1.58 0.694 [0.636, 0.751] $\uparrow$ 0.009
+TPN (Yang et al., 2020) CNN A 74.65 [71.81, 77.48] 0.604 [0.574, 0.635] 71.43 [64.55, 77.78] 0.682 [0.618, 0.744]
+B 76.32 [73.62, 79.28] $\uparrow$ 1.67 0.622 [0.590, 0.658] $\uparrow$ 0.018 77.78 [71.96, 83.60] $\uparrow$ 6.35 0.742 [0.684, 0.798] $\uparrow$ 0.060
+C 76.96 [74.26, 79.67] $\uparrow$ 2.31 0.673 [0.639, 0.710] $\uparrow$ 0.069 75.66 [69.31, 81.48] $\uparrow$ 4.23 0.722 [0.665, 0.784] $\uparrow$ 0.040
+TimeSformer (Bertasius et al., 2021) Transformer A 73.75 [71.04, 76.45] 0.500 [0.500, 0.500] 58.20 [50.79, 65.08] 0.506 [0.500, 0.521]
+B 74.39 [71.69, 77.22] $\uparrow$ 0.64 0.511 [0.502, 0.521] $\uparrow$ 0.011 66.14 [59.79, 72.49] $\uparrow$ 7.94 0.683 [0.620, 0.744] $\uparrow$ 0.177
+C 74.91 [72.20, 77.61] $\uparrow$ 1.16 0.520 [0.509, 0.534] $\uparrow$ 0.020 67.20 [59.79, 74.07] $\uparrow$ 9.00 0.627 [0.570, 0.687] $\uparrow$ 0.121
+MViTv2 (Li et al., 2022b) Transformer A 73.75 [70.91, 76.45] 0.512 [0.498, 0.527] 57.67 [50.26, 64.55] 0.500 [0.500, 0.500]
+B 74.52 [71.69, 77.22] $\uparrow$ 0.77 0.523 [0.507, 0.539] $\uparrow$ 0.011 61.38 [53.97, 68.27] $\uparrow$ 3.71 0.544 [0.514, 0.578] $\uparrow$ 0.044
+C 74.13 [71.30, 76.71] $\uparrow$ 0.38 0.536 [0.515, 0.556] $\uparrow$ 0.024 61.90 [54.50, 68.78] $\uparrow$ 4.23 0.552 [0.518, 0.589] $\uparrow$ 0.052
+VideoSwin (Liu et al., 2022) Transformer A 74.13 [71.30, 76.83] 0.515 [0.502, 0.529] 64.55 [57.14, 71.43] 0.601 [0.543, 0.660]
+B 74.52 [71.69, 77.09] $\uparrow$ 0.39 0.528 [0.511, 0.547] $\uparrow$ 0.013 64.65 [57.67, 71.43] $\uparrow$ 0.10 0.608 [0.544, 0.673] $\uparrow$ 0.007
+C 75.80 [72.84, 78.51] $\uparrow$ 1.67 0.628 [0.592, 0.665] $\uparrow$ 0.113 66.67 [59.79, 73.02] $\uparrow$ 2.12 0.654 [0.584, 0.720] $\uparrow$ 0.053
+UniFormerV2 (Li et al., 2022a) Hybrid A 74.26 [71.56, 76.96] 0.510 [0.501, 0.520] 57.67 [50.26, 64.55] 0.500 [0.500, 0.500]
+B 74.91 [72.07, 77.61] $\uparrow$ 0.65 0.528 [0.511, 0.544] $\uparrow$ 0.018 58.20 [51.32, 65.08] $\uparrow$ 0.53 0.619 [0.560, 0.677] $\uparrow$ 0.019
+C 74.52 [71.69, 77.09] $\uparrow$ 0.26 0.540 [0.520, 0.561] $\uparrow$ 0.030 59.79 [52.38, 66.67] $\uparrow$ 2.12 0.547 [0.492, 0.602] $\uparrow$ 0.047
+FTC (Ahmadi et al., 2023) Hybrid A 74.39 [71.68, 77.10] 0.597 [0.581, 0.624] 63.49 [60.57, 69.81] 0.588 [0.552, 0.651]
+B 76.32 [73.24, 78.33] $\uparrow$ 1.93 0.630 [0.605, 0.663] $\uparrow$ 0.033 66.13 [62.24, 70.99] $\uparrow$ 2.64 0.607 [0.563, 0.674] $\uparrow$ 0.019
+C 77.09 [74.70, 79.54] $\uparrow$ 2.70 0.641 [0.615, 0.675] $\uparrow$ 0.044 68.25 [64.57, 72.10] $\uparrow$ 4.76 0.623 [0.577, 0.691] $\uparrow$ 0.035
+CSG-3DCT (Zhou et al., 2023) Hybrid A 72.33 [69.50, 75.03] 0.616 [0.579, 0.651] 71.43 [64.55, 77.78] 0.677 [0.616, 0.735]
+B 74.00 [71.43, 76.83] $\uparrow$ 1.67 0.622 [0.589, 0.658] $\uparrow$ 0.006 73.02 [66.67, 79.37] $\uparrow$ 1.59 0.688 [0.632, 0.745] $\uparrow$ 0.011
+C 74.78 [71.94, 77.61] $\uparrow$ 2.45 0.622 [0.587, 0.656] $\uparrow$ 0.006 72.49 [66.14, 78.84] $\uparrow$ 1.06 0.688 [0.629, 0.748] $\uparrow$ 0.011
+4.2 Evaluation Metrics
+Existing sequence synthesis works evaluate the synthetic quality by mostly using Fréchet Inception Distance (FID) (Heusel et al., 2017) and Fréchet Video Distance (FVD) (Unterthiner et al., 2018).
+However, researchers have verified that these metrics do not consistently correlate with performance metrics on downstream tasks (Yang et al., 2024; Luo et al., 2024).
+To rigorously validate the effectiveness of our augmentation strategy and to thoroughly assess the value of synthetic data to downstream tasks, we conduct augmentation benchmarks and employ multiple popular classifiers to compare diagnostic accuracy and area under the receiver operating characteristic curve (AUROC).
+Besides, we adopt the proposed VAE-Seq to assess the cross-frame/slice consistency.
+Moreover, motivated by (Huang et al., 2024c), we quantify the smoothness of generated clips using the metric termed Dynamic Smoothness, which is the mean absolute error between the reconstructed frames/slices from the interpolation model (Li et al., 2023) and the original real ones.
+4.3 Statistical Analysis
+In our study, statistical significance between two classification methods evaluated on the same test set is assessed using McNemar’s test (Dietterich, 1998).
+We further report the error bars defined as 95% confidence intervals (CIs) for accuracy and AUROC.
+Uncertainty is quantified using non-parametric percentile bootstrap resampling of the test set (2,000 iterations), where samples are drawn with replacement and the metrics are recomputed for each resample.
+Then, the 95% CIs are obtained from the 2.5th and 97.5th percentiles.
+Therefore, the reported error bars (i.e., CIs) capture test-set sampling variability under a single overall run with fixed experimental conditions.
+4.4 Comparison with Augmentation Methods
+We conducted an extensive set of experiments comparing our method with 15 data augmentation techniques on all five datasets (Table 2).
+Specifically, we grouped these methods into five categories following the augmentation taxonomy in (Cheung and Yeung, 2023), which include: a) Traditional-based, i.e., Basic setting (refer to standard augmentation practices in Table 1), Random elastic deformation; b) Composition-based, i.e., AutoAugment (Cubuk et al., 2019), AdaAug (Cheung and Yeung, 2021), GRA-Video (An et al., 2022), Sample-aware RandAugment (Xiao et al., 2025); c) Mixing-based, i.e., Mixup (Zhang et al., 2017), CutMix (Yun et al., 2019), VideoMix (Yun et al., 2020), LayerMix (Ahmad et al., 2025); d) GAN-based, i.e., TGANv2 (Saito et al., 2020), StyleGAN-V (Skorokhodov et al., 2022); e) Diffusion-based, i.e., VideoComposer (Wang et al., 2023), Endora (Li et al., 2024a), 3D MedDiff (Wang et al., 2025).
+Note that since few studies comprehensively examine the effect of video/3D generative models on downstream tasks, we adopted synthesis-oriented generators from both natural and medical domains as the generative augmentation schemes in our experiments.
+All augmentation methods were based on the Basic setting and employed I3D (Carreira and Zisserman, 2017) as a strong backbone for classification.
+We kept the training datasets for all augmentation methods the same size to ensure fairness.
+As shown in Table 2, our method consistently achieves the highest performance across five datasets on both evaluation metrics, with the majority of performance differences statistically significant ($p<0.05$).
+Our model achieves the best overall accuracy (Carotid: 85.03%, TUSC: 65.41%, ACDC: 87.16%, MosMedData: 72.59%, MRNet: 71.43%) and AUROC (Carotid: 0.813, TUSC: 0.611, ACDC: 0.795, MosMedData: 0.609, MRNet: 0.674), significantly outperforming both traditional augmentations.
+Notably, adding random elastic deformation does not yield noticeable performance improvements across all datasets (Basic: Average Acc.
+70.89%, Elastic deformation: Average Acc.
+71.77%), likely because such local geometric perturbations introduce only limited semantic variation.
+Compared with composition-based augmentations, our method improves the average accuracy by 3.06%, 4.24%, 2.57%, 2.80%, and 5.03%, and AUROC by 0.067, 0.046, 0.049, 0.018, and 0.060 on the Carotid, TUSC, ACDC, MosMedData, and MRNet datasets, respectively.
+Although mixing-based augmentations yield better diagnostic results than composition-based ones in most cases, a clear performance gap remains compared with our method, with differences in mean accuracy/AUROC of 1.53%/0.036, 2.98%/0.033, 2.85%/0.047, 1.93%/0.038, and 2.65%/0.028 across the five datasets.
+Furthermore, as shown in Table 2, generative augmentation paradigms (including GAN- and diffusion-based variants) tend to achieve higher average performance than other augmentation methods, highlighting the potential of leveraging synthetic data to enhance classifier learning.
+Within generative augmentations, GAN-based variants generally underperform diffusion-based ones, indicating that the limited fidelity and diversity of GAN-generated samples constrain the magnitude of the generative augmentation gains.
+In contrast, diffusion-based competitors display stronger overall recognition performance due to their more faithful and diverse synthesis, reaching mean accuracies of 82.08%, 63.36%, 84.71%, 70.83%, and 68.60%, and AUROCs of 0.774, 0.588, 0.754, 0.591, and 0.649 on the Carotid, TUSC, ACDC, MosMedData, and MRNet datasets, respectively.
+Based on the experimental results, with the combination of customized sequence synthesis and pruning of noisy synthetic samples, our diffusion-based approach functions as a powerful and general data augmentation suite across multiple organs and medical modalities.
+Figure 7: Carotid performance radar map, using 4 models trained on Baseline (green) and Joint-train (blue) paradigms in underrepresented high-risk sets.
+Our proposed framework generally enhances diagnostic performance across multiple evaluation metrics within these challenging subsets.
+4.5 Generative Augmentation for Diagnosis with Multiple Sequence Classifiers
 We utilized Real-finetune and Joint-train paradigms to integrate real and synthetic clips for downstream multi-class sequence classification training.
@@
 To demonstrate the generalizability of Ctrl-GenAug, the selected networks involve 3D CNN, transformer, and CNN-Transformer hybrid designs, which specialize in capturing local, global, and mixed features, respectively.
-Table 2 compares the average performance (i.e., accuracy and AUROC) on different training paradigms across a wide range of sequence classifiers, including five CNN-based classic models, three transformer-based models, and three CNN-Transformer hybrid-based methods.
-It is evident that the performance ceiling of existing popular models, from simple CNN-based to complex CNN-Transformer hybrids, can be further elevated using synthetic samples obtained from our proposed framework, under Real-finetune and Joint-train paradigms.
-On the carotid dataset, the accuracy/AUROC can be most boosted by 5.44%/0.077 and 6.80%/0.086 through Real-finetune and Joint-train paradigms, respectively.
-Due to high intra-class variations in TUSC, the baseline exhibits visibly inferior results, with an average accuracy of 62.42% compared to other datasets.
-However, it can be seen from Table 2 that samples synthesized by Ctrl-GenAug make significant contributions and enhance the classification performance to an acceptance level.
-Among all the classifiers, CSN harvests the largest performance gain, which is non-trivially enhanced by 5.41%/0.061 and 3.61%/0.040 accuracy/AUROC under Real-finetune and Joint-train paradigms, respectively.
-Notably, for ACDC, the overall performance improvement remains significant even with a high baseline.
-For instance, CSG-3DCT is boosted from 87.16$\rightarrow$88.63(+1.47%) accuracy and from 0.797$\rightarrow$0.820 (+0.023) AUROC under the Joint-train practice.
-Extensive experimental results show that synthetic sample sets produced by Ctrl-GenAug can be powerful data engines to enhance various diagnostic tasks in medical sequence analysis.
+[…]
+Table 5: Out-domain performance comparison using eleven nets trained on Baseline (A), Real-finetune (B), and Joint-train (C) paradigms in carotid dataset.
+Underlined results show no significant difference compared with Baseline (A), all others indicate significant differences ($p<0.05$).
+95% confidence intervals for both metrics are reported.
+“Hybrid”, the net with CNN-Transformer design.
+Acc., accuracy ($\%$).
+Classifier Backbone Training Paradigm Carotid
+Acc. $\Delta$ AUROC $\Delta$
+I3D (Carreira and Zisserman, 2017) CNN A 70.00 [60.00, 80.00] 0.609 [0.509, 0.721]
+B 71.67 [61.67, 81.67] $\uparrow$ 1.67 0.648 [0.537, 0.771] $\uparrow$ 0.039
+C 75.00 [65.00, 85.00] $\uparrow$ 5.00 0.682 [0.565, 0.794] $\uparrow$ 0.073
+R(2+1)D (Tran et al., 2018) CNN A 66.67 [56.67, 76.71] 0.577 [0.471, 0.690]
+B 71.67 [61.67, 81.67] $\uparrow$ 5.00 0.661 [0.556, 0.773] $\uparrow$ 0.084
+C 75.00 [65.00, 85.00] $\uparrow$ 8.33 0.689 [0.575, 0.810] $\uparrow$ 0.112
+SlowFast (Feichtenhofer et al., 2019) CNN A 73.33 [63.33, 83.33] 0.638 [0.528, 0.750]
+B 75.00 [65.00, 85.00] $\uparrow$ 1.67 0.670 [0.562, 0.787] $\uparrow$ 0.032
+C 75.00 [65.00, 85.00] $\uparrow$ 1.67 0.706 [0.588, 0.817] $\uparrow$ 0.068
+CSN (Tran et al., 2019) CNN A 68.33 [58.33, 78.33] 0.592 [0.494, 0.703]
+B 73.33 [63.33, 83.33] $\uparrow$ 5.00 0.664 [0.556, 0.786] $\uparrow$ 0.072
+C 71.67 [61.67, 81.67] $\uparrow$ 3.34 0.672 [0.552, 0.784] $\uparrow$ 0.080
+TPN (Yang et al., 2020) CNN A 75.00 [65.00, 85.00] 0.667 [0.563, 0.783]
+B 76.67 [66.67, 86.67] $\uparrow$ 1.67 0.731 [0.624, 0.831] $\uparrow$ 0.064
+C 76.67 [66.67, 86.67] $\uparrow$ 1.67 0.678 [0.569, 0.790] $\uparrow$ 0.011
+TimeSformer (Bertasius et al., 2021) Transformer A 68.33 [58.33, 78.33] 0.590 [0.489, 0.689]
+B 71.67 [61.62, 81.67] $\uparrow$ 3.34 0.665 [0.543, 0.787] $\uparrow$ 0.075
+C 73.33 [63.33, 83.33] $\uparrow$ 5.00 0.664 [0.565, 0.773] $\uparrow$ 0.074
+MViTv2 (Li et al., 2022b) Transformer A 66.67 [56.67, 76.67] 0.590 [0.480, 0.699]
+B 68.33 [58.33, 78.33] $\uparrow$ 1.66 0.592 [0.493, 0.699] $\uparrow$ 0.002
+C 70.00 [60.00, 80.00] $\uparrow$ 3.33 0.604 [0.508, 0.706] $\uparrow$ 0.014
+VideoSwin (Liu et al., 2022) Transformer A 65.00 [55.00, 75.00] 0.574 [0.462, 0.688]
+B 70.00 [58.33, 80.00] $\uparrow$ 5.00 0.650 [0.527, 0.767] $\uparrow$ 0.076
+C 68.33 [58.33, 78.33] $\uparrow$ 3.33 0.609 [0.503, 0.728] $\uparrow$ 0.035
+UniFormerV2 (Li et al., 2022a) Hybrid A 66.67 [56.67, 76.67] 0.617 [0.498, 0.735]
+B 73.33 [63.33, 83.33] $\uparrow$ 6.66 0.681 [0.564, 0.803] $\uparrow$ 0.064
+C 68.33 [58.33, 78.33] $\uparrow$ 1.66 0.633 [0.521, 0.756] $\uparrow$ 0.016
+FTC (Ahmadi et al., 2023) Hybrid A 67.08 [56.67, 78.33] 0.616 [0.502, 0.741]
+B 71.31 [66.67, 80.00] $\uparrow$ 4.23 0.632 [0.523, 0.752] $\uparrow$ 0.016
+C 74.68 [66.67, 83.33] $\uparrow$ 7.60 0.670 [0.553, 0.784] $\uparrow$ 0.054
+CSG-3DCT (Zhou et al., 2023) Hybrid A 68.33 [58.33, 78.33] 0.599 [0.491, 0.715]
+B 73.33 [63.33, 83.33] $\uparrow$ 5.00 0.689 [0.569, 0.805] $\uparrow$ 0.090
+C 70.00 [60.00, 80.00] $\uparrow$ 1.67 0.651 [0.530, 0.772] $\uparrow$ 0.052
 4.6 Improved Diagnosis in Underrepresented High-risk Sets
@@
 Four downstream networks were trained on the augmented dataset with Joint-train paradigm, and six metrics were used for class-level evaluation, including sensitivity, specificity, accuracy, F1-score, precision, and AUROC.
-Fig.
-7 quantitatively compares the carotid stenosis diagnostic performance of underrepresented high-risk sets on Baseline and Joint-train paradigms.
-It shows that Baseline performs poorly on underrepresented high-risk sets, especially in terms of sensitivity (average 18.52% in moderate and 36.25% in severe sets).
-In comparison, the classifiers can be greatly enhanced by jointly training with synthetic samples (e.g., SlowFast: 11.11$\rightarrow$44.44sensitivity on the moderate set).
-We also provided the category-level analysis, and summarized the average F1 scores across four classifiers for a global observation and conclusion.
-Specifically, Baseline reached 69.92%, 25.92%, and 43.44% average F1 scores for mild, moderate, and severe cases, respectively.
-The F1 performance of Joint-train in these three categories was 82.44%, 52.73%, and 60.40%.
-We found that well-designed classifiers jointly trained on synthetic samples from Ctrl-GenAug and real ones, can effectively close diagnostic performance gaps between overrepresented (i.e., mild) and underrepresented sets (i.e., moderate and severe) while improving the former ($\sim$13$\uparrow$).
-Figure 7: Carotid performance radar map, using 4 models trained on Baseline (green) and Joint-train (blue) paradigms in underrepresented high-risk sets.
+[…]
+[…]
 4.7 Classification Robustness in Out-domain Conditions
@@
 For the OD data, which did not contain the diagnostic class label, we padded the corresponding conditioning vector with zeros, while solely preserving hospital IDs to form descriptive texts.
-As shown in Table 3, the average accuracy and AUROC of three classifiers display better results on the Real-finetune and Joint-train paradigms compared to the Baseline.
+As shown in Table 5, the average accuracy and AUROC of all eleven classifiers display better results on the Real-finetune and Joint-train paradigms compared to the Baseline, with most of the improvements showing statistical significance ($p<0.05$).
 It validates that diagnostic robustness can be enhanced with the aid of Ctrl-GenAug in scenarios where we only have access to unlabeled cases from additional medical centers due to limited resources.
-[…]
-Method Backbone Training Paradigm Carotid
-Acc. $\Delta$ AUROC $\Delta$
-I3D  (Carreira and Zisserman, 2017) CNN A 70.00 0.609
-B 71.67 $\uparrow$ 1.67 0.648 $\uparrow$ 0.039
-C 75.00 $\uparrow$ 5.00 0.682 $\uparrow$ 0.073
-TPN  (Yang et al., 2020) CNN A 75.00 0.667
-B 76.67 $\uparrow$ 1.67 0.731 $\uparrow$ 0.064
-C 76.67 $\uparrow$ 1.67 0.678 $\uparrow$ 0.011
-CSG-3DCT  (Zhou et al., 2023) Hybrid A 68.33 0.599
-B 73.33 $\uparrow$ 5.00 0.689 $\uparrow$ 0.090
-C 70.00 $\uparrow$ 1.67 0.651 $\uparrow$ 0.052
-Table 4: Ablation study for different conditional controls in TUSC (Stanford-AIMI, 2021), including class label (C), text (T), image prior (I), and motion field (MF).
+Table 6: Ablation study for different conditional controls in multi-organ and multi-modal datasets, including class label (C), text (T), image prior (I), and motion field (MF).
 SlowFast (Feichtenhofer et al., 2019) (C1), CSN (Tran et al., 2019) (C2), and FTC (Ahmadi et al., 2023) (C3) trained on Joint-train paradigm were used for downstream diagnosis.
+The accuracy and AUROC gains over the Baseline are reported.
 “Hybrid”, the classifier with CNN-Transformer design.
 Dataset Synthesis Downstream Diagnosis
-Control Metric Backbone $\Delta$Accuracy $\Delta$AUROC
+Control Metric Backbone $\DeltaAccuracy Δ\DeltaAUROC$
 […]
+Carotid 4.18 89.65% 93.27% C1 CNN $\uparrow$ 2.04% $\uparrow$ 0.009
+C2 CNN $\downarrow$ 0.68% $\downarrow$ 0.009
+C3 Hybrid $\uparrow$ 0.56% $\uparrow$ 0.031
+6.50 89.08% 92.85% C1 CNN $\uparrow$ 1.36% $\uparrow$ 0.046
+C2 CNN $\downarrow$ 0.68% $\downarrow$ 0.011
+C3 Hybrid $\uparrow$ 1.12% $\uparrow$ 0.021
+4.29 88.74% 93.09% C1 CNN $\uparrow$ 2.72% $\uparrow$ 0.078
+C2 CNN $\uparrow$ 4.09% $\uparrow$ 0.094
+C3 Hybrid $\uparrow$ 3.28% $\uparrow$ 0.064
+4.26 91.46% 94.29% C1 CNN $\uparrow$ 3.40% $\uparrow$ 0.069
+C2 CNN $\uparrow$ 6.80% $\uparrow$ 0.086
+C3 Hybrid $\uparrow$ 4.81% $\uparrow$ 0.035
 TUSC (Stanford-AIMI, 2021) 4.18 90.40% 95.03% C1 CNN $\downarrow$ 1.80% $\downarrow$ 0.015
@@
 C3 Hybrid $\uparrow$ 3.57% $\uparrow$ 0.030
-4.5 Analysis of Multimodal Conditions Guidance
-To illustrate the role of different conditional guidance, we conduct comparative experiments with generators that resort to various banks of conditions for training and sampling.
-Table 4 validates that all proposed conditions are effective for enhancing synthesis and diagnosis tasks.
-Specifically, without text control, the downstream diagnostic performance is comparatively poor.
-Fig.
-6(b-c) compares typical synthetic thyroid nodule sequences generated under condition banks with and without text guidance.
-The former is visibly high-fidelity and faithful to the given text prompt (e.g., smooth margin), while the latter exhibits less distinguishable features for diagnosis (e.g., blurry margin).
-That is to say, the proposed Ctrl-GenAug can create more diagnosis-reliable samples by enhancing semantic steerability in the generation process.
-Besides, by incorporating image prior knowledge in the sampling process, the average accuracy and AUROC of three classifiers are improved by 4.20% and 0.040, respectively.
-This proves the domain gap between synthetic and real samples is mitigated by introducing the image prior.
-Moreover, as shown in Table 4, conditioning the generator on the motion field produces more content-consistent and smoother samples, leading to continuous improvements in diagnostic performance.
-It can also be observed that the FVD results show a limited correlation with downstream evaluation metrics, confirming the finding in (Luo et al., 2024).
-Table 5: Ablation study for the sequential augmentation module of the generator in TUSC (Stanford-AIMI, 2021) and ACDC (Bernard et al., 2018).
+ACDC (Bernard et al., 2018) - 8.70 76.74% 92.50% C1 CNN $\uparrow$ 1.46% $\uparrow$ 0.010
+C2 CNN $\uparrow$ 0.37% $\uparrow$ 0.027
+C3 Hybrid $\uparrow$ 1.10% $\uparrow$ 0.051
+- 10.81 79.66% 93.89% C1 CNN $\uparrow$ 0.37% $\uparrow$ 0.020
+C2 CNN $\uparrow$ 0.74% $\uparrow$ 0.012
+C3 Hybrid $\uparrow$ 1.83% $\uparrow$ 0.029
+- 8.18 77.08% 92.77% C1 CNN $\uparrow$ 2.20% $\uparrow$ 0.040
+C2 CNN $\uparrow$ 2.93% $\uparrow$ 0.049
+C3 Hybrid $\uparrow$ 3.43% $\uparrow$ 0.029
+MosMedData (Vladzymyrskyy et al., 2020) - 2.98 83.06% 95.27% C1 CNN $\uparrow$ 2.96% $\uparrow$ 0.004
+C2 CNN $\uparrow$ 1.41% $\downarrow$ 0.091
+C3 Hybrid $\uparrow$ 1.15% $\uparrow$ 0.070
+- 5.05 84.72% 96.35% C1 CNN $\uparrow$ 1.93% $\downarrow$ 0.026
+C2 CNN $\uparrow$ 0.38% $\downarrow$ 0.114
+C3 Hybrid $\uparrow$ 1.81% $\uparrow$ 0.054
+- 2.91 84.07% 96.33% C1 CNN $\uparrow$ 3.73% $\uparrow$ 0.028
+C2 CNN $\uparrow$ 2.57% $\uparrow$ 0.005
+C3 Hybrid $\uparrow$ 2.70% $\uparrow$ 0.044
+MRNet (Bien et al., 2018) - 8.48 75.48% 92.90% C1 CNN $\uparrow$ 2.12% $\uparrow$ 0.018
+C2 CNN $\uparrow$ 0.53% $\uparrow$ 0.045
+C3 Hybrid $\uparrow$ 3.17% $\uparrow$ 0.057
+- 11.67 76.67% 93.61% C1 CNN $\uparrow$ 1.59% $\uparrow$ 0.017
+C2 CNN $\downarrow$ 2.12% $\downarrow$ 0.020
+C3 Hybrid $\uparrow$ 1.06% $\uparrow$ 0.026
+- 8.14 77.10% 93.81% C1 CNN $\uparrow$ 2.12% $\uparrow$ 0.030
+C2 CNN $\uparrow$ 1.58% $\uparrow$ 0.009
+C3 Hybrid $\uparrow$ 4.76% $\uparrow$ 0.035
+Table 7: Ablation study for the sequential augmentation module of the generator in multi-organ and multi-modal datasets.
 SlowFast (Feichtenhofer et al., 2019) (C1), CSN (Tran et al., 2019) (C2), and FTC (Ahmadi et al., 2023) (C3) trained on Joint-train paradigm were used for downstream diagnosis.
+The accuracy and AUROC gains over the Baseline are reported.
 “Hybrid”, the classifier with CNN-Transformer design.
@@
 […]
+Carotid Ours-S 4.53 85.39% 87.45% C1 CNN $\uparrow$ 1.40% $\uparrow$ 0.031
+C2 CNN $\uparrow$ 1.12% $\uparrow$ 0.009
+C3 Hybrid $\downarrow$ 0.68% $\downarrow$ 0.017
+Ours-SK 4.75 89.03% 93.38% C1 CNN $\uparrow$ 2.04% $\uparrow$ 0.044
+C2 CNN $\uparrow$ 1.36% $\uparrow$ 0.053
+C3 Hybrid $\uparrow$ 2.72% $\uparrow$ 0.062
+Ours-SKM 4.26 91.46% 94.29% C1 CNN $\uparrow$ 3.40% $\uparrow$ 0.069
+C2 CNN $\uparrow$ 6.80% $\uparrow$ 0.086
+C3 Hybrid $\uparrow$ 4.81% $\uparrow$ 0.035
 TUSC (Stanford-AIMI, 2021) Ours-S 3.76 87.44% 91.07% C1 CNN $\downarrow$ 0.36% $\downarrow$ 0.009
@@
 C3 Hybrid $\uparrow$ 3.43% $\uparrow$ 0.029
-4.6 Impact of the Sequential Augmentation Module
+MosMedData (Vladzymyrskyy et al., 2020) Ours-S 5.17 75.23% 89.05% C1 CNN $\uparrow$ 1.54% $\downarrow$ 0.002
+C2 CNN $\downarrow$ 0.13% $\downarrow$ 0.102
+C3 Hybrid $\downarrow$ 0.76% $\downarrow$ 0.013
+Ours-SK 2.98 84.19% 96.16% C1 CNN $\uparrow$ 2.83% $\uparrow$ 0.008
+C2 CNN $\uparrow$ 1.28% $\downarrow$ 0.017
+C3 Hybrid $\uparrow$ 1.41% $\uparrow$ 0.022
+Ours-SKM 2.91 84.07% 96.33% C1 CNN $\uparrow$ 3.73% $\uparrow$ 0.028
+C2 CNN $\uparrow$ 2.57% $\uparrow$ 0.005
+C3 Hybrid $\uparrow$ 2.70% $\uparrow$ 0.044
+MRNet (Bien et al., 2018) Ours-S 8.46 73.88% 89.68% C1 CNN $\downarrow$ 0.52% $\downarrow$ 0.060
+C2 CNN $\uparrow$ 0.00% $\downarrow$ 0.001
+C3 Hybrid $\uparrow$ 0.53% $\uparrow$ 0.004
+Ours-SK 9.20 76.81% 93.34% C1 CNN $\uparrow$ 2.65% $\uparrow$ 0.036
+C2 CNN $\uparrow$ 1.58% $\uparrow$ 0.056
+C3 Hybrid $\uparrow$ 2.12% $\uparrow$ 0.029
+Ours-SKM 8.14 77.10% 93.81% C1 CNN $\uparrow$ 2.12% $\uparrow$ 0.030
+C2 CNN $\uparrow$ 1.58% $\uparrow$ 0.009
+C3 Hybrid $\uparrow$ 4.76% $\uparrow$ 0.035
+4.8 Ablation Study
+4.8.1 Analysis of Multimodal Conditions Guidance
+To illustrate the role of different conditional guidance, we used five datasets to conduct comparative experiments with generators that resort to various banks of conditions for training and sampling.
+Table 6 validates that all proposed conditions are effective for enhancing synthesis and diagnosis tasks.
+Specifically, without text control, the downstream diagnostic performance is comparatively poor.
+Fig.
+6(c-d) compares typical synthetic thyroid nodule sequences generated under condition banks with and without text guidance.
+The former is visibly high-fidelity and faithful to the given text prompt (e.g., smooth margin), while the latter exhibits less distinguishable features for diagnosis (e.g., blurry margin).
+That is to say, the proposed Ctrl-GenAug can create more diagnosis-reliable samples by enhancing semantic steerability in the generation process.
+Besides, by incorporating image prior knowledge in the sampling process, the three classifiers achieve consistent performance improvements on all datasets, with average accuracy/AUROC gains of Carotid (+4.40%, +0.045), TUSC (+4.20%, +0.040), ACDC (+1.87%, +0.019), MosMedData (+1.63%, +0.054), and MRNet (+2.64%, +0.017).
+This proves the domain gap between synthetic and real samples is mitigated by introducing the image prior.
+Moreover, as shown in Table 6, conditioning the generator on the motion field produces more content-consistent and smoother samples, leading to continuous improvements in diagnostic performance.
+It can also be observed that the FVD results show a limited correlation with downstream evaluation metrics, confirming the finding in Luo et al.
+(2024).
+4.8.2 Impact of the Sequential Augmentation Module
 To verify the impact of the proposed sequential augmentation module, as shown in Table 7, we compared the performance of three generator variants on the synthesis and downstream tasks.
 Ours-S, Ours-SK, and Ours-SKM denote our ablation studies, including gradually adding sequential attention (’-S’), key-frame/slice attention (’-K’), and motion field attention (’-M’) to the generator with spatial inflation only.
-Table 5 shows that the downstream accuracy and AUROC improvements of Ours-SKM are significantly higher than those of Ours-S and Ours-SK in both datasets.
+Table 7 shows that the holistic downstream accuracy and AUROC improvements of Ours-SKM are significantly higher than those of Ours-S and Ours-SK in five datasets.
 In terms of quantitative assessment of sequential coherence, Ours-SKM achieves better results overall than other variants.
 These prove that this module can help the generator synthesize cross-frame/slice consistent and dynamic-smooth clips, which are diagnosis-promotive for downstream diagnostic tasks.
-Moreover, it can be observed that the FVD results across these three generators show no significant difference, revealing its limited value for evaluating diagnosis-oriented medical sequence synthesis tasks.
-4.7 Effectiveness of the Noisy Synthetic Data Filter
+Moreover, it can be observed that the FVD results across these three generators show no significant difference on four datasets, revealing its limited value for evaluating diagnosis-oriented medical sequence synthesis tasks.
+Figure 8: Carotid stenosis and heart disease diagnostic accuracy (%) comparison using two classifiers on our filtering strategies under Real-finetune and Joint-train paradigms.
+The variant ’CF+SF’ indicates our final filter, and the accuracy gains over the Baseline are depicted.
+4.8.3 Effectiveness of the Noisy Synthetic Data Filter
 As shown in Fig.
@@
 Equipped with both, it achieves an accuracy of 87.07%, outperforming the ’None’ by 6.12%.
-Figure 8: Carotid stenosis and heart disease diagnostic accuracy (%) comparison using two classifiers on our filtering strategies under Real-finetune and Joint-train paradigms.
-The accuracy gains over the Baseline are depicted.
+Table 8: Ablation study for performance-oriented diffusion settings in TUSC (Stanford-AIMI, 2021) and ACDC (Bernard et al., 2018) datasets.
+I3D (Carreira and Zisserman, 2017) trained on Real-finetune paradigm was used for downstream diagnosis.
+The blue-highlighted block represents our default configuration.
+Dataset Diffusion Settings Synthesis Diagnosis
 […]
-Generative Training Paradigm Augmentation Method Carotid TUSC (Stanford-AIMI, 2021) ACDC (Bernard et al., 2018)
-Acc. $\Delta$ AUROC $\Delta$ Acc. $\Delta$ AUROC $\Delta$ Acc. $\Delta$ AUROC $\Delta$
-/ None 75.51 0.718 59.28 0.532 80.55 0.681
-/ TraAug 79.59 $\uparrow$ 4.08 0.737 $\uparrow$ 0.019 60.72 $\uparrow$ 1.44 0.559 $\uparrow$ 0.027 83.49 $\uparrow$ 2.94 0.733 $\uparrow$ 0.052
-Real-finetune TraAug + GenAug 85.03 $\uparrow$ 9.52 0.813 $\uparrow$ 0.095 65.41 $\uparrow$ 6.13 0.611 $\uparrow$ 0.079 87.16 $\uparrow$ 6.61 0.795 $\uparrow$ 0.114
-/ None 78.91 0.735 60.36 0.552 80.92 0.688
-/ TraAug 80.95 $\uparrow$ 2.04 0.751 $\uparrow$ 0.016 61.08 $\uparrow$ 0.72 0.562 $\uparrow$ 0.010 81.28 $\uparrow$ 0.36 0.696 $\uparrow$ 0.008
-Joint-train TraAug + GenAug 84.35 $\uparrow$ 5.44 0.811 $\uparrow$ 0.076 62.88 $\uparrow$ 2.52 0.578 $\uparrow$ 0.026 85.32 $\uparrow$ 4.40 0.762 $\uparrow$ 0.074
-4.8 Impacts of Traditional and Generative Augmentations
-We tested the contributions of different data augmentations in three downstream diagnostic tasks.
-The traditional augmentations used in this experiment can be found in Table 1.
-We report the performance gains of the two variants with data augmentations compared to the augmentation-free setup.
-For experimental fairness, we kept the training datasets for the None and TraAug variants the same size as the method combining traditional and generative augmentations.
-As presented in Table 6, the diagnostic performance shows noticeable improvements using the traditional augmentation method.
-Further coupled with the generative augmentation, the results show more substantial improvements than TraAug.
-We also observe that in the case of using traditional augmentation, adding extra real training samples yields limited performance gains, even a drop for ACDC.
-These prove that generative augmentation can effectively supplement traditional data augmentation and provide intrinsically diverse data characteristics for enhancing deep classifier learning, which support the viewpoints of (Luo et al., 2024).
+TUSC Cosine 12.5 5.19 92.79% 95.30% 63.24% 0.588
+Cosine 7.5 3.90 92.57% 95.76% 64.69% 0.600
+Scaled Linear 12.5 4.57 92.16% 95.45% 64.33% 0.597
+Scaled Linear 7.5 3.69 92.14% 95.59% 65.41% 0.611
+ACDC Cosine 12.5 8.42 73.95% 90.71% 85.32% 0.762
+Cosine 7.5 8.63 73.88% 90.74% 86.05% 0.775
+Scaled Linear 12.5 8.16 76.75% 92.56% 86.06% 0.774
+Scaled Linear 7.5 8.18 77.08% 92.77% 87.16% 0.795
+Table 9: Ablation study for efficiency-performance trade-off across diffusion settings in TUSC (Stanford-AIMI, 2021) and ACDC (Bernard et al., 2018) datasets.
+I3D (Carreira and Zisserman, 2017) trained on Real-finetune paradigm was used for downstream diagnosis.
+All diffusion models are tested on a single NVIDIA GeForce RTX 3090 GPU.
+The blue-highlighted block represents our default configuration.
+DS, Dynamic Smoothness.
+Dataset Diffusion Settings Synthesis Diagnosis
+[…]
+TUSC 3 200 13.26 86.29% 90.30% 53.5s 62.52% 0.575
+3 300 12.33 83.42% 88.19% 87.4s 62.88% 0.580
+4 100 3.70 92.02% 95.53% 39.8s 63.97% 0.592
+4 200 3.69 92.14% 95.59% 68.8s 65.41% 0.611
+4 300 2.69 91.15% 95.23% 100.3s 65.77% 0.622
+ACDC 3 200 10.81 75.15% 88.73% 51.3s 83.85% 0.741
+3 300 10.17 73.52% 88.41% 79.9s 83.85% 0.739
+4 100 8.20 76.86% 92.74% 33.7s 85.69% 0.771
+4 200 8.18 77.08% 92.77% 70.7s 87.16% 0.795
+4 300 7.55 74.59% 91.90% 103.8s 88.06% 0.806
+4.8.4 Impact of Diffusion Factors in Sequence Generator
+To systematically analyze how the key diffusion factors affect the quality of the synthetic databases, we conducted two ablation studies that examine performance-oriented diffusion settings, including the noise schedule and guidance scale (Table 8), as well as efficiency-performance trade-off analyses under the settings of different UNet depths and sampling steps (Table 9).
+As shown in Table 8, the scaled linear noise schedule consistently yields stronger downstream augmentation benefits and also achieves lower FVD than the cosine variant across both datasets.
+We believe that this advantage stems from its smoother and more stable noise evolution during training, which allows the model to more effectively learn and preserve anatomical details throughout the diffusion process.
+Moreover, the results show that using a high guidance scale of 12.5 leads to a noticeable degradation in diagnostic performance compared with the moderate setting of 7.5, with accuracy decreasing from 65.41% to 64.33% and AUROC dropping from 0.611 to 0.597 under the scaled linear schedule on the TUSC dataset.
+This trend suggests that excessive guidance may constrain the sampling trajectory, contributing to in-distribution (i.e., overly close to the real training distribution) synthetic samples, which in turn limit their usefulness for downstream model optimization.
+In contrast, a moderate guidance scale strikes a more favorable balance between semantic alignment and diversity, enabling the synthesizer to generate more diagnosis-promotive samples.
+The depth of the UNet denoiser and the number of sampling steps both exert dual impacts on overall performance and synthesis efficiency.
+Table 9 demonstrates that increasing the UNet depth consistently improves generation fidelity and downstream classification, since deeper architectures can capture richer anatomical patterns and finer dynamic cues during training.
+However, these performance gains come with additional computational cost.
+A similar trend is observed for sampling steps, where additional iterations yield more diagnosis-reliable sequences but also require longer sampling time.
+Notably, configurations with insufficient feature representation capacity, such as a UNet depth of three, benefit minimally from increasing the sampling steps, indicating that additional steps cannot compensate for limited feature extraction ability.
+We further observe that the sequential consistency and smoothness of our synthetic samples are substantially more sensitive to the UNet depth than to other diffusion factors, as it directly determines the network’s capacity to model and propagate temporal/stereoscopic dependencies.
+Overall, our default configuration achieves a balanced compromise between generation quality and inference efficiency, providing a practical choice for large-scale customized medical sequence synthesis and generative augmentation.
+Table 10: Sensitivity analysis on four representative classifier backbones used in the class semantics misalignment filtering across the Carotid, ACDC (Bernard et al., 2018), and MosMedData (Vladzymyrskyy et al., 2020) datasets.
+I3D (Carreira and Zisserman, 2017) trained on Real-finetune and Joint-train paradigms was adopted for downstream diagnosis.
+The accuracy and AUROC improvements over the Baseline are reported, with the best gains underlined.
+Training Paradigm Semantic Filtering Classifier Carotid ACDC MosMedData
+$\DeltaAccuracy Δ\DeltaAUROC$ $\DeltaAccuracy Δ\DeltaAUROC$ $\DeltaAccuracy Δ\DeltaAUROC$
+Real-finetune CSN $\uparrow$ 4.08% $\uparrow$ 0.064 $\uparrow$ 2.56% $\uparrow$ 0.041 $\uparrow$ 4.76% $\uparrow$ 0.027
+VideoSwin $\uparrow$ 4.08% $\uparrow$ 0.066 $\uparrow$ 1.83% $\uparrow$ 0.035 $\uparrow$ 5.66% $\uparrow$ 0.035
+CSG-3DCT $\uparrow$ 5.44% $\uparrow$ 0.056 $\uparrow$ 2.57% $\uparrow$ 0.043 $\uparrow$ 4.25% $\uparrow$ 0.018
+I3D $\uparrow$ 5.44% $\uparrow$ 0.076 $\uparrow$ 3.67% $\uparrow$ 0.062 $\uparrow$ 5.41% $\uparrow$ 0.043
+Joint-train CSN $\uparrow$ 3.40% $\uparrow$ 0.042 $\uparrow$ 1.10% $\uparrow$ 0.021 $\uparrow$ 2.32% $\uparrow$ 0.007
+VideoSwin $\uparrow$ 4.76% $\uparrow$ 0.045 $\uparrow$ 1.47% $\uparrow$ 0.027 $\uparrow$ 2.58% $\uparrow$ 0.005
+CSG-3DCT $\uparrow$ 4.08% $\uparrow$ 0.071 $\uparrow$ 1.83% $\uparrow$ 0.028 $\uparrow$ 2.45% $\uparrow$ 0.006
+I3D $\uparrow$ 4.76% $\uparrow$ 0.074 $\uparrow$ 1.83% $\uparrow$ 0.029 $\uparrow$ 2.32% $\uparrow$ 0.019
+4.8.5 Analysis of Classifier Backbone Choice in the Semantic Filter
+To investigate whether the observed diagnostic performance gains depend on the choice of classifier backbone used in the semantics misalignment filtering stage (refer to Sec.
+6), we performed sensitivity analysis using four backbones across three datasets (Table 10).
+Specifically, we fixed the downstream classifier to I3D and selected four representative backbones in the semantic filtering stage.
+It can be observed that, under both training paradigms on three datasets, the diagnostic performance consistently improves with all backbone selections, and their gains exhibit only marginal variations (with deviations in accuracy gains of at most 1.84% relative to our default setting).
+Overall, these results demonstrate that the performance benefits are not tied to a particular semantic filtering backbone, affirming the robustness and general applicability of our method.
+Figure 9: Diagnostic accuracy (%) comparison using I3D (Carreira and Zisserman, 2017) and FTC (Ahmadi et al., 2023) on three datasets under two annotation settings, both trained under the Real-finetune paradigm.
+Left: generator training with varying proportions of class-labeled real data.
+Right: downstream classifiers trained with different fractions of real labeled data.
+The gray dashed line indicates the Baseline result.
+Our Ctrl-GenAug can substantially reduce the overall annotation burden across both the generator training (Left) and downstream diagnosis stages (Right).
+4.8.6 Analysis of Annotation Workload Reduction
+To investigate the advantages of Ctrl-GenAug in reducing annotation burden, we conducted two dedicated ablation studies, including 1) examining the synthesizer’s dependency on real annotations during training (Figure 9, left), and 2) assessing whether the trained generator can reduce labeling demand in downstream applications (Figure 9, right).
+For the first study, we trained the generator using 25%, 50%, 75%, and 100% of the class-labeled real data, and subsequently used the corresponding synthetic datasets to compare their impact on downstream diagnostic accuracy under identical classifier settings.
+For the second study, we first obtained a synthetic dataset using the generator trained on all real labeled data, and then trained downstream classifiers with 25%, 50%, 75%, and 100% of the class-labeled real data under the Real-finetune paradigm to compare the resulting diagnostic accuracy.
+We also report the performance of classifiers trained exclusively on synthetic data (i.e., using 0% real annotated samples).
+[…]
 5 Conclusion and discussion
 In this study, we present a new and general diffusion-based generative augmentation framework, named Ctrl-GenAug, to facilitate medical sequence classification by leveraging customized and diagnosis-reliable synthetic sequences.
-To the best of our knowledge, this is the first comprehensive study to investigate the impact of the controllable generative augmentation scheme on aiding medical sequence classification.
+To the best of our knowledge, this is the first comprehensive empirical study on controllable generative augmentation for medical sequence classification.
 To improve the quality of synthetic data for promoting downstream classification, we propose a multimodal conditions-guided medical sequence generator that ensures flexibly controllable synthesis across semantic, sequential, and data distribution aspects.
@@
 This can better connect the synthesis task and the downstream one, thus further enhancing our sequence classification performance.
-Extensive experiments on 3 datasets across multiple organs and modalities, with 11 popular networks trained on 3 paradigms, show the effectiveness and generality of Ctrl-GenAug.
+Extensive experiments on 5 medical datasets spanning 4 distinct modalities, including comparisons against 15 augmentation strategies and evaluations using 11 classifiers trained on 3 paradigms, comprehensively validate the robustness and generality of Ctrl-GenAug.
 Furthermore, the empirical analysis validates that Ctrl-GenAug can be effectively leveraged to improve diagnostic performance in underrepresented high-risk populations and out-domain robustness.
-We believe that Ctrl-GenAug can serve as a robust and practical data augmentation tool for various clinical scenarios.
-We then discuss the limitations of Ctrl-GenAug.
+We believe that Ctrl-GenAug can serve as a powerful and practical data augmentation tool for various clinical scenarios.
+Even though our experimental results demonstrate that synthetic data can effectively enhance medical AI models, we recognize that clinical adoption ultimately depends on rigorous validation showing that generated sequences are diagnosis-reliable and interpretable to radiologists.
+Expert evaluation provides an intuitive starting point, where radiologists assess fidelity, sequential coherence, and adherence to conditioning cues (Duan et al., 2025).
+Recording clear provenance for each synthetic sequence (such as its associated disease category and image priors) further enhances transparency and interpretability, which is particularly important for conditional generative methods.
+However, human review alone cannot scale to large synthetic databases, underscoring the need for automated and standardized quality-assessment pipelines.
+Building on prior insights (Huang et al., 2024c), we believe that a hierarchical and multidimensional evaluation suite tailored to medical sequence generation, complemented by small-scale user studies to ensure alignment between automated metrics and radiologists’ judgments, is urgently needed in the future to guarantee data quality and establish clinical credibility.
+This represents an interesting avenue for future exploration.
+It is also worth mentioning that our proposed noisy synthetic data filter contributes an initial effort toward assessing the reliability of synthetic samples before being used in downstream diagnosis, and it can be seamlessly integrated into future large-scale quality-assessment pipelines.
+Ultimately, we suggest that synthetic data should complement rather than replace real clinical data, and its practical use relies on robust quality metrics, rigorous validation protocols, and appropriate regulatory oversight to ensure trustworthy and effective integration into medical AI workflows.
+We further discuss the limitations of Ctrl-GenAug.
 First, the denoising diffusion implicit model (DDIM) sampling process is time-consuming due to the large inference steps.
@@
 Besides, we will incorporate test-time adaptation methods (Huang et al., 2023b; Luo et al., 2023; Huang et al., 2023c; Zhang et al., 2024b) to provide a more user-friendly tool for out-domain applications in clinical settings.
-Last, we will attempt to further extend the proposed method to segmentation and detection tasks (Huang et al., 2025; Tao et al., 2025).
+Last, we will further extend the proposed method to segmentation and detection tasks (Huang et al., 2025; Tao et al., 2025) and work toward establishing a comprehensive benchmark suite tailored for medical sequence generative models.
+Data Availability Statement.
+Our data includes four publicly available datasets: TUSC (Stanford-AIMI, 2021), ACDC (Bernard et al., 2018), MosMedData (Vladzymyrskyy et al., 2020), and MRNet (Bien et al., 2018), as well as a private carotid dataset, which is available from the corresponding author upon reasonable request.
+Acknowledgments.
+This work was supported by the grant from National Natural Science Foundation of China (Nos.
+12326619, 62471313, 62471305), Science and Technology Planning Project of Guangdong Province (Nos.
+2023A0505020002, 2024A1515030143), Frontier Technology Development Program of Jiangsu Province (No.
+BF2024078), Guangdong Basic and Applied Basic Research Foundation (Nos.
+2025A1515011448, 2025A1515011821), Shenzhen Science and Technology Program (No.
+JCYJ20240813143302004), Shenzhen Natural Science Foundation (No.
+JCYJ20250604181708011), the Royal Academy of Engineering, United Kingdom under the RAEng Chair in Emerging Technologies (INSILEXCiET1919/19), the ERC Advanced Grant UKRI Frontier Research Guarantee (INSILICO EP/Y030494/1), the UK Centre of Excellence on in-silico Regulatory Scienceand Innovation (UK CEiRSI) (10139527), the National Institute for Health and CareResearch (NIHR) Manchester Biomedical Research Centre (BRC) (NIHR203308), the BHF Manchester Centre of Research Excellence (RE/24/130017), and the CRUKRadNet Manchester (C1994/A28701).
 Supplementary Material
 Figure S1: Additional synthetic results on Carotid (a-d), TUSC (Stanford-AIMI, 2021) (e-h), and ACDC (Bernard et al., 2018) (i-l).
+We used the multimodal conditions joint training strategy (Huang et al., 2023a) during training and flexibly dropped several during sampling.
+Figure S2: Additional synthetic results on MosMedData (Vladzymyrskyy et al., 2020) (a-d) and MRNet (Bien et al., 2018) (e-h).
 We used the multimodal conditions joint training strategy (Huang et al., 2023a) during training and flexibly dropped several during sampling.
@@
 Study Conditional Control Downstream Task Downstream Arch. Downstream Net Num. Medical Modality Data Type
-(Shang et al., 2023) disease class,
+Shang et al. (2023) disease class,
 image readability
@@
 2 Fundus image
-(Farooq et al., 2024) text classification CNN-based,
+Farooq et al. (2024) text classification CNN-based,
 Transformer-based
 2 Dermoscopy image
-(Zhang et al., 2024c) text, edge segmentation CNN-Transformer hybrid-based 1 US, CT, MRI image
-(Sagers et al., 2023) text classification CNN-based 1 Dermoscopy image
-(Khosravi et al., 2024) attribute, 14 pathology labels classification CNN-based 1 X-ray image
-(Yu et al., 2024a) pathology labels,
+Zhang et al. (2024c) text, edge segmentation CNN-Transformer hybrid-based 1 US, CT, MRI image
+Sagers et al. (2023) text classification CNN-based 1 Dermoscopy image
+Khosravi et al. (2024) attribute, 14 pathology labels classification CNN-based 1 X-ray image
+Yu et al. (2024a) pathology labels,
 lesion bounding boxes,
@@
 classification Transformer-based 1 US image
-(Ktena et al., 2024) disease class, attribute classification CNN-based 1 Histopathology,
+Ktena et al. (2024) disease class, attribute classification CNN-based 1 Histopathology,
 X-ray,
@@
 image
-(Luo et al., 2024) disease class classification CNN-based 10 Fundus,
+Luo et al. (2024) disease class classification CNN-based 10 Fundus,
 Dermoscopy,
@@
 image
-(Li et al., 2024a) multi-scale latent
+Li et al. (2024a) multi-scale latent
 representation
 classification Transformer-based 1 Endoscopy video
-(Gong et al., 2024) mask segmentation CNN-based 1 MRI volume
-(Peng et al., 2024) mask segmentation CNN-based, CNN-Transformer hybrid-based 3 CT volume
-(Kebaili et al., 2024) tumor attribute, mask segmentation CNN-based 1 MRI volume
-(Yu et al., 2024b) mask segmentation CNN-based 1 CT volume
-(Chen et al., 2024) mask segmentation CNN-based,
+Gong et al. (2024) mask segmentation CNN-based 1 MRI volume
+Peng et al. (2024) mask segmentation CNN-based, CNN-Transformer hybrid-based 3 CT volume
+Kebaili et al. (2024) tumor attribute, mask segmentation CNN-based 1 MRI volume
+Yu et al. (2024b) mask segmentation CNN-based 1 CT volume
+Chen et al. (2024) mask segmentation CNN-based,
 CNN-Transformer hybrid-based
 3 CT volume
-(Dorjsembe et al., 2024a) mask segmentation CNN-based 1 MRI volume
-(Huang et al., 2024a) mask segmentation CNN-based 2 OCT volume
+Dorjsembe et al. (2024a) mask segmentation CNN-based 1 MRI volume
+Huang et al. (2024a) mask segmentation CNN-based 2 OCT volume
 Ours disease class, text,